### v22

# The Riemann Hypotheses is false

ABSTRACT. Let $\Theta$ denote the supremum of the real parts of the zeros of the Riemann zeta function. We demonstrate that $\Theta=1$, which entails the existence of infinitely many Riemann zeros off the critical line (thus disproving the Riemann Hypothesis (RH)). The paper is concluded by a brief discussion of why our argument doesn’t work for both Weil and Beurling zeta functions whose analogues of the RH are known to be true.
2020 Mathematics Subject Classifications: 11M26, 11M06.
Keywords and phrases: Riemann zeta function, Riemann Hypothesis, disproof.
Introduction. The Riemann zeta function is a function of the complex variable $s$, defined in the half-plane $\Re(s)>1$ by $\zeta(s):=\sum_{n=1}^{\infty}n^{-s}$ and in the whole complex plane by holomorphic continuation. Euler noticed that for $\Re(s)>1$, $\zeta(s)$ can be expressed as a product $\prod_{p}(1-p^{-s})^{-1}$ over the entire set of primes, which entails that $\zeta(s)\neq 0$ for $\Re(s)>1$. It can be shown that $\zeta(s)$ extends to $\mathbb{C}$ as a meromorphic function with only a simple pole at $s=1$, with residue $1$, and satisfies the functional equation $\xi(s)=\xi(1-s)$, where $\xi(z)=\frac{1}{2}z(z-1)\pi^{-z/2}\Gamma(\frac{1}{2}z)\zeta(z)$ and $\Gamma(w)=\int_{0}^{\infty}e^{-x}x^{w-1}\mathrm{d}x$. Let $\chi$ be a primitive Dirichlet character and $L(s,\chi)$ be the corresponding L-function. Most of the analytic properties for $\zeta$ are also known to hold for $L(s,\chi)$ [5, Chapter 10]. Let $\Theta_{\chi}$ be the supremum of the real parts of the zeros of $L(s,\chi)$. The Generalised Riemann Hypothesis (GRH) is equivalent to the statement that $\Theta_{\chi}=\frac{1}{2}$ for each $\chi$, and the Generalised Prime Number Theorem is equivalent to the statement that $L(s,\chi)\neq 0$ at $\Re(s)=1$ for each $\chi$. For a far more thorough discussion of the RH and GRH, the interested reader is kindly referred to Borwein et al [1].
Definitions. Let: $\rho$ denote a non-trivial zero of the Riemann zeta function, $\gamma=\Im(\rho)$ and $\Theta$ be the supremum of the real parts of the $\rho$’s. Denote by $\Lambda$ the von Mangoldt function.
Theorem 1. One has $\Theta=1$.
Proof.
Fix $\varepsilon\in(0,1)$. Let $T\in\mathbb{R}_{>0}$ be arbitrarily large and let neither $T$ nor $T/\log^{2}T$ be an integer. By Lemma 12.2 of [5], note that we can choose $T$ and $H\in[2T,2T+1]$ such that
$\max\Big{(}\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iT)\Big{|},\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iH)\Big{|}\Big{)}\ll\log^{2}T$ (1)
uniformly for $-1\leq\sigma\leq 2$. For $n\geq 2$, the residue theorem gives us
$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{1+\frac{1}{\log n}+iT}^{1+\frac{1}{\log n}+iH}+\int_{1+\frac{1}{\log n}+iH}^{-\varepsilon+iH}+\int_{-\varepsilon+iH}^{-\varepsilon+iT}+\int_{-\varepsilon+iT}^{1+\frac{1}{\log n}+iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
$=I_{1}+I_{2}+I_{3}+I_{4},$ (3)
say. For $0<a<b\leq 2a$ and $t\in\mathbb{R}_{>0}$, note that
$\sum_{a<n\leq b}n^{it}\ll t^{1/2}+at^{-1/2}$ (4)
[6, p.107]. Hence for $t\asymp T,$ we have
$\sum_{\frac{T}{\log^{2}T}<n\leq T}n^{it}$ $=\Big{(}\sum_{\frac{T}{\log^{2}T}<n\leq\frac{2T}{\log^{2}T}}+\sum_{\frac{2T}{\log^{2}T}<n\leq\frac{4T}{\log^{2}T}}+\cdots+\sum_{\frac{T}{2}<n\leq T}\Big{)}n^{it}$ (5)
$\ll T^{\frac{1}{2}}\log^{4}T.$ (6)
By partial summation, one deduces from (6) that
$\sum_{\frac{T}{\log^{2}T}<n\leq T}n^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{4}T$ (7)
for $\sigma\in\mathbb{R},|\sigma|\leq A$ where $A>0$ is a constant. Combining (1) and (7) yields $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{6}T.$ By the log-derivative of the functional equation for $\zeta$ (see e.g. [3, p.403]), note that $\frac{\zeta^{\prime}}{\zeta}(-\varepsilon+it)\ll\log t$ uniformly for $t\geq 2$. Together with (7), this implies that $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{3}\ll T^{1/2}.$ Thus we arrive at
$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{4}T.$ (8)
Note that $-\frac{\zeta^{\prime}}{\zeta}(s)=\sum_{m=2}^{\infty}\frac{\Lambda(m)}{m^{s}}$ for $\Re(s)>1$. Putting this into $I_{1}$ gives
$I_{1}=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{e\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{e\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
Recall that neither $T$ nor $T/\log^{2}T$ is an integer, thus both $\frac{1}{\log(T/m)}$ and $\frac{1}{\log(T/m\log^{2}T)}$ are well-defined for every positive integer $m$. Hence by partial summation, it follows from (6) that for $m\neq n$, one has
$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{iT}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}|\log(T/m)|}.$ (10)
See Appendix 1 for a more rigorous derivation of (10) from (6). Recall that $H=2T+O(1)$. Thus, summing (9) for $\frac{T}{\log^{2}T}<n\leq T$ and invoking (10) yields
$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{1}$ $=-\frac{T}{2\pi}\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{\Lambda(n)}{n}+O\Bigg{(}T^{\frac{1}{2}}\log^{4}T\sum_{m=2}^{\infty}\frac{\Lambda(m)}{m^{1+\frac{1}{\log T}}|\log(T/m)|}\Bigg{)}$ (11)
$=-\frac{T}{2\pi}\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{\Lambda(n)}{n}+O(T^{\frac{1}{2}}\log^{6}T)$ (12)
$=-\frac{T}{\pi}\log\log T+o(T)$ (13)
since $\sum_{m<x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ [5, p.182], where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant. By Theorem 1 of [3], we have
$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ (14)
uniformly for $2\leq n\ll T.$ Define $N(y):=\sum_{0<\gamma\leq y}1$. Note that $N(y)=(y/2\pi)\log(y/2\pi)-(y/2\pi)+O(\log y)$ [5, Corollary 14.3]. Thus from (14) we obtain
$\sum_{1\leq n\leq\frac{T}{\log^{2}T}}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=N(H)-N(T)-\frac{T}{2\pi}\sum_{2\leq n\leq\frac{T}{\log^{2}T}}\frac{\Lambda(n)}{n}+o(T)$ (15)
$=\frac{T}{\pi}\log\log T+\frac{T}{2\pi}(B-1+\log 2)+o(T),$ (16)
since $H=2T+O(1)$. Combining (8), (13) and (16) gives
$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}$ (17)
$=\Big{(}\sum_{T\leq\gamma_{T}\leq H}\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{2}T}}+\sum_{\frac{T}{\log^{2}T}<n\leq T}\Big{)}\Big{)}n^{\rho_{T}-1}$ (18)
$\gg T.$ (19)
Suppose that $\Theta<1$. Then by Theorem 4.11 of [6], we have
$\sum_{n\leq T}n^{\rho_{T}-1}=\zeta(1-\rho_{T})+\frac{T^{\rho_{T}}}{\rho_{T}}+O(T^{\Theta-1})\ll T^{\Theta-1}$ (20)
uniformly for $T\leq\gamma_{T}\leq H$. Hence
$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}\ll T^{\Theta}\log T.$ (21)
But (21) is a contradiction to (19), thus we deduce that our supposition must be false. ∎
Corollary 1. There exists an infinite sequence of zeros of $\zeta$, whose real parts converge to 1.
Proof.
Combine Theorem 1 with the fact that $\zeta(s)\neq 0$ for $\Re(s)\geq 1$ [5, Theorem 6.6]. ∎
Appendix 1. Proof of (10) from (6):
Proof.
For $n\geq m+1$, note that $\frac{1}{\log(n/m)}=\frac{1}{\log n\Big{(}1-\frac{\log m}{\log n}\Big{)}}=\frac{1}{\log n}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log n}\Big{)}^{k}.$ Thus
$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}=\sum_{k=0}^{\infty}(\log m)^{k}\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{iT}}{m^{1+\frac{1}{\log n}+iT}(\log n)^{k+1}}.$ (22)
By partial summation, it follows from (6) that
$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{iT}}{m^{1+\frac{1}{\log n}+iT}(\log n)^{k+1}}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}|\log(T/m)|(\log T)^{k+1}}.$ (23)
Inserting (23) into the right-hand side of (22) and invoking the fact that $\frac{1}{\log T}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log T}\Big{)}^{k}=\frac{1}{\log(T/m)}$ gives
$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}\log(T/m)}$ (24)
as desired, for $T>n\geq m+1\geq\frac{T}{\log^{2}T}$. The case $\frac{T}{\log^{2}T}<n\leq m-1<T$ is treated similarly. ∎
Remark 1. Notice that our argument actually shows that, for all large enough $T\in\mathbb{R}^{+}$, there exists some $\rho_{T}$ with $T\leq\gamma_{T}\leq 2T+1$ and $\Re(\rho_{T})=1-o(1)$.
Remark 2. Our argument crucially relies on:
(i) The functional equation for $\zeta$, and
(ii) Theorem 4.11 of [6].
Both (i) and (ii) do not hold for Beurling zeta functions corresponding to Beurling number systems whose integers are not the ordinary integers. Hence our approach wouldn’t work for such, whose analogues of the RH are known to be true. Note that (ii) also doesn’t hold for the Weil zeta functions, which are also known to satisfy the RH.

### v24

@@
 $=I_{1}+I_{2}+I_{3}+I_{4},$ (3)
-say. For $0<a<b\leq 2a$ and $t\in\mathbb{R}_{>0}$, note that
+say. From now on, let $T\leq t\leq H$. For $0<a<b\leq 2a$, note that
 $\sum_{a<n\leq b}n^{it}\ll t^{1/2}+at^{-1/2}$ (4)
-[6, p.107]. Hence for $t\asymp T,$ we have
+[6, p.107]. Hence
 $\sum_{\frac{T}{\log^{2}T}<n\leq T}n^{it}$ $=\Big{(}\sum_{\frac{T}{\log^{2}T}<n\leq\frac{2T}{\log^{2}T}}+\sum_{\frac{2T}{\log^{2}T}<n\leq\frac{4T}{\log^{2}T}}+\cdots+\sum_{\frac{T}{2}<n\leq T}\Big{)}n^{it}$ (5)
@@
 Combine Theorem 1 with the fact that $\zeta(s)\neq 0$ for $\Re(s)\geq 1$ [5, Theorem 6.6]. ∎
-Appendix 1. Proof of (10) from (6):
+Appendix 1. Derivation of (10) from (6):
 Proof.
-For $n\geq m+1$, note that $\frac{1}{\log(n/m)}=\frac{1}{\log n\Big{(}1-\frac{\log m}{\log n}\Big{)}}=\frac{1}{\log n}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log n}\Big{)}^{k}.$ Thus
-$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}=\sum_{k=0}^{\infty}(\log m)^{k}\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{iT}}{m^{1+\frac{1}{\log n}+iT}(\log n)^{k+1}}.$ (22)
-By partial summation, it follows from (6) that
-$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{iT}}{m^{1+\frac{1}{\log n}+iT}(\log n)^{k+1}}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}|\log(T/m)|(\log T)^{k+1}}.$ (23)
-Inserting (23) into the right-hand side of (22) and invoking the fact that $\frac{1}{\log T}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log T}\Big{)}^{k}=\frac{1}{\log(T/m)}$ gives
-$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}\log(T/m)}$ (24)
-as desired, for $T>n\geq m+1\geq\frac{T}{\log^{2}T}$. The case $\frac{T}{\log^{2}T}<n\leq m-1<T$ is treated similarly. ∎
+For $T>n\geq m+1\geq\frac{T}{\log^{2}T}$, note that $\frac{1}{\log(n/m)}=\frac{1}{\log n\Big{(}1-\frac{\log m}{\log n}\Big{)}}=\frac{1}{\log n}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log n}\Big{)}^{k}.$ Thus
+$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=m^{-it}\sum_{k=0}^{\infty}(\log m)^{k}\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}(\log n)^{k+1}}.$ (22)
+Define $c_{n}:=n^{it},C(x):=\sum_{n\leq x}c_{n}$ and $f(n)=\frac{1}{m^{1+\frac{1}{\log n}}(\log n)^{k+1}}$. By partial summation, note that
+$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}(\log n)^{k+1}}$ $=\sum_{m+1\leq n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}(\log n)^{k+1}}$ (23)
+$=\int_{m+1}^{T}(C(x)-C(m+1))f^{\prime}(x)\mathrm{d}x$ (24)
+$=\Bigg{(}\int_{m+1}^{2(m+1)}+\int_{2(m+1)}^{4(m+1)}+\cdots+\int_{T/2}^{T}\Bigg{)}(C(x)-C(m+1))f^{\prime}(x)\mathrm{d}x$ (25)
+$\ll T^{1/2}\log^{4}T\Bigg{(}\int_{m+1}^{2(m+1)}+\int_{2(m+1)}^{4(m+1)}+\cdots+\int_{T/2}^{T}\Bigg{)}|f^{\prime}(x)|\mathrm{d}x$ (26)
+$\ll\frac{T^{1/2}\log^{4}T}{m^{1+\frac{1}{\log T}}\log^{k+1}(\frac{T}{\log^{2}T})}$ (27)
+where (26) was obtained from (25) by applying (6), which entails that $C(x)-C(m+1)\ll T^{1/2}\log^{4}T$ uniformly for $m+1\leq x\leq T$.
+Inserting (26) into the right-hand side of (22) and invoking the fact that
+$\frac{1}{\log(T/\log^{2}T)}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log(T/\log^{2}T)}\Big{)}^{k}=\frac{1}{\log(T/m\log^{2}T)},$ (28)
+gives
+$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}$ $\ll\frac{T^{1/2}\log^{4}T}{m^{1+\frac{1}{\log T}}\log(\frac{T}{m\log^{2}T})}$ (29)
+$\ll\frac{T^{1/2}\log^{4}T}{m^{1+\frac{1}{\log T}}\log(T/m)}$ (30)
+as desired. This completes the proof for $T>n\geq m+1\geq\frac{T}{\log^{2}T}$. The case $\frac{T}{\log^{2}T}<n\leq m-1<T$ is treated similarly. ∎
 Remark 1. Notice that our argument actually shows that, for all large enough $T\in\mathbb{R}^{+}$, there exists some $\rho_{T}$ with $T\leq\gamma_{T}\leq 2T+1$ and $\Re(\rho_{T})=1-o(1)$.

### v25

@@
 Definitions.
-Let: $\rho$ denote a non-trivial zero of the Riemann zeta function, $\gamma=\Im(\rho)$ and $\Theta$ be the supremum of the real parts of the $\rho$’s.
+Let: $\rho$ denote a non-trivial zero of the Riemann zeta function, $\gamma=\Im(\rho)$ and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s.
 Denote by $\Lambda$ the von Mangoldt function.
@@
 Proof.
-Fix $\varepsilon\in(0,1)$.
-Let $T\in\mathbb{R}_{>0}$ be arbitrarily large and let neither $T$ nor $T/\log^{2}T$ be an integer.
-By Lemma 12.2 of [5], note that we can choose $T$ and $H\in[2T,2T+1]$ such that
-$\max\Big{(}\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iT)\Big{|},\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iH)\Big{|}\Big{)}\ll\log^{2}T$ (1)
-uniformly for $-1\leq\sigma\leq 2$. For $n\geq 2$, the residue theorem gives us
-$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{1+\frac{1}{\log n}+iT}^{1+\frac{1}{\log n}+iH}+\int_{1+\frac{1}{\log n}+iH}^{-\varepsilon+iH}+\int_{-\varepsilon+iH}^{-\varepsilon+iT}+\int_{-\varepsilon+iT}^{1+\frac{1}{\log n}+iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
+Suppose that $\Theta<1$ and fix $0<\varepsilon<1-\Theta$.
+Let $T\in\mathbb{R}_{>0}$ be arbitrarily large and $X\geq 2$.
+By Lemma 12.2 of [5], we have $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX)\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$.
+Thus we can choose $T$ and $H\in[3T,3T+1]$ such that
+$\max\Big{(}\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+2iT)\Big{|},\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iH)\Big{|}\Big{)}\ll\log^{2}T$ (1)
+uniformly for $-1\leq\sigma\leq 2$. By the residue theorem, we have
+$\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{\Theta+\varepsilon+2iT}^{\Theta+\varepsilon+iH}+\int_{\Theta+\varepsilon+iH}^{-1+iH}+\int_{-1+iH}^{-1+2iT}+\int_{-1+2iT}^{\Theta+\varepsilon+2iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
 $=I_{1}+I_{2}+I_{3}+I_{4},$ (3)
-say. From now on, let $T\leq t\leq H$. For $0<a<b\leq 2a$, note that
-$\sum_{a<n\leq b}n^{it}\ll t^{1/2}+at^{-1/2}$ (4)
+say. From now on, let $2T\leq t\leq H$ and $k=k(\varepsilon)>4$ be a fixed real number such that $k(1-\Theta-\varepsilon)>2$. For $0<a<b\leq 2a$, note that
+$\sum_{a\leq n\leq b}n^{it}\ll t^{1/2}+at^{-1/2}$ (4)
 [6, p.107]. Hence
-$\sum_{\frac{T}{\log^{2}T}<n\leq T}n^{it}$ $=\Big{(}\sum_{\frac{T}{\log^{2}T}<n\leq\frac{2T}{\log^{2}T}}+\sum_{\frac{2T}{\log^{2}T}<n\leq\frac{4T}{\log^{2}T}}+\cdots+\sum_{\frac{T}{2}<n\leq T}\Big{)}n^{it}$ (5)
-$\ll T^{\frac{1}{2}}\log^{4}T.$ (6)
+$\sum_{\frac{T}{\log^{k}T}<n\leq T}n^{it}$ $=\Big{(}\sum_{\frac{T}{\log^{k}T}<n\leq\frac{2T}{\log^{k}T}}+\sum_{\frac{2T}{\log^{k}T}<n\leq\frac{4T}{\log^{k}T}}+\cdots+\sum_{\frac{T}{2}<n\leq T}\Big{)}n^{it}$ (5)
+$\ll T^{\frac{1}{2}}\log^{k+1}T.$ (6)
 By partial summation, one deduces from (6) that
-$\sum_{\frac{T}{\log^{2}T}<n\leq T}n^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{4}T$ (7)
+$\sum_{\frac{T}{\log^{k}T}<n\leq T}n^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{k+1}T$ (7)
 for $\sigma\in\mathbb{R},|\sigma|\leq A$ where $A>0$ is a constant.
-Combining (1) and (7) yields $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{6}T.$ By the log-derivative of the functional equation for $\zeta$ (see e.g.
-[3, p.403]), note that $\frac{\zeta^{\prime}}{\zeta}(-\varepsilon+it)\ll\log t$ uniformly for $t\geq 2$.
-Together with (7), this implies that $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{3}\ll T^{1/2}.$ Thus we arrive at
-$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{4}T.$ (8)
-Note that $-\frac{\zeta^{\prime}}{\zeta}(s)=\sum_{m=2}^{\infty}\frac{\Lambda(m)}{m^{s}}$ for $\Re(s)>1$. Putting this into $I_{1}$ gives
-$I_{1}=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{e\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{e\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
-Recall that neither $T$ nor $T/\log^{2}T$ is an integer, thus both $\frac{1}{\log(T/m)}$ and $\frac{1}{\log(T/m\log^{2}T)}$ are well-defined for every positive integer $m$.
-Hence by partial summation, it follows from (6) that for $m\neq n$, one has
-$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}|\log(T/m)|}.$ (10)
-See Appendix 1 for a more rigorous derivation of (10) from (6). Recall that $H=2T+O(1)$. Thus, summing (9) for $\frac{T}{\log^{2}T}<n\leq T$ and invoking (10) yields
-$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{1}$ $=-\frac{T}{2\pi}\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{\Lambda(n)}{n}+O\Bigg{(}T^{\frac{1}{2}}\log^{4}T\sum_{m=2}^{\infty}\frac{\Lambda(m)}{m^{1+\frac{1}{\log T}}|\log(T/m)|}\Bigg{)}$ (11)
-$=-\frac{T}{2\pi}\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{\Lambda(n)}{n}+O(T^{\frac{1}{2}}\log^{6}T)$ (12)
-$=-\frac{T}{\pi}\log\log T+o(T)$ (13)
+Combining (1) and (7) yields $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{2}+I_{4}\ll T^{1/2}.$ Let $\sigma\geq\varepsilon$.
+By the log-derivative of the functional equation for $\zeta$ (see e.g.
+[3, p.403]), note that
+$\frac{\zeta^{\prime}}{\zeta}(-\sigma+it)\ll\log t$ (8)
+uniformly for $t\geq 2$. Combining (7) and (8) gives $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{3}\ll T^{1/2}\log^{k+2}T.$ Thus we arrive at
+$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{k+2}T.$ (9)
+Let $s=\Theta+\varepsilon+iY$ and $c=1-\Theta-\varepsilon+\frac{1}{\log T}>0$, where $Y\in[2T,H]$. By Theorem 3.12 of [6], note that
+$-\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\Lambda(m)}{m^{s}}=\frac{1}{2\pi i}\int_{c-iT}^{c+iT}\frac{\zeta^{\prime}}{\zeta}(s+w)\frac{(T/2\log^{k}T)^{w}}{w}\mathrm{d}w+O(T^{c-1}\log^{1-kc}T).$ (10)
+Notice that the above integral has a simple pole at $w=1-s$ with residue $\frac{\zeta^{\prime}}{\zeta}(s)-\frac{(T/2\log^{k}T)^{1-s}}{1-s}.$ Hence, shifting the line of integration in (10) to $\Re(s)=-1$ then invoking (1) and (8) to bound the integrals from Cauchy’s residue theorem, gives
+$\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\Lambda(m)}{m^{s}}-\frac{(T/2\log^{k}T)^{1-s}}{1-s}+O(T^{c-1}\log^{1-kc}T).$ (11)
+Putting (11) into $I_{1}$ gives
+$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m\leq\frac{T}{2\log^{k}T}}\Bigg{(}\frac{n^{\Theta+\varepsilon-1+iT}\Lambda(m)}{m^{\Theta+\varepsilon+iT}\log(n/m)}-\frac{n^{\Theta+\varepsilon-1+iH}\Lambda(m)}{m^{\Theta+\varepsilon+iH}\log(n/m)}\Bigg{)}$ (12)
+$-\frac{1}{2\pi i}\int_{\Theta+\varepsilon+iT}^{\Theta+\varepsilon+iH}\frac{(T/2n\log^{k}T)^{1-s}}{1-s}\mathrm{d}s+O(n^{\Theta+\varepsilon-1}T^{c}\log^{2-kc}T).$ (13)
+Since $1\leq m\leq\frac{T}{2\log^{k}T}$, note that $0<\frac{1}{\log(n/m)}\leq\frac{1}{\log 2}$ uniformly for $\frac{T}{\log^{k}T}\leq n\leq T$. Hence by partial summation, it follows from (6) that
+$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{n^{iT}}{m^{\Theta+\varepsilon+iT}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{k+1}T}{m^{\Theta+\varepsilon}\log(T/m)}$ (14)
+uniformly for $2\leq m\leq\frac{T}{2\log^{k}T}$.
+The same bound holds for $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{n^{iH}}{m^{\Theta+\varepsilon+iH}\log(n/m)}.$ By partial summation, it also follows from (6) that
+$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\int_{\Theta+\varepsilon+iT}^{\Theta+\varepsilon+iH}\frac{(T/2n\log^{k}T)^{1-s}}{1-s}\mathrm{d}s\ll T^{1/2}\log^{k+1}T.$ (15)
+Recall that $H=3T+O(1)$ and $kc>2$. Thus, summing (12) and (13) for $\frac{T}{\log^{k}T}\leq n\leq T$ yields
+$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{1}$ $=-\frac{T}{2\pi}\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{\Lambda(n)}{n}+O\Bigg{(}T^{\frac{1}{2}}\log^{k+1}T\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\log m}{m^{\Theta+\varepsilon}\log(T/m)}\Bigg{)}+o(T)$ (16)
+$=-\frac{T}{2\pi}\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{\Lambda(n)}{n}+o(T)$ (17)
+$=-\frac{kT}{2\pi}\log\log T+o(T)$ (18)
 since $\sum_{m<x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ [5, p.182], where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant. By Theorem 1 of [3], we have
-$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ (14)
+$\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ (19)
 uniformly for $2\leq n\ll T.$ Define $N(y):=\sum_{0<\gamma\leq y}1$. Note that $N(y)=(y/2\pi)\log(y/2\pi)-(y/2\pi)+O(\log y)$ [5, Corollary 14.3]. Thus from (19) we obtain
-$\sum_{1\leq n\leq\frac{T}{\log^{2}T}}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=N(H)-N(T)-\frac{T}{2\pi}\sum_{2\leq n\leq\frac{T}{\log^{2}T}}\frac{\Lambda(n)}{n}+o(T)$ (15)
-$=\frac{T}{\pi}\log\log T+\frac{T}{2\pi}(B-1+\log 2)+o(T),$ (16)
-since $H=2T+O(1)$. Combining (8), (13) and (16) gives
-$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}$ (17)
-$=\Big{(}\sum_{T\leq\gamma_{T}\leq H}\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{2}T}}+\sum_{\frac{T}{\log^{2}T}<n\leq T}\Big{)}\Big{)}n^{\rho_{T}-1}$ (18)
+$\sum_{1\leq n\leq\frac{T}{\log^{k}T}}\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=N(H)-N(2T)-\frac{T}{2\pi}\sum_{2\leq n\leq\frac{T}{\log^{k}T}}\frac{\Lambda(n)}{n}+o(T)$ (20)
+$=\frac{kT}{2\pi}\log\log T+\frac{T}{2\pi}(B-1+\log 2)+o(T),$ (21)
+since $H=3T+O(1)$. Combining (21), (18) and (9) gives
+$\sum_{1\leq n\leq T}\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\sum_{2T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}$ (22)
+$=\Big{(}\sum_{2T\leq\gamma_{T}\leq H}\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{k}T}}+\sum_{\frac{T}{\log^{k}T}<n\leq T}\Big{)}\Big{)}n^{\rho_{T}-1}$ (23)
 $\gg T.$ (24)
-Suppose that $\Theta<1$. Then by Theorem 4.11 of [6], we have
+By Theorem 4.11 of [6], we have
 $\sum_{n\leq T}n^{\rho_{T}-1}=\zeta(1-\rho_{T})+\frac{T^{\rho_{T}}}{\rho_{T}}+O(T^{\Theta-1})\ll T^{\Theta-1}$ (25)
-uniformly for $T\leq\gamma_{T}\leq H$. Hence
-$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}\ll T^{\Theta}\log T.$ (21)
+uniformly for $2T\leq\gamma_{T}\leq H$. Hence
+$\sum_{1\leq n\leq T}\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=\sum_{2T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}\ll T^{\Theta}\log T.$ (26)
 But (26) is a contradiction to (24), thus we deduce that our supposition must be false. ∎
@@
 Combine Theorem 1 with the fact that $\zeta(s)\neq 0$ for $\Re(s)\geq 1$ [5, Theorem 6.6]. ∎
-Appendix 1. Derivation of (10) from (6):
-Proof.
-For $T>n\geq m+1\geq\frac{T}{\log^{2}T}$, note that $\frac{1}{\log(n/m)}=\frac{1}{\log n\Big{(}1-\frac{\log m}{\log n}\Big{)}}=\frac{1}{\log n}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log n}\Big{)}^{k}.$ Thus
-$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=m^{-it}\sum_{k=0}^{\infty}(\log m)^{k}\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}(\log n)^{k+1}}.$ (22)
-Define $c_{n}:=n^{it},C(x):=\sum_{n\leq x}c_{n}$ and $f(n)=\frac{1}{m^{1+\frac{1}{\log n}}(\log n)^{k+1}}$. By partial summation, note that
-$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}(\log n)^{k+1}}$ $=\sum_{m+1\leq n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}(\log n)^{k+1}}$ (23)
-$=\int_{m+1}^{T}(C(x)-C(m+1))f^{\prime}(x)\mathrm{d}x$ (24)
-$=\Bigg{(}\int_{m+1}^{2(m+1)}+\int_{2(m+1)}^{4(m+1)}+\cdots+\int_{T/2}^{T}\Bigg{)}(C(x)-C(m+1))f^{\prime}(x)\mathrm{d}x$ (25)
-$\ll T^{1/2}\log^{4}T\Bigg{(}\int_{m+1}^{2(m+1)}+\int_{2(m+1)}^{4(m+1)}+\cdots+\int_{T/2}^{T}\Bigg{)}|f^{\prime}(x)|\mathrm{d}x$ (26)
-$\ll\frac{T^{1/2}\log^{4}T}{m^{1+\frac{1}{\log T}}\log^{k+1}(\frac{T}{\log^{2}T})}$ (27)
-where (26) was obtained from (25) by applying (6), which entails that $C(x)-C(m+1)\ll T^{1/2}\log^{4}T$ uniformly for $m+1\leq x\leq T$.
-Inserting (26) into the right-hand side of (22) and invoking the fact that
-$\frac{1}{\log(T/\log^{2}T)}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log(T/\log^{2}T)}\Big{)}^{k}=\frac{1}{\log(T/m\log^{2}T)},$ (28)
-gives
-$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}$ $\ll\frac{T^{1/2}\log^{4}T}{m^{1+\frac{1}{\log T}}\log(\frac{T}{m\log^{2}T})}$ (29)
-$\ll\frac{T^{1/2}\log^{4}T}{m^{1+\frac{1}{\log T}}\log(T/m)}$ (30)
-as desired. This completes the proof for $T>n\geq m+1\geq\frac{T}{\log^{2}T}$. The case $\frac{T}{\log^{2}T}<n\leq m-1<T$ is treated similarly. ∎
-Remark 1. Notice that our argument actually shows that, for all large enough $T\in\mathbb{R}^{+}$, there exists some $\rho_{T}$ with $T\leq\gamma_{T}\leq 2T+1$ and $\Re(\rho_{T})=1-o(1)$.
+Remark 1. Notice that our argument actually shows that, for all large enough $T\in\mathbb{R}^{+}$, there exists some $\rho_{T}$ with $12T\leq\gamma_{T}\leq 3T+$ 1 and $\Re(\rho_{T})=1-o(1)$.
 Remark 2. Our argument crucially relies on:
-(i) The functional equation for $\zeta$, and
-(ii) Theorem 4.11 of [6].
-Both (i) and (ii) do not hold for Beurling zeta functions corresponding to Beurling number systems whose integers are not the ordinary integers.
+(i) The functional equation for $\zeta$,
+(ii) Theorem 3.12 of [6], and
+(iii) Theorem 4.11 of [6].
+Both (i) and (iii) do not hold for Beurling zeta functions corresponding to Beurling number systems whose integers are not the ordinary integers.
 Hence our approach wouldn’t work for such, whose analogues of the RH are known to be true.
-Note that (ii) also doesn’t hold for the Weil zeta functions, which are also known to satisfy the RH.
+Note that both (i) and (ii) also do not hold for the Weil zeta functions, which are also known to satisfy the RH.

### v26

@@
 $=I_{1}+I_{2}+I_{3}+I_{4},$ (3)
-say. From now on, let $2T\leq t\leq H$ and $k=k(\varepsilon)>4$ be a fixed real number such that $k(1-\Theta-\varepsilon)>2$. For $0<a<b\leq 2a$, note that
+say. From now on, let $2T\leq t\leq H$ and $k=k(\varepsilon)>6$ be a fixed real number such that $k(1-\Theta-\varepsilon)\geq 3$. For $0<a<b\leq 2a$, note that
 $\sum_{a\leq n\leq b}n^{it}\ll t^{1/2}+at^{-1/2}$ (4)
@@
 for $\sigma\in\mathbb{R},|\sigma|\leq A$ where $A>0$ is a constant.
-Combining (1) and (7) yields $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{2}+I_{4}\ll T^{1/2}.$ Let $\sigma\geq\varepsilon$.
+Combining (1) and (7) yields $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{2}+I_{4}\ll T^{1/2}.$ Let $\sigma\geq\varepsilon/2$.
 By the log-derivative of the functional equation for $\zeta$ (see e.g.
 [3, p.403]), note that
-$\frac{\zeta^{\prime}}{\zeta}(-\sigma+it)\ll\log t$ (8)
-uniformly for $t\geq 2$. Combining (7) and (8) gives $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{3}\ll T^{1/2}\log^{k+2}T.$ Thus we arrive at
+$\frac{\zeta^{\prime}}{\zeta}(-\sigma+it)\ll\log t.$ (8)
+Combining (7) and (8) gives $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{3}\ll T^{1/2}\log^{k+2}T.$ Thus we arrive at
 $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{k+2}T.$ (9)
 Let $s=\Theta+\varepsilon+iY$ and $c=1-\Theta-\varepsilon+\frac{1}{\log T}>0$, where $Y\in[2T,H]$. By Theorem 3.12 of [6], note that
-$-\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\Lambda(m)}{m^{s}}=\frac{1}{2\pi i}\int_{c-iT}^{c+iT}\frac{\zeta^{\prime}}{\zeta}(s+w)\frac{(T/2\log^{k}T)^{w}}{w}\mathrm{d}w+O(T^{c-1}\log^{1-kc}T).$ (10)
-Notice that the above integral has a simple pole at $w=1-s$ with residue $\frac{\zeta^{\prime}}{\zeta}(s)-\frac{(T/2\log^{k}T)^{1-s}}{1-s}.$ Hence, shifting the line of integration in (10) to $\Re(s)=-1$ then invoking (1) and (8) to bound the integrals from Cauchy’s residue theorem, gives
-$\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\Lambda(m)}{m^{s}}-\frac{(T/2\log^{k}T)^{1-s}}{1-s}+O(T^{c-1}\log^{1-kc}T).$ (11)
+$-\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\Lambda(m)}{m^{s}}=\frac{1}{2\pi i}\int_{c-4iT}^{c+4iT}\frac{\zeta^{\prime}}{\zeta}(s+w)\frac{(T/2\log^{k}T)^{w}}{w}\mathrm{d}w+O(T^{c-1}\log^{1-kc}T).$ (10)
+Notice that the above integrand has simple poles at $w=0,w=\rho$ and $w=1-s$.
+Hence, shifting the line of integration in (10) to $\Re(w)=-1$ then invoking (1) and (8) to bound the integrals from Cauchy’s residue theorem, gives
+$\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\Lambda(m)}{m^{s}}-\frac{(T/2\log^{k}T)^{1-s}}{1-s}+\sum_{2T\leq\gamma_{T}\leq H}\frac{(T/2\log^{k}T)^{\rho-s}}{\rho-s}+O(T^{c-1}\log^{1-kc}T).$ (11)
 Putting (11) into $I_{1}$ gives
-$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m\leq\frac{T}{2\log^{k}T}}\Bigg{(}\frac{n^{\Theta+\varepsilon-1+iT}\Lambda(m)}{m^{\Theta+\varepsilon+iT}\log(n/m)}-\frac{n^{\Theta+\varepsilon-1+iH}\Lambda(m)}{m^{\Theta+\varepsilon+iH}\log(n/m)}\Bigg{)}$ (12)
-$-\frac{1}{2\pi i}\int_{\Theta+\varepsilon+iT}^{\Theta+\varepsilon+iH}\frac{(T/2n\log^{k}T)^{1-s}}{1-s}\mathrm{d}s+O(n^{\Theta+\varepsilon-1}T^{c}\log^{2-kc}T).$ (13)
+$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m\leq\frac{T}{2\log^{k}T}}\Bigg{(}\frac{n^{\Theta+\varepsilon-1+2iT}\Lambda(m)}{m^{\Theta+\varepsilon+2iT}\log(n/m)}-\frac{n^{\Theta+\varepsilon-1+iH}\Lambda(m)}{m^{\Theta+\varepsilon+iH}\log(n/m)}\Bigg{)}$ (12)
+$-\frac{1}{2\pi i}\int_{\Theta+\varepsilon+2iT}^{\Theta+\varepsilon+iH}\frac{(T/2n\log^{k}T)^{1-s}}{1-s}\mathrm{d}s+\frac{1}{2\pi i}\sum_{2T\leq\gamma_{T}\leq H}(T/2\log^{k}T)^{\rho_{T}}\int_{\Theta+\varepsilon+2iT}^{\Theta+\varepsilon+iH}\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s$ (13)
+$+O(n^{\Theta+\varepsilon-1}T^{c}\log^{2-kc}T).$ (14)
+By the residue theorem again, we have
+$\frac{1}{2\pi i}\int_{\Theta+\varepsilon+2iT}^{\Theta+\varepsilon+iH}\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s$ $=\sum_{2T\leq\gamma_{T}\leq H}(T/2\log^{k}T)^{-\rho_{T}}n^{\rho_{T}-1}-\frac{1}{2\pi i}\Bigg{(}\int_{\Theta+\varepsilon+iH}^{\Theta+\varepsilon+2iT}+\int_{\Theta+\varepsilon+2iT}^{-\varepsilon+2iT}+$ (15)
+$\int_{-\varepsilon+2iT}^{-\varepsilon+iH}\Bigg{)}\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s$ (16)
+$=\sum_{2T\leq\gamma_{T}\leq H}(T/2\log^{k}T)^{-\rho_{T}}n^{\rho_{T}-1}-\frac{1}{2\pi i}\Bigg{(}\int_{\Theta+\varepsilon+iH}^{\Theta+\varepsilon+2iT}+\int_{\Theta+\varepsilon+2iT}^{-\varepsilon+2iT}\Bigg{)}$ (17)
+$\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s+O(n^{-1-\varepsilon}T^{-\varepsilon}).$ (18)
+Inserting (18), (17), (16) and (15) into the summation in (13), gives
+$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m\leq\frac{T}{2\log^{k}T}}\Bigg{(}\frac{n^{\Theta+\varepsilon-1+2iT}\Lambda(m)}{m^{\Theta+\varepsilon+2iT}\log(n/m)}-\frac{n^{\Theta+\varepsilon-1+iH}\Lambda(m)}{m^{\Theta+\varepsilon+iH}\log(n/m)}\Bigg{)}$ (19)
+$-\frac{1}{2\pi i}\int_{\Theta+\varepsilon+2iT}^{-\varepsilon+2iT}\frac{(T/2n\log^{k}T)^{1-s}}{1-s}\mathrm{d}s+\sum_{2T\leq\gamma_{T}\leq H}(n^{\rho_{T}-1}+O(n^{-1-\varepsilon}T^{-\varepsilon})).$ (20)
 Since $1\leq m\leq\frac{T}{2\log^{k}T}$, note that $0<\frac{1}{\log(n/m)}\leq\frac{1}{\log 2}$ uniformly for $\frac{T}{\log^{k}T}\leq n\leq T$. Hence by partial summation, it follows from (6) that
-$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{n^{iT}}{m^{\Theta+\varepsilon+iT}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{k+1}T}{m^{\Theta+\varepsilon}\log(T/m)}$ (14)
+$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{n^{2iT}}{m^{\Theta+\varepsilon+2iT}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{k+1}T}{m^{\Theta+\varepsilon}\log(T/m)}$ (21)
 uniformly for $2\leq m\leq\frac{T}{2\log^{k}T}$.
-The same bound holds for $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{n^{iH}}{m^{\Theta+\varepsilon+iH}\log(n/m)}.$ By partial summation, it also follows from (6) that
-$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\int_{\Theta+\varepsilon+iT}^{\Theta+\varepsilon+iH}\frac{(T/2n\log^{k}T)^{1-s}}{1-s}\mathrm{d}s\ll T^{1/2}\log^{k+1}T.$ (15)
-Recall that $H=3T+O(1)$ and $kc>2$. Thus, summing (12) and (13) for $\frac{T}{\log^{k}T}\leq n\leq T$ yields
+The same bound holds for $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{n^{iH}}{m^{\Theta+\varepsilon+iH}\log(n/m)}.$ By partial summation again, it follows from (7) that
+$\sum_{\frac{T}{2\log^{k}T}\leq n\leq T}\int_{\Theta+\varepsilon+2iT}^{-\varepsilon+2iT}\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s\ll T^{1/2}\log^{k+2}T.$ (22)
+By Theorem 4.11 of [6], we have
+$\sum_{n\leq T}n^{\rho_{T}-1}=\zeta(1-\rho_{T})+\frac{T^{\rho_{T}}}{\rho_{T}}+O(T^{\Theta-1})\ll T^{\Theta-1}$ (23)
+uniformly for $2T\leq\gamma_{T}\leq H$. Hence
+$\sum_{1\leq n\leq T}\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=\sum_{2T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}\ll T^{\Theta}\log T.$ (24)
+Recall that $H=3T+O(1)$ and $kc>3$. Summing (19) and (20) for $\frac{T}{\log^{k}T}\leq n\leq T$ and invoking the above estimates therefore yields
 $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{1}$ $=-\frac{T}{2\pi}\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{\Lambda(n)}{n}+O\Bigg{(}T^{\frac{1}{2}}\log^{k+1}T\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\log m}{m^{\Theta+\varepsilon}\log(T/m)}\Bigg{)}+o(T)$ (25)
@@
 $=-\frac{kT}{2\pi}\log\log T+o(T)$ (27)
-since $\sum_{m<x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ [5, p.182], where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant. By Theorem 1 of [3], we have
+since $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ [5, p.182], where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant. By Theorem 1 of [3], we have
 $\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ (28)
@@
 $\gg T.$ (33)
-By Theorem 4.11 of [6], we have
-$\sum_{n\leq T}n^{\rho_{T}-1}=\zeta(1-\rho_{T})+\frac{T^{\rho_{T}}}{\rho_{T}}+O(T^{\Theta-1})\ll T^{\Theta-1}$ (25)
-uniformly for $2T\leq\gamma_{T}\leq H$. Hence
-$\sum_{1\leq n\leq T}\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=\sum_{2T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}\ll T^{\Theta}\log T.$ (26)
 But (33) is a contradiction to (24), thus we deduce that our supposition must be false. ∎

### v27

@@
 Keywords and phrases: Riemann zeta function, Riemann Hypothesis, disproof.
+ [^1] Data availability statement: My manuscript has no associated data.
 Introduction.
@@
 For a far more thorough discussion of the RH and GRH, the interested reader is kindly referred to Borwein et al [1].
-Definitions.
-Let: $\rho$ denote a non-trivial zero of the Riemann zeta function, $\gamma=\Im(\rho)$ and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s.
-Denote by $\Lambda$ the von Mangoldt function.
+Definitions. Let: $\rho$ denote a non-trivial zero of $\zeta$, $\gamma=\Im(\rho)$ and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s.
 Theorem 1. One has $\Theta=1$.
 Proof.
-Suppose that $\Theta<1$ and fix $0<\varepsilon<1-\Theta$.
+Suppose that $\Theta<1$ and fix an arbitrarily small $0<\varepsilon<1-\Theta$.
 Let $T\in\mathbb{R}_{>0}$ be arbitrarily large and $X\geq 2$.
 By Lemma 12.2 of [5], we have $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX)\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$.
-Thus we can choose $T$ and $H\in[3T,3T+1]$ such that
+Thus we can choose $T$ and $H\in[2T,2T+1]$ such that
 $\max\Big{(}\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+2iT)\Big{|},\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iH)\Big{|}\Big{)}\ll\log^{2}T$ (1)
 uniformly for $-1\leq\sigma\leq 2$. By the residue theorem, we have
-$\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{\Theta+\varepsilon+2iT}^{\Theta+\varepsilon+iH}+\int_{\Theta+\varepsilon+iH}^{-1+iH}+\int_{-1+iH}^{-1+2iT}+\int_{-1+2iT}^{\Theta+\varepsilon+2iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
+$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{1+\frac{1}{\log n}+iT}^{1+\frac{1}{\log n}+iH}+\int_{1+\frac{1}{\log n}+iH}^{-1+iH}+\int_{-1+iH}^{-1+iT}+\int_{-1+iT}^{1+\frac{1}{\log n}+iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
 $=I_{1}+I_{2}+I_{3}+I_{4},$ (3)
-say. From now on, let $2T\leq t\leq H$ and $k=k(\varepsilon)>6$ be a fixed real number such that $k(1-\Theta-\varepsilon)\geq 3$. For $0<a<b\leq 2a$, note that
+say. From now on, let $T\leq t\leq H$. For $0<a<b\leq 2a$, note that
 $\sum_{a\leq n\leq b}n^{it}\ll t^{1/2}+at^{-1/2}$ (4)
 [6, p.107]. Hence
-$\sum_{\frac{T}{\log^{k}T}<n\leq T}n^{it}$ $=\Big{(}\sum_{\frac{T}{\log^{k}T}<n\leq\frac{2T}{\log^{k}T}}+\sum_{\frac{2T}{\log^{k}T}<n\leq\frac{4T}{\log^{k}T}}+\cdots+\sum_{\frac{T}{2}<n\leq T}\Big{)}n^{it}$ (5)
-$\ll T^{\frac{1}{2}}\log^{k+1}T.$ (6)
-By partial summation, one deduces from (6) that
-$\sum_{\frac{T}{\log^{k}T}<n\leq T}n^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{k+1}T$ (7)
+$\sum_{\frac{T}{\log^{2}T}<n<T}n^{it}=\Big{(}\sum_{\frac{T}{\log^{2}T}<n\leq\frac{2T}{\log^{2}T}}+\sum_{\frac{2T}{\log^{2}T}<n\leq\frac{4T}{\log^{2}T}}+\cdots+\sum_{T/2<n<T}\Big{)}n^{it}\ll T^{1/2}\log^{3}T.$ (5)
+By partial summation, it follows from (5) that
+$\sum_{\frac{T}{\log^{2}T}<n<T}n^{\sigma-1+it}\ll T^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{3}T$ (6)
 for $\sigma\in\mathbb{R},|\sigma|\leq A$ where $A>0$ is a constant.
-Combining (1) and (7) yields $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{2}+I_{4}\ll T^{1/2}.$ Let $\sigma\geq\varepsilon/2$.
-By the log-derivative of the functional equation for $\zeta$ (see e.g.
-[3, p.403]), note that
-$\frac{\zeta^{\prime}}{\zeta}(-\sigma+it)\ll\log t.$ (8)
-Combining (7) and (8) gives $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{3}\ll T^{1/2}\log^{k+2}T.$ Thus we arrive at
-$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{k+2}T.$ (9)
-Let $s=\Theta+\varepsilon+iY$ and $c=1-\Theta-\varepsilon+\frac{1}{\log T}>0$, where $Y\in[2T,H]$. By Theorem 3.12 of [6], note that
-$-\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\Lambda(m)}{m^{s}}=\frac{1}{2\pi i}\int_{c-4iT}^{c+4iT}\frac{\zeta^{\prime}}{\zeta}(s+w)\frac{(T/2\log^{k}T)^{w}}{w}\mathrm{d}w+O(T^{c-1}\log^{1-kc}T).$ (10)
-Notice that the above integrand has simple poles at $w=0,w=\rho$ and $w=1-s$.
-Hence, shifting the line of integration in (10) to $\Re(w)=-1$ then invoking (1) and (8) to bound the integrals from Cauchy’s residue theorem, gives
-$\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\Lambda(m)}{m^{s}}-\frac{(T/2\log^{k}T)^{1-s}}{1-s}+\sum_{2T\leq\gamma_{T}\leq H}\frac{(T/2\log^{k}T)^{\rho-s}}{\rho-s}+O(T^{c-1}\log^{1-kc}T).$ (11)
-Putting (11) into $I_{1}$ gives
-$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m\leq\frac{T}{2\log^{k}T}}\Bigg{(}\frac{n^{\Theta+\varepsilon-1+2iT}\Lambda(m)}{m^{\Theta+\varepsilon+2iT}\log(n/m)}-\frac{n^{\Theta+\varepsilon-1+iH}\Lambda(m)}{m^{\Theta+\varepsilon+iH}\log(n/m)}\Bigg{)}$ (12)
-$-\frac{1}{2\pi i}\int_{\Theta+\varepsilon+2iT}^{\Theta+\varepsilon+iH}\frac{(T/2n\log^{k}T)^{1-s}}{1-s}\mathrm{d}s+\frac{1}{2\pi i}\sum_{2T\leq\gamma_{T}\leq H}(T/2\log^{k}T)^{\rho_{T}}\int_{\Theta+\varepsilon+2iT}^{\Theta+\varepsilon+iH}\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s$ (13)
-$+O(n^{\Theta+\varepsilon-1}T^{c}\log^{2-kc}T).$ (14)
-By the residue theorem again, we have
-$\frac{1}{2\pi i}\int_{\Theta+\varepsilon+2iT}^{\Theta+\varepsilon+iH}\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s$ $=\sum_{2T\leq\gamma_{T}\leq H}(T/2\log^{k}T)^{-\rho_{T}}n^{\rho_{T}-1}-\frac{1}{2\pi i}\Bigg{(}\int_{\Theta+\varepsilon+iH}^{\Theta+\varepsilon+2iT}+\int_{\Theta+\varepsilon+2iT}^{-\varepsilon+2iT}+$ (15)
-$\int_{-\varepsilon+2iT}^{-\varepsilon+iH}\Bigg{)}\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s$ (16)
-$=\sum_{2T\leq\gamma_{T}\leq H}(T/2\log^{k}T)^{-\rho_{T}}n^{\rho_{T}-1}-\frac{1}{2\pi i}\Bigg{(}\int_{\Theta+\varepsilon+iH}^{\Theta+\varepsilon+2iT}+\int_{\Theta+\varepsilon+2iT}^{-\varepsilon+2iT}\Bigg{)}$ (17)
-$\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s+O(n^{-1-\varepsilon}T^{-\varepsilon}).$ (18)
-Inserting (18), (17), (16) and (15) into the summation in (13), gives
-$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m\leq\frac{T}{2\log^{k}T}}\Bigg{(}\frac{n^{\Theta+\varepsilon-1+2iT}\Lambda(m)}{m^{\Theta+\varepsilon+2iT}\log(n/m)}-\frac{n^{\Theta+\varepsilon-1+iH}\Lambda(m)}{m^{\Theta+\varepsilon+iH}\log(n/m)}\Bigg{)}$ (19)
-$-\frac{1}{2\pi i}\int_{\Theta+\varepsilon+2iT}^{-\varepsilon+2iT}\frac{(T/2n\log^{k}T)^{1-s}}{1-s}\mathrm{d}s+\sum_{2T\leq\gamma_{T}\leq H}(n^{\rho_{T}-1}+O(n^{-1-\varepsilon}T^{-\varepsilon})).$ (20)
-Since $1\leq m\leq\frac{T}{2\log^{k}T}$, note that $0<\frac{1}{\log(n/m)}\leq\frac{1}{\log 2}$ uniformly for $\frac{T}{\log^{k}T}\leq n\leq T$. Hence by partial summation, it follows from (6) that
-$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{n^{2iT}}{m^{\Theta+\varepsilon+2iT}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{k+1}T}{m^{\Theta+\varepsilon}\log(T/m)}$ (21)
-uniformly for $2\leq m\leq\frac{T}{2\log^{k}T}$.
-The same bound holds for $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{n^{iH}}{m^{\Theta+\varepsilon+iH}\log(n/m)}.$ By partial summation again, it follows from (7) that
-$\sum_{\frac{T}{2\log^{k}T}\leq n\leq T}\int_{\Theta+\varepsilon+2iT}^{-\varepsilon+2iT}\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s\ll T^{1/2}\log^{k+2}T.$ (22)
+Combining (1) and (6) yields $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{4}\ll T^{1/2}.$ By the log-derivative of the functional equation for $\zeta$ (see e.g [3, p.403]), note that
+$\frac{\zeta^{\prime}}{\zeta}(-\sigma+it)\ll\log t$ (7)
+for $\sigma\geq\varepsilon.$ Combining (6) and (7) gives $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{3}\ll T^{1/2}\log^{4}T.$ Thus we arrive at
+$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{4}T.$ (8)
+For $n\geq m+1$, note that $\frac{1}{\log(n/m)}=\frac{1}{\log n\Big{(}1-\frac{\log m}{\log n}\Big{)}}=\frac{1}{\log n}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log n}\Big{)}^{k}.$ Thus we have
+$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=\sum_{k=0}^{\infty}(\log m)^{k}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}(\log n)^{k+1}}.$ (9)
+By partial summation, it follows from (5) that
+$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}(\log n)^{k+1}}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}(\log T)^{k+1}}.$ (10)
+Inserting (10) into the right-hand side of (9) and invoking the fact that $\frac{1}{\log T}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log T}\Big{)}^{k}=\frac{1}{\log(T/m)}$, gives
+$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}\log(T/m)}$ (11)
+for $T>n\geq m+1>\frac{T}{\log^{2}T}$.
+The case $\frac{T}{\log^{2}T}<n\leq m-1<T$ is treated similarly by writing $\frac{1}{\log(n/m)}=-\frac{1}{\log m\Big{(}1-\frac{\log n}{\log m}\Big{)}}=-\frac{1}{\log m}\sum_{k=0}^{\infty}\Big{(}\frac{\log n}{\log m}\Big{)}^{k}$ and proceeding exactly as before, so that the same bound holds.
+Hence for $n\geq m+1$ and $n\leq m-1$, we have
+$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}\log(T/m)}.$ (12)
+For $\Re(s)>1$, note that $-\frac{\zeta^{\prime}}{\zeta}(s)=\sum_{m=1}^{\infty}\frac{\Lambda(m)}{m^{s}}$. Putting this into $I_{1}$ gives
+$I_{1}=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{e\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{e\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (13)
+Summing (13) for $\frac{T}{\log^{2}T}<n<T$ and invoking (12) yields
+$\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+O(T^{1/2}\log^{6}T)$ (14)
+$=-\frac{T}{\pi}\log\log T+o(T)$ (15)
+since $H=2T+O(1)$ and $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant [5, p.182]. By Theorem 1 of [3], we have
+$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ (16)
+uniformly for $2\leq n\ll T$.
+Define $N(y):=\sum_{0<\gamma\leq y}1$.
+Note that $N(y)=(y/2\pi)\log(y/2\pi)-y/2\pi+O(\log y)$ [5, Corollary 14.3].
+Thus, summing (16) for $1\leq n\leq\frac{T}{\log^{2}T}$ gives
+$\sum_{1\leq n\leq\frac{T}{\log^{2}T}}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=N(H)-N(T)-\frac{T}{2\pi}\sum_{2\leq n\leq\frac{T}{\log^{2}T}}\frac{\Lambda(n)}{n}+o(T)$ (17)
+$=\frac{T}{\pi}\log\log T+\frac{T}{2\pi}(B-1+\log 2)+o(T).$ (18)
+Combining (18), (15), (8) and (3) yields
+$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{2}T}}+\sum_{\frac{T}{\log^{2}T}<n\leq T}\Big{)}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ (19)
+$\gg T.$ (20)
 By Theorem 4.11 of [6], we have
-$\sum_{n\leq T}n^{\rho_{T}-1}=\zeta(1-\rho_{T})+\frac{T^{\rho_{T}}}{\rho_{T}}+O(T^{\Theta-1})\ll T^{\Theta-1}$ (23)
-uniformly for $2T\leq\gamma_{T}\leq H$. Hence
-$\sum_{1\leq n\leq T}\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=\sum_{2T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}\ll T^{\Theta}\log T.$ (24)
-Recall that $H=3T+O(1)$ and $kc>3$. Summing (19) and (20) for $\frac{T}{\log^{k}T}\leq n\leq T$ and invoking the above estimates therefore yields
-$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{1}$ $=-\frac{T}{2\pi}\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{\Lambda(n)}{n}+O\Bigg{(}T^{\frac{1}{2}}\log^{k+1}T\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\log m}{m^{\Theta+\varepsilon}\log(T/m)}\Bigg{)}+o(T)$ (25)
-$=-\frac{T}{2\pi}\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{\Lambda(n)}{n}+o(T)$ (26)
-$=-\frac{kT}{2\pi}\log\log T+o(T)$ (27)
-since $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ [5, p.182], where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant. By Theorem 1 of [3], we have
-$\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ (28)
-uniformly for $2\leq n\ll T.$ Define $N(y):=\sum_{0<\gamma\leq y}1$. Note that $N(y)=(y/2\pi)\log(y/2\pi)-(y/2\pi)+O(\log y)$ [5, Corollary 14.3]. Thus from (28) we obtain
-$\sum_{1\leq n\leq\frac{T}{\log^{k}T}}\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=N(H)-N(2T)-\frac{T}{2\pi}\sum_{2\leq n\leq\frac{T}{\log^{k}T}}\frac{\Lambda(n)}{n}+o(T)$ (29)
-$=\frac{kT}{2\pi}\log\log T+\frac{T}{2\pi}(B-1+\log 2)+o(T),$ (30)
-since $H=3T+O(1)$. Combining (30), (27) and (9) gives
-$\sum_{1\leq n\leq T}\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\sum_{2T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}$ (31)
-$=\Big{(}\sum_{2T\leq\gamma_{T}\leq H}\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{k}T}}+\sum_{\frac{T}{\log^{k}T}<n\leq T}\Big{)}\Big{)}n^{\rho_{T}-1}$ (32)
-$\gg T.$ (33)
-But (33) is a contradiction to (24), thus we deduce that our supposition must be false. ∎
+$\sum_{1\leq n\leq T}n^{\rho_{T}-1}=\zeta(1-\rho_{T})+\frac{T^{\rho_{T}}}{\rho_{T}}+O(T^{\Theta-1})\ll T^{\Theta-1}$ (21)
+uniformly for $T\leq\gamma_{T}\leq H.$ Thus
+$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}$ (22)
+$\ll T^{\Theta}\log T.$ (23)
+But (23) is a contradiction to (20), thus our supposition must be false. This completes the proof. ∎
 Corollary 1. There exists an infinite sequence of zeros of $\zeta$, whose real parts converge to 1.
@@
 Combine Theorem 1 with the fact that $\zeta(s)\neq 0$ for $\Re(s)\geq 1$ [5, Theorem 6.6]. ∎
-Remark 1. Notice that our argument actually shows that, for all large enough $T\in\mathbb{R}^{+}$, there exists some $\rho_{T}$ with $12T\leq\gamma_{T}\leq 3T+$ 1 and $\Re(\rho_{T})=1-o(1)$.
+Remark 1. Notice that our argument actually shows that, for all large enough $T\in\mathbb{R}^{+}$, there exists some $\rho_{T}$ with $T\leq\gamma_{T}\leq 2T+1$ and $\Re(\rho_{T})=1-o(1)$.
 Remark 2. Our argument crucially relies on:
-(i) The functional equation for $\zeta$,
-(ii) Theorem 3.12 of [6], and
-(iii) Theorem 4.11 of [6].
-Both (i) and (iii) do not hold for Beurling zeta functions corresponding to Beurling number systems whose integers are not the ordinary integers.
+(i) The functional equation for $\zeta$, and
+(ii) Theorem 4.11 of [6]
+Both (i) and (ii) do not hold for Beurling zeta functions corresponding to Beurling number systems whose integers are not the ordinary integers.
 Hence our approach wouldn’t work for such, whose analogues of the RH are known to be true.
-Note that both (i) and (ii) also do not hold for the Weil zeta functions, which are also known to satisfy the RH.
+Also, (ii) does not hold for the Weil zeta functions, which are known to satisfy the RH.

### v28

@@
 Keywords and phrases: Riemann zeta function, Riemann Hypothesis, disproof.
- [^1] Data availability statement: My manuscript has no associated data.
+ [^1] Data availability statement: This manuscript has no associated data.
 Introduction.
@@
 By partial summation, it follows from (5) that
-$\sum_{\frac{T}{\log^{2}T}<n<T}n^{\sigma-1+it}\ll T^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{3}T$ (6)
+$\sum_{\frac{T}{\log^{2}T}<n<T}n^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{3}T$ (6)
 for $\sigma\in\mathbb{R},|\sigma|\leq A$ where $A>0$ is a constant.
@@
 Inserting (10) into the right-hand side of (9) and invoking the fact that $\frac{1}{\log T}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log T}\Big{)}^{k}=\frac{1}{\log(T/m)}$, gives
-$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}\log(T/m)}$ (11)
+$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}|\log(T/m)|}$ (11)
 for $T>n\geq m+1>\frac{T}{\log^{2}T}$.
@@
 Hence for $n\geq m+1$ and $n\leq m-1$, we have
-$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}\log(T/m)}.$ (12)
+$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}|\log(T/m)|}.$ (12)
 For $\Re(s)>1$, note that $-\frac{\zeta^{\prime}}{\zeta}(s)=\sum_{m=1}^{\infty}\frac{\Lambda(m)}{m^{s}}$. Putting this into $I_{1}$ gives
@@
 Also, (ii) does not hold for the Weil zeta functions, which are known to satisfy the RH.
+Remark 3.
+(i) From the literature on arXiv and elsewhere, it doesn’t appear like this approach on RH has been tried before.
+However, the idea of applying Gonek’s formula (16) in estimating sums over zeros of zeta in general, is not new.
+As one can confirm from the literature, it has proven to be a useful tool for such tasks.
+It is therefore instructive not to rush to bound the summation in (13), e.g.
+by applying $n^{it}\ll 1$.
+Doing so will definitely make one lose important information needed for a sharper bound for the sum over $n$.
+(ii) The key part of our argument is the estimate in (12), whose derivation is a routine application of partial summation from (5).
+It actually suffices for the reader to only verify this part.
+Infact for our purposes, the bound in (12) only has to be $\ll\frac{T}{m^{1+\frac{1}{\log T}}(\log^{3}T)|\log(T/m)|}$.

### v29

@@
 Suppose that $\Theta<1$ and fix an arbitrarily small $0<\varepsilon<1-\Theta$.
-Let $T\in\mathbb{R}_{>0}$ be arbitrarily large and $X\geq 2$.
-By Lemma 12.2 of [5], we have $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX)\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$.
+Let $T\in\mathbb{R}_{>0}$ be arbitrarily large.
+Let $X\geq 2$.
+By Lemma 12.2 of [5], we know that there exists some $X_{0}\in[X,X+1]$ such that $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX_{0})\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$.
 Thus we can choose $T$ and $H\in[2T,2T+1]$ such that
@@
 (i) The functional equation for $\zeta$, and
-(ii) Theorem 4.11 of [6]
+(ii) Theorem 4.11 of [6].
 Both (i) and (ii) do not hold for Beurling zeta functions corresponding to Beurling number systems whose integers are not the ordinary integers.

### v30

@@
 Let $\Theta_{\chi}$ be the supremum of the real parts of the zeros of $L(s,\chi)$.
-The Generalised Riemann Hypothesis (GRH) is equivalent to the statement that $\Theta_{\chi}=\frac{1}{2}$ for each $\chi$, and the Generalised Prime Number Theorem is equivalent to the statement that $L(s,\chi)\neq 0$ at $\Re(s)=1$ for each $\chi$.
+The Generalised Riemann Hypothesis (1GRH) is equivalent to the statement that $\Theta_{\chi}=\frac{1}{2}$ for each $\chi$, and the Generalised Prime Number Theorem is equivalent to the statement that $L(s,\chi)\neq 0$ at $\Re(s)=1$ for each $\chi$.
 For a far more thorough discussion of the RH and GRH, the interested reader is kindly referred to Borwein et al [1].
@@
 Proof.
-Suppose that $\Theta<1$ and fix an arbitrarily small $0<\varepsilon<1-\Theta$.
-Let $T\in\mathbb{R}_{>0}$ be arbitrarily large.
-Let $X\geq 2$.
+Suppose that $\Theta<1$.
+Let: $T\in\mathbb{R}_{>0}$ be arbitrarily large and $X\geq 2$.
 By Lemma 12.2 of [5], we know that there exists some $X_{0}\in[X,X+1]$ such that $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX_{0})\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$.
@@
 $\max\Big{(}\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+2iT)\Big{|},\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iH)\Big{|}\Big{)}\ll\log^{2}T$ (1)
-uniformly for $-1\leq\sigma\leq 2$. By the residue theorem, we have
+uniformly for $-1\leq\sigma\leq 2$. Let $n\geq\frac{T}{\log^{2}T}$. By the residue theorem, we have
 $\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{1+\frac{1}{\log n}+iT}^{1+\frac{1}{\log n}+iH}+\int_{1+\frac{1}{\log n}+iH}^{-1+iH}+\int_{-1+iH}^{-1+iT}+\int_{-1+iT}^{1+\frac{1}{\log n}+iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
@@
 for $\sigma\in\mathbb{R},|\sigma|\leq A$ where $A>0$ is a constant.
-Combining (1) and (6) yields $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{4}\ll T^{1/2}.$ By the log-derivative of the functional equation for $\zeta$ (see e.g [3, p.403]), note that
+Combining (1) and (6) yields $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{2}T.$ By the log-derivative of the functional equation for $\zeta$ (see e.g [3, p.403]), note that
 $\frac{\zeta^{\prime}}{\zeta}(-\sigma+it)\ll\log t$ (7)
-for $\sigma\geq\varepsilon.$ Combining (6) and (7) gives $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{3}\ll T^{1/2}\log^{4}T.$ Thus we arrive at
+for $\sigma\geq\sigma_{0}>0.$ Combining (6) and (7) gives $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{3}\ll T^{1/2}\log^{4}T.$ Thus we arrive at
 $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{4}T.$ (8)
-For $n\geq m+1$, note that $\frac{1}{\log(n/m)}=\frac{1}{\log n\Big{(}1-\frac{\log m}{\log n}\Big{)}}=\frac{1}{\log n}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log n}\Big{)}^{k}.$ Thus we have
-$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=\sum_{k=0}^{\infty}(\log m)^{k}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}(\log n)^{k+1}}.$ (9)
-By partial summation, it follows from (5) that
-$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}(\log n)^{k+1}}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}(\log T)^{k+1}}.$ (10)
-Inserting (10) into the right-hand side of (9) and invoking the fact that $\frac{1}{\log T}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log T}\Big{)}^{k}=\frac{1}{\log(T/m)}$, gives
-$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}|\log(T/m)|}$ (11)
-for $T>n\geq m+1>\frac{T}{\log^{2}T}$.
-The case $\frac{T}{\log^{2}T}<n\leq m-1<T$ is treated similarly by writing $\frac{1}{\log(n/m)}=-\frac{1}{\log m\Big{(}1-\frac{\log n}{\log m}\Big{)}}=-\frac{1}{\log m}\sum_{k=0}^{\infty}\Big{(}\frac{\log n}{\log m}\Big{)}^{k}$ and proceeding exactly as before, so that the same bound holds.
-Hence for $n\geq m+1$ and $n\leq m-1$, we have
-$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}|\log(T/m)|}.$ (12)
 For $\Re(s)>1$, note that $-\frac{\zeta^{\prime}}{\zeta}(s)=\sum_{m=1}^{\infty}\frac{\Lambda(m)}{m^{s}}$. Putting this into $I_{1}$ gives
 $I_{1}=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{e\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{e\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
-Summing (13) for $\frac{T}{\log^{2}T}<n<T$ and invoking (12) yields
-$\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+O(T^{1/2}\log^{6}T)$ (14)
+Let $\varepsilon=\varepsilon_{T}=T^{-1/4}$.
+We split the sum on the right-hand side of (9) for $m\leq\frac{n}{1+\varepsilon},\frac{n}{1+\varepsilon}\leq m\leq n-1,n+1\leq m\leq n(1+\varepsilon)$ and $m>n(1+\varepsilon)$.
+For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is uniformly $\ll T^{1/4}$ so it follows from (5) by routine partial summation (Riemann-Stieltjies integration) that
+$\sum_{\frac{T}{\log^{2}T}<n\leq\frac{m}{1+\varepsilon}<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (10)
+and
+$\sum_{\frac{T}{\log^{2}T}<m(1+\varepsilon)\leq n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (11)
+It now remains to treat the ranges $\frac{m}{1+\varepsilon}\leq n\leq m-1$ and $m+1\leq n\leq m(1+\varepsilon)$, where $n\geq\frac{T}{\log^{2}T}$. Note that
+$\sum_{\frac{m}{1+\varepsilon}<n<m-1}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\sum_{\frac{m}{1+\varepsilon}<n<m-1}\frac{1}{m|\log(n/m)|}\ll 1$ (12)
+and
+$\sum_{m+1<n<m(1+\varepsilon)}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\sum_{m+1<n<m(1+\varepsilon)}\frac{1}{m\log(n/m)}\ll 1.$ (13)
+Notice that (13) is obtained by applying the fact that
+$\sum_{m+1<n<m(1+\varepsilon)}\frac{1}{m\log(n/m)}\ll\frac{1}{m}\int_{m+1}^{m(1+\varepsilon)}\frac{\mathrm{d}x}{\log(x/m)}\ll\frac{x/m}{m\log(x/m)}\Bigg{|}_{m+1}^{m(1+\varepsilon)}\ll 1.$ (14)
+A similar argument yields (12). Summing (9) for $\frac{T}{\log^{2}T}<n<T$ and invoking (13), (12), (11) and (10) together with $\sum_{m\leq x}\Lambda(m)=x+O(x^{\Theta}\log^{2}x)$ [5, p.430], yields
+$\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+o(T)$ (15)
 $=-\frac{T}{\pi}\log\log T+o(T)$ (16)
@@
 As one can confirm from the literature, it has proven to be a useful tool for such tasks.
-It is therefore instructive not to rush to bound the summation in (13), e.g.
-by applying $n^{it}\ll 1$.
-Doing so will definitely make one lose important information needed for a sharper bound for the sum over $n$.
-(ii) The key part of our argument is the estimate in (12), whose derivation is a routine application of partial summation from (5).
+It is therefore instructive to sum (9) carefully, e.g.
+by not rushing to naively bound $n^{it}$ by 1 for $n$ not too close to $m$.
+Doing so will definitely make one lose important information needed for a sufficiently sharp bound for the sum over $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$.
+(ii) The key parts of our argument, are the estimates in (10) and (11), whose respective derivations are routine applications of partial summation from (5).
 It actually suffices for the reader to only verify this part.
-Infact for our purposes, the bound in (12) only has to be $\ll\frac{T}{m^{1+\frac{1}{\log T}}(\log^{3}T)|\log(T/m)|}$.

### v31

@@
-The Riemann Hypotheses is false
+The Riemann Hypothesis is false
 ABSTRACT.
@@
 Let $\Theta_{\chi}$ be the supremum of the real parts of the zeros of $L(s,\chi)$.
-The Generalised Riemann Hypothesis (1GRH) is equivalent to the statement that $\Theta_{\chi}=\frac{1}{2}$ for each $\chi$, and the Generalised Prime Number Theorem is equivalent to the statement that $L(s,\chi)\neq 0$ at $\Re(s)=1$ for each $\chi$.
+The Generalised Riemann Hypothesis (GRH) is equivalent to the statement that $\Theta_{\chi}=\frac{1}{2}$ for each $\chi$, and the Generalised Prime Number Theorem is equivalent to the statement that $L(s,\chi)\neq 0$ at $\Re(s)=1$ for each $\chi$.
 For a far more thorough discussion of the RH and GRH, the interested reader is kindly referred to Borwein et al [1].
@@
 Thus we can choose $T$ and $H\in[2T,2T+1]$ such that
-$\max\Big{(}\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+2iT)\Big{|},\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iH)\Big{|}\Big{)}\ll\log^{2}T$ (1)
+$\max\Big{(}\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iT)\Big{|},\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iH)\Big{|}\Big{)}\ll\log^{2}T$ (1)
 uniformly for $-1\leq\sigma\leq 2$. Let $n\geq\frac{T}{\log^{2}T}$. By the residue theorem, we have
@@
 by not rushing to naively bound $n^{it}$ by 1 for $n$ not too close to $m$.
-Doing so will definitely make one lose important information needed for a sufficiently sharp bound for the sum over $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$.
+Doing so will definitely make one lose important information needed for a sufficiently sharp bound for the sums over $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$.
 (ii) The key parts of our argument, are the estimates in (10) and (11), whose respective derivations are routine applications of partial summation from (5).
 It actually suffices for the reader to only verify this part.
+Remark 4.
+The reader might wonder if our approach can give a lower bound for $\Theta_{T}:$ the supremum of the real parts of the zeros $\rho_{T}$ with $T\leq\gamma_{T}\leq H.$ The author thinks that this approach is not powerful enough to do that.
+Indeed, Theorem 4.11 of Titchmarsh [6] only works to approximate $\zeta(\sigma+it)$ with a Dirichlet polynomial of length $>Ct/2\pi$ where $C>1$ is a given constant, provided that $\sigma\geq\sigma_{0}>0.$ We therefore cannot apply that result to find a lower bound for $\Theta_{T},$ since $\lim\inf_{T\rightarrow\infty}\Re(1-\rho_{T})=0$.

### v32

@@
 By partial summation, it follows from (5) that
-$\sum_{\frac{T}{\log^{2}T}<n<T}n^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{3}T$ (6)
-for $\sigma\in\mathbb{R},|\sigma|\leq A$ where $A>0$ is a constant.
-Combining (1) and (6) yields $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{2}T.$ By the log-derivative of the functional equation for $\zeta$ (see e.g [3, p.403]), note that
+$\sum_{\frac{T}{\log^{2}T}<n<T}n^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{7}T$ (6)
+for $\sigma\in\mathbb{R},|\sigma|\leq 2$.
+Combining (1) and (6) yields $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{7}T.$ By the log-derivative of the functional equation for $\zeta$ (see e.g [3, p.403]), note that
 $\frac{\zeta^{\prime}}{\zeta}(-\sigma+it)\ll\log t$ (7)
-for $\sigma\geq\sigma_{0}>0.$ Combining (6) and (7) gives $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{3}\ll T^{1/2}\log^{4}T.$ Thus we arrive at
-$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{4}T.$ (8)
+for $\sigma\geq\sigma_{0}>0.$ Combining (6) and (7) gives $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{3}\ll T^{-1/2}\log^{8}T.$ Thus we arrive at
+$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{7}T.$ (8)
 For $\Re(s)>1$, note that $-\frac{\zeta^{\prime}}{\zeta}(s)=\sum_{m=1}^{\infty}\frac{\Lambda(m)}{m^{s}}$. Putting this into $I_{1}$ gives
 $I_{1}=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{e\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{e\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
-Let $\varepsilon=\varepsilon_{T}=T^{-1/4}$.
-We split the sum on the right-hand side of (9) for $m\leq\frac{n}{1+\varepsilon},\frac{n}{1+\varepsilon}\leq m\leq n-1,n+1\leq m\leq n(1+\varepsilon)$ and $m>n(1+\varepsilon)$.
-For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is uniformly $\ll T^{1/4}$ so it follows from (5) by routine partial summation (Riemann-Stieltjies integration) that
+From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}<n<T$.
+For $m\geq T(1+\varepsilon)$ and $m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll T^{1/4}$.
+Thus it follows from (5) by routine partial summation (Riemann-Stieltjes integration) that for such $m$ we have
+$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (10)
+For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll T^{1/4}$ so it follows again from (5) by partial summation that
 $\sum_{\frac{T}{\log^{2}T}<n\leq\frac{m}{1+\varepsilon}<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (11)
@@
 $\sum_{\frac{T}{\log^{2}T}<m(1+\varepsilon)\leq n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (12)
-It now remains to treat the ranges $\frac{m}{1+\varepsilon}\leq n\leq m-1$ and $m+1\leq n\leq m(1+\varepsilon)$, where $n\geq\frac{T}{\log^{2}T}$. Note that
-$\sum_{\frac{m}{1+\varepsilon}<n<m-1}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\sum_{\frac{m}{1+\varepsilon}<n<m-1}\frac{1}{m|\log(n/m)|}\ll 1$ (12)
+We now treat the ranges $\frac{m}{1+\varepsilon}<n\leq m-1$ and $m+1\leq n<m(1+\varepsilon)$. Note that for $m\ll T$, we have
+$\sum_{\frac{m}{1+\varepsilon}<n\leq m-1}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\sum_{\frac{m}{1+\varepsilon}<n\leq m-1}\frac{1}{m|\log(n/m)|}\ll\log T$ (13)
 and
-$\sum_{m+1<n<m(1+\varepsilon)}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\sum_{m+1<n<m(1+\varepsilon)}\frac{1}{m\log(n/m)}\ll 1.$ (13)
-Notice that (13) is obtained by applying the fact that
-$\sum_{m+1<n<m(1+\varepsilon)}\frac{1}{m\log(n/m)}\ll\frac{1}{m}\int_{m+1}^{m(1+\varepsilon)}\frac{\mathrm{d}x}{\log(x/m)}\ll\frac{x/m}{m\log(x/m)}\Bigg{|}_{m+1}^{m(1+\varepsilon)}\ll 1.$ (14)
-A similar argument yields (12). Summing (9) for $\frac{T}{\log^{2}T}<n<T$ and invoking (13), (12), (11) and (10) together with $\sum_{m\leq x}\Lambda(m)=x+O(x^{\Theta}\log^{2}x)$ [5, p.430], yields
+$\sum_{m+1\leq n<m(1+\varepsilon)}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\sum_{m+1\leq n<m(1+\varepsilon)}\frac{1}{m\log(n/m)}\ll\log T.$ (14)
+Notice that (14) is obtained by writing $n=m+r$ for $11\leq r\leq\varepsilon m-$ 1 and applying the fact that $\log(1+\delta)\sim\delta$ uniformly for $1/m\leq\delta\leq\varepsilon$, hence
+$\sum_{m+1\leq n<m(1+\varepsilon)}\frac{1}{m\log(n/m)}\ll\sum_{1\leq r\leq\varepsilon m-1}\frac{1}{r}\ll\log T.$ (15)
+A similar argument yields (13).
+Summing (9) for $\frac{T}{\log^{2}T}<n<T$ and invoking (14), (13), (12), (11) and (10) together with $\sum_{m\leq x}\Lambda(m)=x+O(x^{\Theta}\log^{2}x)$ [5, p.430], yields
 $\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+o(T)$ (16)

### v33

@@
 From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}<n<T$.
-For $m\geq T(1+\varepsilon)$ and $m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll T^{1/4}$.
+For $m\geq T(1+\varepsilon)$ and $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll T^{1/4}$.
 Thus it follows from (5) by routine partial summation (Riemann-Stieltjes integration) that for such $m$ we have
@@
 Doing so will definitely make one lose important information needed for a sufficiently sharp bound for the sums over $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$.
-(ii) The key parts of our argument, are the estimates in (10) and (11), whose respective derivations are routine applications of partial summation from (5).
+(ii) The key parts of our argument, are the estimates in (10), (11) and (12), whose respective derivations are routine applications of partial summation from (5).
 It actually suffices for the reader to only verify this part.

### v34

@@
 $=I_{1}+I_{2}+I_{3}+I_{4},$ (3)
-say. From now on, let $T\leq t\leq H$. For $0<a<b\leq 2a$, note that
-$\sum_{a\leq n\leq b}n^{it}\ll t^{1/2}+at^{-1/2}$ (4)
-[6, p.107]. Hence
-$\sum_{\frac{T}{\log^{2}T}<n<T}n^{it}=\Big{(}\sum_{\frac{T}{\log^{2}T}<n\leq\frac{2T}{\log^{2}T}}+\sum_{\frac{2T}{\log^{2}T}<n\leq\frac{4T}{\log^{2}T}}+\cdots+\sum_{T/2<n<T}\Big{)}n^{it}\ll T^{1/2}\log^{3}T.$ (5)
-By partial summation, it follows from (5) that
-$\sum_{\frac{T}{\log^{2}T}<n<T}n^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{7}T$ (6)
+say. From now on, let $T\leq t\leq H$. By the approximate functional equation for $\zeta$ (Theorem 4.15 of [6]), note that
+$\zeta(\sigma+it)=\sum_{n\leq x}n^{-\sigma-it}+\chi(\sigma+it)\sum_{n\leq y}n^{\sigma-1+it}+O(x^{-\sigma}\log t+t^{\frac{1}{2}-\sigma}y^{\sigma-1})$ (4)
+where $0\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h$ is a constant and $|\chi(\sigma+it)|=\Big{|}\frac{\zeta(\sigma+it)}{\zeta(1-\sigma-it)}\Big{|}\asymp(t/2\pi)^{1/2-\sigma}$ [6, p.78].
+Thus since $|\zeta(it)|\asymp t^{1/2}|\zeta(1-it)|\ll T^{1/2}\log T$, it follows from (4) that
+$S(x):=\sum_{n\leq x}n^{it}\ll T^{1/2}\log^{3}T$ (5)
+uniformly for $\frac{T}{\log^{2}T}\leq x\leq T$. Let $f(n)=n^{\sigma-1}$. Then by partial summation, we obtain from (5) that
+$\sum_{\frac{T}{\log^{2}T}\leq n\leq T}n^{\sigma-1+it}=S(T)f(T)-S(T/\log^{2}T)f(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)f^{\prime}(x)\mathrm{d}x\ll T^{\sigma-\frac{1}{2}}\log^{7}T$ (6)
 for $\sigma\in\mathbb{R},|\sigma|\leq 2$.
@@
 For $m\geq T(1+\varepsilon)$ and $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll T^{1/4}$.
-Thus it follows from (5) by routine partial summation (Riemann-Stieltjes integration) that for such $m$ we have
-$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (10)
-For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll T^{1/4}$ so it follows again from (5) by partial summation that
+Let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=\frac{m^{-\frac{1}{\log n}-1}(\log(m)\log(n)-\log^{2}(m)-\log^{2}n)}{n\log^{2}(n)\log^{2}(n/m)}$.
+Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) by routine partial summation that
+$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}$ $=S(T)g(T)-S(T/\log^{2}T)g(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)g^{\prime}(x)\mathrm{d}x$ (10)
+$\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}-T^{1/2}\log^{3}T\int_{T/\log^{2}T}^{T}g^{\prime}(x)\mathrm{d}x$ (11)
+$\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (12)
+for $m\geq T(1+\varepsilon)$ and $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$.
+For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll T^{1/4}$ so it similarly follows from (5) by partial summation that
 $\sum_{\frac{T}{\log^{2}T}<n\leq\frac{m}{1+\varepsilon}<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (13)
@@
 Remark 4.
-The reader might wonder if our approach can give a lower bound for $\Theta_{T}:$ the supremum of the real parts of the zeros $\rho_{T}$ with $T\leq\gamma_{T}\leq H.$ The author thinks that this approach is not powerful enough to do that.
-Indeed, Theorem 4.11 of Titchmarsh [6] only works to approximate $\zeta(\sigma+it)$ with a Dirichlet polynomial of length $>Ct/2\pi$ where $C>1$ is a given constant, provided that $\sigma\geq\sigma_{0}>0.$ We therefore cannot apply that result to find a lower bound for $\Theta_{T},$ since $\lim\inf_{T\rightarrow\infty}\Re(1-\rho_{T})=0$.
+The reader might wonder if our approach can give a lower bound for $\Theta_{T}:$ the supremum of the real parts of the zeros $\rho_{T}$ with $T\leq\gamma_{T}\leq H.$ For that, one would need Theorem 4.11 of Titchamrsh [6] to also hold for $\sigma=\sigma_{t}\sim 0$ with an error term $\ll(\log T)^{-1}$.
+However, this is not possible (as can be seen from the approximate functional equation for $\zeta$).
+It can also be seen that Titchmarsh’s proof of his Theorem 4.11 doesn’t hold for $\sigma=\sigma_{t}\sim 0$.
+Indeed, he begins his argument by quoting the Euler-McLaurin summation formula for $\zeta(\sigma+it)$ (equation 3.5.3).
+This is initially defined for $\sigma>1$, and then is extended by holomorphic continuation to $\sigma>0$.
+Recall that a function can only be holomorphically extended on open sets, and that’s precisely why Titchmarsh needs the assumption that $\sigma\geq\sigma_{0}>0$.

### v35

@@
 say. From now on, let $T\leq t\leq H$. By the approximate functional equation for $\zeta$ (Theorem 4.15 of [6]), note that
-$\zeta(\sigma+it)=\sum_{n\leq x}n^{-\sigma-it}+\chi(\sigma+it)\sum_{n\leq y}n^{\sigma-1+it}+O(x^{-\sigma}\log t+t^{\frac{1}{2}-\sigma}y^{\sigma-1})$ (4)
-where $0\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h$ is a constant and $|\chi(\sigma+it)|=\Big{|}\frac{\zeta(\sigma+it)}{\zeta(1-\sigma-it)}\Big{|}\asymp(t/2\pi)^{1/2-\sigma}$ [6, p.78].
+$\zeta(\sigma+it)=\sum_{n\leq x}n^{-\sigma-it}+\chi(\sigma+it)\sum_{v\leq y}v^{\sigma-1+it}+O(x^{-\sigma}\log t+t^{\frac{1}{2}-\sigma}y^{\sigma-1})$ (4)
+where $0\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h$ is a constant and $|\chi(\sigma+it)|=\Big{|}\frac{\zeta(\sigma+it)}{\zeta(1-\sigma-it)}\Big{|}=\Big{|}\Big{(}\frac{2\pi}{t}\Big{)}^{\sigma-\frac{1}{2}+it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{1/2-\sigma}$ [6, p.78].
 Thus since $|\zeta(it)|\asymp t^{1/2}|\zeta(1-it)|\ll T^{1/2}\log T$, it follows from (4) that
@@
 $\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (12)
-for $m\geq T(1+\varepsilon)$ and $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$.
+uniformly for $m\geq T(1+\varepsilon)$ and $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$.
 For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll T^{1/4}$ so it similarly follows from (5) by partial summation that
@@
 A similar argument yields (15).
-Summing (9) for $\frac{T}{\log^{2}T}<n<T$ and invoking (16), (15), (14), (13) and (12) together with $\sum_{m\leq x}\Lambda(m)=x+O(x^{\Theta}\log^{2}x)$ [5, p.430], yields
+Summing (9) for $\frac{T}{\log^{2}T}<n<T$, changing the order of summation and invoking (16), (15), (14), (13) and (12) together with $\sum_{m\leq x}\Lambda(m)=x+O(x^{\Theta}\log^{2}x)$ [5, p.430], yields
 $\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+o(T)$ (18)
@@
 Remark 4.
-The reader might wonder if our approach can give a lower bound for $\Theta_{T}:$ the supremum of the real parts of the zeros $\rho_{T}$ with $T\leq\gamma_{T}\leq H.$ For that, one would need Theorem 4.11 of Titchamrsh [6] to also hold for $\sigma=\sigma_{t}\sim 0$ with an error term $\ll(\log T)^{-1}$.
+The reader might wonder if our approach can give a lower bound for $\Theta_{T}:$ the supremum of the real parts of the zeros $\rho_{T}$ with $T\leq\gamma_{T}\leq H.$ For that, one would need Theorem 4.11 of Titchamrsh [6] to also hold for $\sigma=\sigma_{t}=o(1)$ with an error term $\ll(\log T)^{-1}$.
 However, this is not possible (as can be seen from the approximate functional equation for $\zeta$).
-It can also be seen that Titchmarsh’s proof of his Theorem 4.11 doesn’t hold for $\sigma=\sigma_{t}\sim 0$.
+It can also be seen that Titchmarsh’s proof of his Theorem 4.11 doesn’t hold for $\sigma=\sigma_{t}=o(1)$.
 Indeed, he begins his argument by quoting the Euler-McLaurin summation formula for $\zeta(\sigma+it)$ (equation 3.5.3).
 This is initially defined for $\sigma>1$, and then is extended by holomorphic continuation to $\sigma>0$.
-Recall that a function can only be holomorphically extended on open sets, and that’s precisely why Titchmarsh needs the assumption that $\sigma\geq\sigma_{0}>0$.
+Recall that a function can only be holomorphically extended on open sets, and that’s precisely why Titchmarsh assumes that $\sigma$ is bounded below by some constant $\sigma_{0}>0$.

### v36

@@
 $\frac{\zeta^{\prime}}{\zeta}(-\sigma+it)\ll\log t$ (7)
-for $\sigma\geq\sigma_{0}>0.$ Combining (6) and (7) gives $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{3}\ll T^{-1/2}\log^{8}T.$ Thus we arrive at
-$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{7}T.$ (8)
+for $\sigma\geq\sigma_{0}>0.$ Combining (6) and (7) gives $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{3}\ll T^{1/2}\log^{8}T.$ Thus we arrive at
+$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{8}T.$ (8)
 For $\Re(s)>1$, note that $-\frac{\zeta^{\prime}}{\zeta}(s)=\sum_{m=1}^{\infty}\frac{\Lambda(m)}{m^{s}}$. Putting this into $I_{1}$ gives
 $I_{1}=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{e\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{e\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
-From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}<n<T$.
-For $m\geq T(1+\varepsilon)$ and $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll T^{1/4}$.
+From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}<n<T$ (unless specified otherwise).
+For $2\leq m\leq T(1-\varepsilon)$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll T^{1/4}$.
 Let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=\frac{m^{-\frac{1}{\log n}-1}(\log(m)\log(n)-\log^{2}(m)-\log^{2}n)}{n\log^{2}(n)\log^{2}(n/m)}$.
 Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) by routine partial summation that
-$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}$ $=S(T)g(T)-S(T/\log^{2}T)g(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)g^{\prime}(x)\mathrm{d}x$ (10)
+$\sum_{2\leq n\leq m(1-\varepsilon)}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}$ $=S(T)g(T)-S(T/\log^{2}T)g(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)g^{\prime}(x)\mathrm{d}x$ (10)
 $\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}-T^{1/2}\log^{3}T\int_{T/\log^{2}T}^{T}g^{\prime}(x)\mathrm{d}x$ (11)
 $\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (12)
-uniformly for $m\geq T(1+\varepsilon)$ and $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$.
+uniformly for $m\geq T(1+\varepsilon)$ and $2\leq m\leq T(1-\varepsilon)$.
 For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll T^{1/4}$ so it similarly follows from (5) by partial summation that
@@
 $\sum_{\frac{T}{\log^{2}T}<m(1+\varepsilon)\leq n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (14)
-We now treat the ranges $\frac{m}{1+\varepsilon}<n\leq m-1$ and $m+1\leq n<m(1+\varepsilon)$. Note that for $m\ll T$, we have
-$\sum_{\frac{m}{1+\varepsilon}<n\leq m-1}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\sum_{\frac{m}{1+\varepsilon}<n\leq m-1}\frac{1}{m|\log(n/m)|}\ll\log T$ (15)
-and
-$\sum_{m+1\leq n<m(1+\varepsilon)}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\sum_{m+1\leq n<m(1+\varepsilon)}\frac{1}{m\log(n/m)}\ll\log T.$ (16)
-Notice that (16) is obtained by writing $n=m+r$ for $11\leq r\leq\varepsilon m-$ 1 and applying the fact that $\log(1+\delta)\sim\delta$ uniformly for $1/m\leq\delta\leq\varepsilon$, hence
-$\sum_{m+1\leq n<m(1+\varepsilon)}\frac{1}{m\log(n/m)}\ll\sum_{1\leq r\leq\varepsilon m-1}\frac{1}{r}\ll\log T.$ (17)
-A similar argument yields (15).
-Summing (9) for $\frac{T}{\log^{2}T}<n<T$, changing the order of summation and invoking (16), (15), (14), (13) and (12) together with $\sum_{m\leq x}\Lambda(m)=x+O(x^{\Theta}\log^{2}x)$ [5, p.430], yields
+We now treat the ranges $\frac{m}{1+\varepsilon}<n\leq m-1$ and $m+1\leq n<m(1+\varepsilon)$. For $m\ll T$, note that
+$\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\log T$ (15)
+$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\log T$ (16)
+Notice that (16) is obtained by writing $m=n+r$ for $1\leq r\leq n\varepsilon$ and applying the fact that $\log(1+\delta)\sim\delta$ uniformly for $1/m\leq\delta\leq\varepsilon$, hence
+$\sum_{\frac{n}{1+\varepsilon}\leq m<n-1}\frac{1}{m\log(n/m)}\ll\sum_{1\leq r\leq\varepsilon m-1}\frac{1}{r}\ll\log T.$ (17)
+A similar argument yields (15). Define $\psi(x):=\sum_{m\leq x}\Lambda(m)$. Note that $\psi(x)=x+O(x^{\Theta}\log^{2}x)$ [5, p.430]. Thus from (16) it follows that
+$\sum_{T(1-\varepsilon)<m<T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}$ $\ll\sum_{T(1-\varepsilon)<m<T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)}{m^{1+\frac{1}{\log n}}\log(n/m)}$ (18)
+$\ll\log T\sum_{T(1-\varepsilon)<m<T(1+\varepsilon)}\Lambda(m)$ (19)
+$=(\log T)(\psi(T+T\varepsilon)-\psi(T-T\varepsilon)+O(\log T))$ (20)
+$=(\log T)(2\varepsilon T+O(T^{\Theta}\log^{2}T))$ (21)
+$=o(T).$ (22)
+By arguing similarly, one obtains from (15) that
+$\sum_{T(1-\varepsilon)<m<T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=o(T).$ (23)
+Summing (9) for $\frac{T}{\log^{2}T}<n<T$, changing the order of summation and invoking (23), (22), (14), (13) and (12) yields
 $\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+o(T)$ (24)
@@
 Doing so will definitely make one lose important information needed for a sufficiently sharp bound for the sums over $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$.
-(ii) The key parts of our argument, are the estimates in (12), (13) and (14), whose respective derivations are routine applications of partial summation from (5).
-It actually suffices for the reader to only verify this part.
+Note that the analogue of (26) from the Riemann explicit formula is $\sum_{T\leq\gamma_{T}\leq H}\frac{n^{\rho_{T}-1}}{\rho_{T}}\ll\frac{(\log T)(\log\log T)}{T}.$ This has no main term, so it’s not quite useful for our purposes.
+(ii) The key parts of our argument, are the estimates in (12), (13), (14), (22) and (23).
+The first three are are routine applications of partial summation from (5), whilst (22) and (23) follow easily from the well-known estimates for $\psi(x)$.
 Remark 4.
-The reader might wonder if our approach can give a lower bound for $\Theta_{T}:$ the supremum of the real parts of the zeros $\rho_{T}$ with $T\leq\gamma_{T}\leq H.$ For that, one would need Theorem 4.11 of Titchamrsh [6] to also hold for $\sigma=\sigma_{t}=o(1)$ with an error term $\ll(\log T)^{-1}$.
+The reader might wonder if our approach can give a lower bound for $\Theta_{T}:$ the supremum of the real parts of the zeros $\rho_{T}$ with $T\leq\gamma_{T}\leq H.$ For that, one would need Theorem 4.11 of Titchmarsh [6] to also hold for $\sigma=\sigma_{t}=o(1)$ with an error term $\ll(\log T)^{-1}$.
 However, this is not possible (as can be seen from the approximate functional equation for $\zeta$).

### v37

@@
 Definitions. Let: $\rho$ denote a non-trivial zero of $\zeta$, $\gamma=\Im(\rho)$ and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s.
+Lemma 1. (Theorem 4.11 of [6]). We have $\zeta(s)=\sum_{n\leq x}n^{-s}-\frac{x^{1-s}}{1-s}+O(x^{-\sigma})$ uniformly for $\sigma\geq\sigma_{0}>0,|\Im(s)|<2\pi x/C$, when $C$ is a given constant $>1$.
+Lemma 2, (Theorem 1 of [3]). One has $\sum_{T\leq\gamma_{T}\leq 2T+1}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ uniformly for $2\leq n\ll T.$
 Theorem 1. One has $\Theta=1$.
 Proof.
-Suppose that $\Theta<1$.
 Let: $T\in\mathbb{R}_{>0}$ be arbitrarily large and $X\geq 2$.
@@
 $\max\Big{(}\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iT)\Big{|},\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iH)\Big{|}\Big{)}\ll\log^{2}T$ (1)
-uniformly for $-1\leq\sigma\leq 2$. Let $n\geq\frac{T}{\log^{2}T}$. By the residue theorem, we have
-$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{1+\frac{1}{\log n}+iT}^{1+\frac{1}{\log n}+iH}+\int_{1+\frac{1}{\log n}+iH}^{-1+iH}+\int_{-1+iH}^{-1+iT}+\int_{-1+iT}^{1+\frac{1}{\log n}+iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
+uniformly for $-1\leq\sigma\leq 2$. By the residue theorem, we have
+$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{1+\frac{1}{\log n}+iT}^{1+\frac{1}{\log n}+\varepsilon+iH}+\int_{1+\frac{1}{\log n}+iH}^{-1+\varepsilon+iH}+\int_{-1+iH}^{-1+iT}+\int_{-1+iT}^{\Theta+\varepsilon+iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
 $=I_{1}+I_{2}+I_{3}+I_{4},$ (3)
-say. From now on, let $T\leq t\leq H$. By the approximate functional equation for $\zeta$ (Theorem 4.15 of [6]), note that
-$\zeta(\sigma+it)=\sum_{n\leq x}n^{-\sigma-it}+\chi(\sigma+it)\sum_{v\leq y}v^{\sigma-1+it}+O(x^{-\sigma}\log t+t^{\frac{1}{2}-\sigma}y^{\sigma-1})$ (4)
-where $0\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h$ is a constant and $|\chi(\sigma+it)|=\Big{|}\frac{\zeta(\sigma+it)}{\zeta(1-\sigma-it)}\Big{|}=\Big{|}\Big{(}\frac{2\pi}{t}\Big{)}^{\sigma-\frac{1}{2}+it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{1/2-\sigma}$ [6, p.78].
-Thus since $|\zeta(it)|\asymp t^{1/2}|\zeta(1-it)|\ll T^{1/2}\log T$, it follows from (4) that
-$S(x):=\sum_{n\leq x}n^{it}\ll T^{1/2}\log^{3}T$ (5)
-uniformly for $\frac{T}{\log^{2}T}\leq x\leq T$. Let $f(n)=n^{\sigma-1}$. Then by partial summation, we obtain from (5) that
-$\sum_{\frac{T}{\log^{2}T}\leq n\leq T}n^{\sigma-1+it}=S(T)f(T)-S(T/\log^{2}T)f(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)f^{\prime}(x)\mathrm{d}x\ll T^{\sigma-\frac{1}{2}}\log^{7}T$ (6)
-for $\sigma\in\mathbb{R},|\sigma|\leq 2$.
-Combining (1) and (6) yields $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{7}T.$ By the log-derivative of the functional equation for $\zeta$ (see e.g [3, p.403]), note that
+say. From now on, let $T\leq t\leq H$ and $\sigma=\Re(s)$. By the approximate functional equation for $\zeta$ (Theorem 4.15 of [6]), note that
+$\zeta(\sigma-it)=\sum_{n\leq x}n^{-\sigma+it}+\chi(\sigma-it)\sum_{v\leq y}v^{\sigma-1-it}+O(x^{-\sigma}+t^{\frac{1}{2}-\sigma}y^{\sigma-1})$ (4)
+where $00\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h>$ 0 is a constant and $|\chi(\sigma-it)|=\Big{|}\Big{(}\frac{t}{2\pi}\Big{)}^{\sigma-\frac{1}{2}-it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{\frac{1}{2}-\sigma}$ [6, p.78].
+Thus since $|\zeta(-it)|\asymp t^{1/2}|\zeta(1-it)|\ll T^{1/2}\log T$, it follows by taking $\sigma=0$ in (5) that
+$S(x):=\sum_{n\leq x}n^{it}\ll T^{\frac{1}{2}\log^{3}T}$ (5)
+uniformly for $\frac{T}{\log^{2}T}\leq x\leq T$. Let $f(n)=n^{\sigma-1}.$ Then by Abel summation, it follows that
+$\sum_{n\leq x}n^{\sigma-1+it}=S(T)f(T)-S(T/\log^{2}T)f(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)f^{\prime}(x)\mathrm{d}x\ll T^{\sigma-\frac{1}{2}}\log^{7}T$ (6)
+for $-1\leq\sigma\leq 2.$ Combining (1) and (6) gives $\sum_{\frac{T}{\log^{2}T}\leq n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{7}T.$ By the log-derivative of the functional equation for $\zeta$, we have
 $\frac{\zeta^{\prime}}{\zeta}(-\sigma+it)\ll\log T$ (7)
-for $\sigma\geq\sigma_{0}>0.$ Combining (6) and (7) gives $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{3}\ll T^{1/2}\log^{8}T.$ Thus we arrive at
-$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{8}T.$ (8)
-For $\Re(s)>1$, note that $-\frac{\zeta^{\prime}}{\zeta}(s)=\sum_{m=1}^{\infty}\frac{\Lambda(m)}{m^{s}}$. Putting this into $I_{1}$ gives
-$I_{1}=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{e\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{e\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
-From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}<n<T$ (unless specified otherwise).
-For $2\leq m\leq T(1-\varepsilon)$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll T^{1/4}$.
-Let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=\frac{m^{-\frac{1}{\log n}-1}(\log(m)\log(n)-\log^{2}(m)-\log^{2}n)}{n\log^{2}(n)\log^{2}(n/m)}$.
-Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) by routine partial summation that
-$\sum_{2\leq n\leq m(1-\varepsilon)}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}$ $=S(T)g(T)-S(T/\log^{2}T)g(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)g^{\prime}(x)\mathrm{d}x$ (10)
+for $\sigma\geq\sigma_{0}>0$. Combining (6) and (7) yields $\sum_{\frac{T}{\log^{2}T}\leq n\leq T}I_{3}\ll T^{1/2}\log^{7}T.$ Thus we arrive at
+$\sum_{\frac{T}{\log^{2}T}\leq n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{7}T.$ (8)
+Note that $\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m=1}^{\infty}\frac{\Lambda(m)}{m^{s}}$ for $\Re(s)=\sigma>1$. Inserting this into $I_{1}$ gives
+$I_{1}=\frac{(T-H))(\Lambda(n))}{2\pi n}+\frac{e}{2\pi i}\sum_{m=1,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
+From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$, $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise).
+For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{\log(n/m)}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$.
+Let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=-\frac{m^{-1-\frac{1}{\log n}}(\log(n)\log(m)-\log^{2}(n)-\log^{2}(m))}{n\log^{2}(n)\log^{2}(n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
+$\sum_{\frac{T}{\log^{2}T}\leq n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}$ $=S(T)g(T)-S(T/\log^{2}T)g(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)g^{\prime}(x)\mathrm{d}x$ (10)
 $\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}-T^{1/2}\log^{3}T\int_{T/\log^{2}T}^{T}g^{\prime}(x)\mathrm{d}x$ (11)
 $\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (12)
-uniformly for $m\geq T(1+\varepsilon)$ and $2\leq m\leq T(1-\varepsilon)$.
-For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll T^{1/4}$ so it similarly follows from (5) by partial summation that
-$\sum_{\frac{T}{\log^{2}T}<n\leq\frac{m}{1+\varepsilon}<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (13)
+uniformly for $1\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon).$ For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll 1/\varepsilon=T^{1/4}$ so it similarly follows from (5) by Abel summation that
+$\sum_{\frac{T}{\log^{2}T}\leq n\leq\frac{m}{1+\varepsilon}\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (13)
 and
-$\sum_{\frac{T}{\log^{2}T}<m(1+\varepsilon)\leq n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (14)
-We now treat the ranges $\frac{m}{1+\varepsilon}<n\leq m-1$ and $m+1\leq n<m(1+\varepsilon)$. For $m\ll T$, note that
-$\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\log T$ (15)
-$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\log T$ (16)
-Notice that (16) is obtained by writing $m=n+r$ for $1\leq r\leq n\varepsilon$ and applying the fact that $\log(1+\delta)\sim\delta$ uniformly for $1/m\leq\delta\leq\varepsilon$, hence
-$\sum_{\frac{n}{1+\varepsilon}\leq m<n-1}\frac{1}{m\log(n/m)}\ll\sum_{1\leq r\leq\varepsilon m-1}\frac{1}{r}\ll\log T.$ (17)
-A similar argument yields (15). Define $\psi(x):=\sum_{m\leq x}\Lambda(m)$. Note that $\psi(x)=x+O(x^{\Theta}\log^{2}x)$ [5, p.430]. Thus from (16) it follows that
-$\sum_{T(1-\varepsilon)<m<T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}$ $\ll\sum_{T(1-\varepsilon)<m<T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)}{m^{1+\frac{1}{\log n}}\log(n/m)}$ (18)
-$\ll\log T\sum_{T(1-\varepsilon)<m<T(1+\varepsilon)}\Lambda(m)$ (19)
-$=(\log T)(\psi(T+T\varepsilon)-\psi(T-T\varepsilon)+O(\log T))$ (20)
-$=(\log T)(2\varepsilon T+O(T^{\Theta}\log^{2}T))$ (21)
-$=o(T).$ (22)
-By arguing similarly, one obtains from (15) that
-$\sum_{T(1-\varepsilon)<m<T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=o(T).$ (23)
-Summing (9) for $\frac{T}{\log^{2}T}<n<T$, changing the order of summation and invoking (23), (22), (14), (13) and (12) yields
-$\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+o(T)$ (24)
+$\sum_{\frac{T}{\log^{2}T}\leq m(1+\varepsilon)\leq n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (14)
+We now treat the ranges $m+1\leq n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq n\leq m-1$.
+For $m+1\leq n\leq m(1+\varepsilon)$, put $n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(n/m)=r/m(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=m/r(1+O(r/m))$ hence
+$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}$ $=\frac{\Lambda(m)}{m^{it}}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}$ (15)
+$=\frac{\Lambda(m)}{em^{it}}\sum_{0<r\leq\varepsilon m-1}\Big{(}\frac{(m+r)^{it}}{r}+O(r/m^{2})\Big{)}$ (16)
+$=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq\varepsilon m-1}\frac{(1+\frac{r}{m})^{it}}{r}+O(\varepsilon^{2}\log m).$ (17)
+If $f(x)$ is a real function that is differentiable thrice and $f^{\prime}(x)$ is steadily decreasing on $[a,b]$, note that
+$\sum_{a<n\leq b}g(n)e^{2\pi if(n)}\sim e^{-\frac{1}{4}\pi i}\sum_{\alpha<v\leq\beta}\frac{e^{2\pi i(f(x_{v})-vx_{v})}}{|f^{\prime\prime}(x_{v})|^{\frac{1}{2}}}g(x_{v})\ll\sum_{\alpha<v\leq\beta}\Big{|}\frac{g(x_{v})}{f^{\prime\prime}(x_{v})^{\frac{1}{2}}}\Big{|}$ (18)
+where $f^{\prime}(x_{v})=v,\alpha=f^{\prime}(a)$ and $\beta=f^{\prime}(b)$ [6, p.78].
+Taking $g(u)=1/u,f(u)=\frac{t\log(1+\frac{u}{m})}{2\pi},f^{\prime}(u)=\frac{t}{2\pi(m+u)},f^{\prime\prime}(u)=-\frac{t}{2\pi(m+u)^{2}},x_{v}=\frac{t}{2\pi v}-m,f^{\prime\prime}(x_{v})=\frac{v^{2}}{t}$ gives, for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$,
+$\sum_{0<r\leq\varepsilon m-1}\frac{(1+\frac{r}{m})^{it}}{r}=\sum_{0<r\leq\varepsilon m-1}\frac{e^{it\log(1+\frac{r}{m})}}{r}\ll\sum_{\alpha<v\leq\frac{t}{2\pi m(1+\varepsilon)}}\frac{\sqrt{t}}{\frac{t}{2\pi}-mv}\ll\sum_{\alpha<v\leq\log^{2}t}\frac{1}{\varepsilon\sqrt{t}}\ll T^{-1/4}\log^{2}T.$ (19)
+Inserting (19) into the right-hand side of (17), summing over $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$ and applying the Prime Number Theorem that $\sum_{m\leq x}\Lambda(m)\sim x,$ yields
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll T^{3/4}\log^{2}T.$ (20)
+For $\frac{m}{1+\varepsilon}\leq n\leq m-1$, put $n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ thus $\log(n/m)=-r/m(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=-m/r(1+O(r/m))$.
+Thus by arguing as above, one arrives at
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll T^{3/4}\log^{2}T.$ (21)
+Summing (9) for $\frac{T}{\log^{2}T}<n<T$, changing the order of summation, invoking (21), (20), (14), (13) and (12) yields
+$\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi n}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+o(T)$ (22)
 $=-\frac{T}{\pi}\log\log T+o(T)$ (23)
-since $H=2T+O(1)$ and $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant [5, p.182]. By Theorem 1 of [3], we have
+since $H=2T+O(1)$ and $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant [5, p.182]. By Lemma 2, we have
 $\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ (24)
@@
 Combining (26), (23), (8) and (3) yields
-$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{2}T}}+\sum_{\frac{T}{\log^{2}T}<n\leq T}\Big{)}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ (29)
+$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{2}T}}+\sum_{\frac{T}{\log^{2}T}<n<T}\Big{)}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ (27)
 $\gg T.$ (28)
-By Theorem 4.11 of [6], we have
-$\sum_{1\leq n\leq T}n^{\rho_{T}-1}=\zeta(1-\rho_{T})+\frac{T^{\rho_{T}}}{\rho_{T}}+O(T^{\Theta-1})\ll T^{\Theta-1}$ (31)
-uniformly for $T\leq\gamma_{T}\leq H.$ Thus
-$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}$ (32)
-$\ll T^{\Theta}\log T.$ (33)
+Suppose that $\Theta<1$. Then by Lemma 1 we have
+$\sum_{1\leq n\leq T}n^{\rho_{T}-1}\ll T^{\Theta-1}$ (29)
+uniformly for $T\leq\gamma_{T}\leq H.$ Hence
+$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}\ll T^{\Theta}\log T.$ (30)
 But (30) is a contradiction to (28), thus our supposition must be false. This completes the proof. ∎
@@
 Remark 3.
-(i) From the literature on arXiv and elsewhere, it doesn’t appear like this approach on RH has been tried before.
+From the literature on arXiv and elsewhere, it doesn’t appear like this approach on RH has been tried before.
 However, the idea of applying Gonek’s formula (24) in estimating sums over zeros of zeta in general, is not new.
@@
 Note that the analogue of (21) from the Riemann explicit formula is $\sum_{T\leq\gamma_{T}\leq H}\frac{n^{\rho_{T}-1}}{\rho_{T}}\ll\frac{(\log T)(\log\log T)}{T}.$ This has no main term, so it’s not quite useful for our purposes.
-(ii) The key parts of our argument, are the estimates in (12), (13), (14), (22) and (23).
-The first three are are routine applications of partial summation from (5), whilst (22) and (23) follow easily from the well-known estimates for $\psi(x)$.
 Remark 4.

### v38

@@
 $\zeta(\sigma-it)=\sum_{n\leq x}n^{-\sigma+it}+\chi(\sigma-it)\sum_{v\leq y}v^{\sigma-1-it}+O(x^{-\sigma}+t^{\frac{1}{2}-\sigma}y^{\sigma-1})$ (4)
-where $00\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h>$ 0 is a constant and $|\chi(\sigma-it)|=\Big{|}\Big{(}\frac{t}{2\pi}\Big{)}^{\sigma-\frac{1}{2}-it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{\frac{1}{2}-\sigma}$ [6, p.78].
+where $00\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h>$ 0 is a constant and $|\chi(\sigma-it)|=\Big{|}\Big{(}\frac{t}{2\pi}\Big{)}^{\sigma-\frac{1}{2}-it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{\frac{1}{2}-\sigma}$ [7, p.78].
 Thus since $|\zeta(-it)|\asymp t^{1/2}|\zeta(1-it)|\ll T^{1/2}\log T$, it follows by taking $\sigma=0$ in (4) that
@@
 Note that $\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m=1}^{\infty}\frac{\Lambda(m)}{m^{s}}$ for $\Re(s)=\sigma>1$. Inserting this into $I_{1}$ gives
-$I_{1}=\frac{(T-H))(\Lambda(n))}{2\pi n}+\frac{e}{2\pi i}\sum_{m=1,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
-From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$, $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise).
+$I_{1}=\frac{(T-H))\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=1,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
+From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise).
 For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{\log(n/m)}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$.
@@
 $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq\varepsilon m-1}\frac{(1+\frac{r}{m})^{it}}{r}+O(\varepsilon^{2}\log m).$ (17)
-If $f(x)$ is a real function that is differentiable thrice and $f^{\prime}(x)$ is steadily decreasing on $[a,b]$, note that
-$\sum_{a<n\leq b}g(n)e^{2\pi if(n)}\sim e^{-\frac{1}{4}\pi i}\sum_{\alpha<v\leq\beta}\frac{e^{2\pi i(f(x_{v})-vx_{v})}}{|f^{\prime\prime}(x_{v})|^{\frac{1}{2}}}g(x_{v})\ll\sum_{\alpha<v\leq\beta}\Big{|}\frac{g(x_{v})}{f^{\prime\prime}(x_{v})^{\frac{1}{2}}}\Big{|}$ (18)
-where $f^{\prime}(x_{v})=v,\alpha=f^{\prime}(a)$ and $\beta=f^{\prime}(b)$ [6, p.78].
-Taking $g(u)=1/u,f(u)=\frac{t\log(1+\frac{u}{m})}{2\pi},f^{\prime}(u)=\frac{t}{2\pi(m+u)},f^{\prime\prime}(u)=-\frac{t}{2\pi(m+u)^{2}},x_{v}=\frac{t}{2\pi v}-m,f^{\prime\prime}(x_{v})=\frac{v^{2}}{t}$ gives, for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$,
-$\sum_{0<r\leq\varepsilon m-1}\frac{(1+\frac{r}{m})^{it}}{r}=\sum_{0<r\leq\varepsilon m-1}\frac{e^{it\log(1+\frac{r}{m})}}{r}\ll\sum_{\alpha<v\leq\frac{t}{2\pi m(1+\varepsilon)}}\frac{\sqrt{t}}{\frac{t}{2\pi}-mv}\ll\sum_{\alpha<v\leq\log^{2}t}\frac{1}{\varepsilon\sqrt{t}}\ll T^{-1/4}\log^{2}T.$ (19)
+Note that if $f(t)$ is differentiable on $[a,b]$, $f^{\prime}(t)$ is monotonic decreasing on $[a,b]$, $f^{\prime}(a)=\beta,f^{\prime}(b)=\alpha,$ then
+$\sum_{a<r\leq b}e^{2\pi if(r)}=\sum_{|v|\leq N}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x+o(1)$ (18)
+as $N\rightarrow\infty$, where
+$\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x\ll\Big{|}\frac{1}{\alpha-v}-\frac{1}{\beta-v}\Big{|}=\frac{\beta-\alpha}{|(\alpha-v)(\beta-v)|}$ (19)
+see e.g. [6, p.10]. Taking $f(u)=\frac{t\log(1+\frac{u}{m})}{2\pi},a=0,b=m\varepsilon-1$ and $f^{\prime}(u)=\frac{t}{2\pi(m+u)}$ gives
+$\sum_{-N\leq v\leq 0}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $\ll\sum_{-N\leq v\leq 0}\frac{\beta-\alpha}{(\alpha-v)(\beta-v)}$ (20)
+$=(\beta-\alpha)\sum_{0\leq v\leq N}\frac{1}{(\alpha+v)(\beta+v)}$ (21)
+$\ll(\beta-\alpha)\sum_{0\leq v\leq N}\frac{1}{1+v^{2}}$ (22)
+$\ll\beta-\alpha$ (23)
+$\sim\frac{t\varepsilon}{2\pi m}.$ (24)
+Choose a rational $t\in[T,H]$ so that $\alpha=\frac{t}{2\pi(m+M)}$ and $\beta=\frac{t}{2\pi m}$ are never integers, where $m\in\Big{[}\frac{T}{(1+\varepsilon)\log^{2}T},T(1+\varepsilon)\Big{]}$ is an integer and $M$ is the largest integer $\leq m\varepsilon-1$.
+Thus $v\neq\alpha$ in particular.
+Let $K_{0}$ be the largest integer $<\alpha$.
+We also have
+$\sum_{1\leq v\leq\alpha}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x=\sum_{1\leq v\leq K_{0}}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $=\sum_{1\leq v\leq K_{0}}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ (25)
+$\ll\sum_{1\leq v\leq K_{0}}\frac{\beta-\alpha}{(\alpha-v)(\beta-v)}$ (26)
+$\ll(\beta-\alpha)\int_{1}^{K_{0}}\frac{1}{(\alpha-v)(\beta-v)}\mathrm{d}v$ (27)
+$\ll(\beta-\alpha)\log t$ (28)
+$\ll(\beta-\alpha)\log t$ (29)
+$\sim\frac{t\varepsilon}{2\pi m}\log t.$ (30)
+Let $K_{1}$ be the least integer $>\beta$. Note that
+$\sum_{\beta<v\leq N}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $\ll\sum_{K_{1}\leq v\leq N}\frac{\beta-\alpha}{(v-\alpha)(v-\beta)}$ (31)
+$\ll(\beta-\alpha)\int_{K_{1}}^{N}\frac{1}{(v-\alpha)(v-\beta)}\mathrm{d}v$ (32)
+$\ll(\beta-\alpha)\log t$ (33)
+$\ll(\beta-\alpha)\log t$ (34)
+$\sim\frac{t\varepsilon}{2\pi m}\log t.$ (35)
+Since $\beta-\alpha\sim\frac{t\varepsilon}{2\pi m}\ll\frac{\log^{2}T}{T^{1/4}}=o(1)$, note that the range $\alpha<v\leq\beta$ contains no integers. Thus
+$\sum_{\alpha<v\leq\beta}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x=0.$ (36)
+Combining (36), (35), (30), (25) and (19), yields $\sum_{a<r\leq b}e^{2\pi if(r)}=o(1)$ and by Abel summation it follows that
+$\sum_{a<r\leq b}\frac{1}{r}e^{2\pi if(r)}=\sum_{0<r\leq\varepsilon m-1}\frac{1}{r}e^{it\log(1+\frac{r}{m})}=o(1).$ (37)
 Inserting (37) into the right-hand side of (18), summing over $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$ and applying the Prime Number Theorem that $\sum_{m\leq x}\Lambda(m)\sim x,$ yields
-$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll T^{3/4}\log^{2}T.$ (20)
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=o(T).$ (38)
 For $\frac{m}{1+\varepsilon}\leq n\leq m-1$, put $n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ thus $\log(n/m)=-r/m(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=-m/r(1+O(r/m))$.
 Thus by arguing as above, one arrives at
-$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll T^{3/4}\log^{2}T.$ (21)
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=o(T).$ (39)
 Summing (9) for $\frac{T}{\log^{2}T}<n<T$, changing the order of summation, invoking (39), (38), (14), (13) and (12) yields
-$\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi n}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+o(T)$ (22)
+$\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+o(T)$ (40)
 $=-\frac{T}{\pi}\log\log T+o(T)$ (41)

### v39

@@
 For a far more thorough discussion of the RH and GRH, the interested reader is kindly referred to Borwein et al [1].
-Definitions. Let: $\rho$ denote a non-trivial zero of $\zeta$, $\gamma=\Im(\rho)$ and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s.
-Lemma 1. (Theorem 4.11 of [7]). We have $\zeta(s)=\sum_{n\leq x}n^{-s}-\frac{x^{1-s}}{1-s}+O(x^{-\sigma})$ uniformly for $\sigma\geq\sigma_{0}>0,|\Im(s)|<2\pi x/C$, when $C$ is a given constant $>1$.
-Lemma 2, (Theorem 1 of [3]). One has $\sum_{T\leq\gamma_{T}\leq 2T+1}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ uniformly for $2\leq n\ll T.$
+Definitions. Let: $\rho$ denote a non-trivial zero of $\zeta,\sigma=\Re(s)$, $\gamma=\Im(\rho)$ and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s.
+Lemma 1.
+(Theorem 4.11 of [7]).
+We have $\zeta(s)=\sum_{n\leq x}n^{-s}-\frac{x^{1-s}}{1-s}+O(x^{-\sigma})$ uniformly for $\sigma\geq\sigma_{0}>0$ and $|\Im(s)|<2\pi x/C$, where $C$ is a given constant $>1$.
+Lemma 2, (Theorem 1 of [3]). One has $\sum_{0\leq\gamma\leq y}n^{\rho-1}=-\frac{y\Lambda(n)}{2\pi n}+O((\log y)\log\log 3n)$ uniformly for $2\leq n\ll y.$
 Theorem 1. One has $\Theta=1$.
@@
 uniformly for $-1\leq\sigma\leq 2$. By the residue theorem, we have
-$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{1+\frac{1}{\log n}+iT}^{1+\frac{1}{\log n}+\varepsilon+iH}+\int_{1+\frac{1}{\log n}+iH}^{-1+\varepsilon+iH}+\int_{-1+iH}^{-1+iT}+\int_{-1+iT}^{\Theta+\varepsilon+iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
+$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{1+\frac{1}{\log n}+iT}^{1+\frac{1}{\log n}+iH}+\int_{1+\frac{1}{\log n}+iH}^{-1+iH}+\int_{-1+iH}^{-1+iT}+\int_{-1+iT}^{1+\frac{1}{\log n}+iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
 $=I_{1}+I_{2}+I_{3}+I_{4},$ (3)
-say. From now on, let $T\leq t\leq H$ and $\sigma=\Re(s)$. By the approximate functional equation for $\zeta$ (Theorem 4.15 of [7]), note that
+say. From now on, let $T\leq t\leq H$. By the approximate functional equation for $\zeta$ (Theorem 4.15 of [7]), note that
 $\zeta(\sigma-it)=\sum_{n\leq x}n^{-\sigma+it}+\chi(\sigma-it)\sum_{v\leq y}v^{\sigma-1-it}+O(x^{-\sigma}+t^{\frac{1}{2}-\sigma}y^{\sigma-1})$ (4)
@@
 $\sum_{\frac{T}{\log^{2}T}\leq n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{7}T.$ (8)
-Note that $\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m=1}^{\infty}\frac{\Lambda(m)}{m^{s}}$ for $\Re(s)=\sigma>1$. Inserting this into $I_{1}$ gives
-$I_{1}=\frac{(T-H))\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=1,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
+Note that $\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m=1}^{\infty}\frac{\Lambda(m)}{m^{s}}$ for $\sigma>1$. Inserting this into $I_{1}$ gives
+$I_{1}=\frac{(T-H)\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=1,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
 From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise).
-For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{\log(n/m)}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$.
+For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$.
 Let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=-\frac{m^{-1-\frac{1}{\log n}}(\log(n)\log(m)-\log^{2}(n)-\log^{2}(m))}{n\log^{2}(n)\log^{2}(n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
@@
 uniformly for $1\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon).$ For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll 1/\varepsilon=T^{1/4}$ so it similarly follows from (5) by Abel summation that
-$\sum_{\frac{T}{\log^{2}T}\leq n\leq\frac{m}{1+\varepsilon}\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (13)
+$\sum_{\frac{T}{\log^{2}T}\leq n\leq\frac{m}{1+\varepsilon}\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (13)
 and
@@
 We now treat the ranges $m+1\leq n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq n\leq m-1$.
-For $m+1\leq n\leq m(1+\varepsilon)$, put $n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(n/m)=r/m(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=m/r(1+O(r/m))$ hence
+For $m+1\leq n\leq m(1+\varepsilon)$, put $n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(n/m)=\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=\frac{m}{r}(1+O(r/m))$ hence
 $\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}$ $=\frac{\Lambda(m)}{m^{it}}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}$ (15)
@@
 $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq\varepsilon m-1}\frac{(1+\frac{r}{m})^{it}}{r}+O(\varepsilon^{2}\log m).$ (17)
-Note that if $f(t)$ is differentiable on $[a,b]$, $f^{\prime}(t)$ is monotonic decreasing on $[a,b]$, $f^{\prime}(a)=\beta,f^{\prime}(b)=\alpha,$ then
+Note that if $f(t)$ is a real differentiable function on $[a,b]$, $f^{\prime}(t)$ is monotonic decreasing on $[a,b]$, $f^{\prime}(a)=\beta$ and $f^{\prime}(b)=\alpha,$ then
 $\sum_{a<r\leq b}e^{2\pi if(r)}=\sum_{|v|\leq N}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x+o(1)$ (18)
@@
 $\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x\ll\Big{|}\frac{1}{\alpha-v}-\frac{1}{\beta-v}\Big{|}=\frac{\beta-\alpha}{|(\alpha-v)(\beta-v)|}$ (19)
-see e.g. [6, p.10]. Taking $f(u)=\frac{t\log(1+\frac{u}{m})}{2\pi},a=0,b=m\varepsilon-1$ and $f^{\prime}(u)=\frac{t}{2\pi(m+u)}$ gives
+see e.g. [6, pp.10-11]. Taking $f(r)=\frac{t\log(1+\frac{r}{m})}{2\pi},a=0,b=m\varepsilon-1$ and $f^{\prime}(r)=\frac{t}{2\pi(m+r)}$ gives
 $\sum_{-N\leq v\leq 0}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $\ll\sum_{-N\leq v\leq 0}\frac{\beta-\alpha}{(\alpha-v)(\beta-v)}$ (20)
@@
 We also have
-$\sum_{1\leq v\leq\alpha}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x=\sum_{1\leq v\leq K_{0}}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $=\sum_{1\leq v\leq K_{0}}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ (25)
+$\sum_{1\leq v\leq\alpha}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $=\sum_{1\leq v\leq K_{0}}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ (25)
 $\ll\sum_{1\leq v\leq K_{0}}\frac{\beta-\alpha}{(\alpha-v)(\beta-v)}$ (26)
 $\ll(\beta-\alpha)\int_{1}^{K_{0}}\frac{1}{(\alpha-v)(\beta-v)}\mathrm{d}v$ (27)
-$\ll(\beta-\alpha)\log t$ (28)
 $\ll(\beta-\alpha)\log t$ (28)
@@
 $\ll(\beta-\alpha)\log t$ (32)
-$\ll(\beta-\alpha)\log t$ (34)
 $\sim\frac{t\varepsilon}{2\pi m}\log t.$ (33)
@@
 $\sum_{\alpha<v\leq\beta}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x=0.$ (34)
-Combining (36), (35), (30), (25) and (19), yields $\sum_{a<r\leq b}e^{2\pi if(r)}=o(1)$ and by Abel summation it follows that
+Combining (34), (33), (29), (24) and (18) yields $\sum_{a<r\leq b}e^{2\pi if(r)}=o(1)$. Thus by Abel summation it follows that
 $\sum_{a<r\leq b}\frac{1}{r}e^{2\pi if(r)}=\sum_{0<r\leq\varepsilon m-1}\frac{1}{r}e^{it\log(1+\frac{r}{m})}=o(1).$ (35)
@@
 $\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=o(T).$ (36)
-For $\frac{m}{1+\varepsilon}\leq n\leq m-1$, put $n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ thus $\log(n/m)=-r/m(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=-m/r(1+O(r/m))$.
+For $\frac{m}{1+\varepsilon}\leq n\leq m-1$, put $n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ thus $\log(n/m)=-\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=-\frac{r}{m}(1+O(r/m))$.
 Thus by arguing as above, one arrives at
@@
 uniformly for $T\leq\gamma_{T}\leq H.$ Hence
-$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}\ll T^{\Theta}\log T.$ (48)
+$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}\ll T^{\Theta}\log T.$ (46)
 But (46) is a contradiction to (44), thus our supposition must be false. This completes the proof. ∎

### v40

@@
 $\zeta(\sigma-it)=\sum_{n\leq x}n^{-\sigma+it}+\chi(\sigma-it)\sum_{v\leq y}v^{\sigma-1-it}+O(x^{-\sigma}+t^{\frac{1}{2}-\sigma}y^{\sigma-1})$ (4)
-where $00\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h>$ 0 is a constant and $|\chi(\sigma-it)|=\Big{|}\Big{(}\frac{t}{2\pi}\Big{)}^{\sigma-\frac{1}{2}-it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{\frac{1}{2}-\sigma}$ [7, p.78].
+where $00\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h>$ 0 is a constant and $|\chi(\sigma-it)|=|\chi(\sigma+it)|=\Big{|}\Big{(}\frac{t}{2\pi}\Big{)}^{\sigma-\frac{1}{2}-it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{\frac{1}{2}-\sigma}$ [7, p.78].
 Thus since $|\zeta(-it)|\asymp t^{1/2}|\zeta(1-it)|\ll T^{1/2}\log T$, it follows by taking $\sigma=0$ in (4) that
@@
 $\sum_{\frac{T}{\log^{2}T}\leq n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{7}T.$ (8)
-Note that $\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m=1}^{\infty}\frac{\Lambda(m)}{m^{s}}$ for $\sigma>1$. Inserting this into $I_{1}$ gives
-$I_{1}=\frac{(T-H)\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=1,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
+Note that $\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m=2}^{\infty}\frac{\Lambda(m)}{m^{s}}$ for $\sigma>1$. Inserting this into $I_{1}$ gives
+$I_{1}=\frac{(T-H)\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
 From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise).
 For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$.
-Let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=-\frac{m^{-1-\frac{1}{\log n}}(\log(n)\log(m)-\log^{2}(n)-\log^{2}(m))}{n\log^{2}(n)\log^{2}(n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
+For $m\neq n$, let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=\frac{m^{-1-\frac{1}{\log n}}(\log(n)\log(m)-\log^{2}(n)-\log^{2}(m))}{n\log^{2}(n)\log^{2}(n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
 $\sum_{\frac{T}{\log^{2}T}\leq n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}$ $=S(T)g(T)-S(T/\log^{2}T)g(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)g^{\prime}(x)\mathrm{d}x$ (10)
@@
 $\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (12)
-uniformly for $1\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon).$ For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll 1/\varepsilon=T^{1/4}$ so it similarly follows from (5) by Abel summation that
+uniformly for $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon).$ For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll 1/\varepsilon=T^{1/4}$ so it similarly follows from (5) by Abel summation that
 $\sum_{\frac{T}{\log^{2}T}\leq n\leq\frac{m}{1+\varepsilon}\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (13)
@@
 We now treat the ranges $m+1\leq n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq n\leq m-1$.
-For $m+1\leq n\leq m(1+\varepsilon)$, put $n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(n/m)=\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=\frac{m}{r}(1+O(r/m))$ hence
+For $m+1\leq n\leq m(1+\varepsilon)$, put $n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(n/m)=\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=\frac{m}{r}(1+O(r/m))=\frac{m}{r}+O(1)$ hence
 $\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}$ $=\frac{\Lambda(m)}{m^{it}}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}$ (15)
-$=\frac{\Lambda(m)}{em^{it}}\sum_{0<r\leq\varepsilon m-1}\Big{(}\frac{(m+r)^{it}}{r}+O(r/m^{2})\Big{)}$ (16)
-$=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq\varepsilon m-1}\frac{(1+\frac{r}{m})^{it}}{r}+O(\varepsilon^{2}\log m).$ (17)
+$=\frac{\Lambda(m)}{em^{it}}\sum_{0<r\leq\varepsilon m-1}\Big{(}\frac{(m+r)^{it}}{r}+O(1/m)\Big{)}$ (16)
+$=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq\varepsilon m-1}\frac{(1+\frac{r}{m})^{it}}{r}+O(\varepsilon\log m).$ (17)
 Note that if $f(t)$ is a real differentiable function on $[a,b]$, $f^{\prime}(t)$ is monotonic decreasing on $[a,b]$, $f^{\prime}(a)=\beta$ and $f^{\prime}(b)=\alpha,$ then
@@
 $\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=o(T).$ (36)
-For $\frac{m}{1+\varepsilon}\leq n\leq m-1$, put $n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ thus $\log(n/m)=-\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=-\frac{r}{m}(1+O(r/m))$.
+For $\frac{m}{1+\varepsilon}\leq n\leq m-1$, put $n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ thus $\log(n/m)=-\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=-\frac{m}{r}(1+O(r/m))=-\frac{m}{r}+O(1)$.
 Thus by arguing as above, one arrives at

### v41

@@
 Definitions. Let: $\rho$ denote a non-trivial zero of $\zeta,\sigma=\Re(s)$, $\gamma=\Im(\rho)$ and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s.
+Definitions. Let: $\rho$ denote a non-trivial zero of $\zeta,\sigma=\Re(s)$, $\gamma=\Im(\rho)$ and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s.
 Lemma 1.
@@
 $\zeta(\sigma-it)=\sum_{n\leq x}n^{-\sigma+it}+\chi(\sigma-it)\sum_{v\leq y}v^{\sigma-1-it}+O(x^{-\sigma}+t^{\frac{1}{2}-\sigma}y^{\sigma-1})$ (4)
-where $00\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h>$ 0 is a constant and $|\chi(\sigma-it)|=|\chi(\sigma+it)|=\Big{|}\Big{(}\frac{t}{2\pi}\Big{)}^{\sigma-\frac{1}{2}-it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{\frac{1}{2}-\sigma}$ [7, p.78].
-Thus since $|\zeta(-it)|\asymp t^{1/2}|\zeta(1-it)|\ll T^{1/2}\log T$, it follows by taking $\sigma=0$ in (4) that
+where $00\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h>$ 0 is a constant, $\chi(s)=\frac{\zeta(s)}{\zeta(1-s)}$ and $|\chi(\sigma-it)|=|\chi(\sigma+it)|=\Big{|}\Big{(}\frac{t}{2\pi}\Big{)}^{\sigma-\frac{1}{2}-it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{\frac{1}{2}-\sigma}$ [7, p.78].
+Thus since $|\zeta(-it)|\asymp t^{1/2}|\zeta(1+it)|\ll T^{1/2}\log T$, it follows by taking $\sigma=0$ in (4) that
 $S(x):=\sum_{n\leq x}n^{it}\ll T^{\frac{1}{2}}\log^{3}T$ (5)
@@
 $\sum_{\alpha<v\leq\beta}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x=0.$ (34)
-Combining (34), (33), (29), (24) and (18) yields $\sum_{a<r\leq b}e^{2\pi if(r)}=o(1)$. Thus by Abel summation it follows that
-$\sum_{a<r\leq b}\frac{1}{r}e^{2\pi if(r)}=\sum_{0<r\leq\varepsilon m-1}\frac{1}{r}e^{it\log(1+\frac{r}{m})}=o(1).$ (35)
+Combining (34), (33), (29), (24) and (18) yields $\sum_{a<r\leq b}e^{2\pi if(r)}=o(1)$.
+Infact, by adapting the argument between (18) and (34), one finds that $\sum_{0<r\leq J}e^{2\pi if(r)}=o(1)$ uniformly for $11\leq J\leq b=m\varepsilon-$ 1.
+Thus by Abel summation it follows that
+$\sum_{a<r\leq b}\frac{1}{r}e^{2\pi if(r)}=\sum_{0<r\leq b}\frac{1}{r}e^{it\log(1+\frac{r}{m})}=o(1).$ (35)
 Inserting (35) into the right-hand side of (17), summing over $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$ and applying the Prime Number Theorem that $\sum_{m\leq x}\Lambda(m)\sim x,$ yields

### v42

@@
 For a far more thorough discussion of the RH and GRH, the interested reader is kindly referred to Borwein et al [1].
-Definitions. Let: $\rho$ denote a non-trivial zero of $\zeta,\sigma=\Re(s)$, $\gamma=\Im(\rho)$ and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s.
 Definitions. Let: $\rho$ denote a non-trivial zero of $\zeta,\sigma=\Re(s)$, $\gamma=\Im(\rho)$ and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s.
@@
 Note that $\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m=2}^{\infty}\frac{\Lambda(m)}{m^{s}}$ for $\sigma>1$. Inserting this into $I_{1}$ gives
-$I_{1}=\frac{(T-H)\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
+$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}$ (9)
+say.
 From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise).
@@
 $\sum_{\frac{T}{\log^{2}T}\leq m(1+\varepsilon)\leq n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (14)
-We now treat the ranges $m+1\leq n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq n\leq m-1$.
+We now treat the ranges $m+1\leq n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq n\leq m-1$, where $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$.
 For $m+1\leq n\leq m(1+\varepsilon)$, put $n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(n/m)=\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=\frac{m}{r}(1+O(r/m))=\frac{m}{r}+O(1)$ hence
@@
 $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq\varepsilon m-1}\frac{(1+\frac{r}{m})^{iT}}{r}+O(\varepsilon\log m).$ (17)
-Note that if $f(t)$ is a real differentiable function on $[a,b]$, $f^{\prime}(t)$ is monotonic decreasing on $[a,b]$, $f^{\prime}(a)=\beta$ and $f^{\prime}(b)=\alpha,$ then
-$\sum_{a<r\leq b}e^{2\pi if(r)}=\sum_{|v|\leq N}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x+o(1)$ (18)
-as $N\rightarrow\infty$, where
-$\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x\ll\Big{|}\frac{1}{\alpha-v}-\frac{1}{\beta-v}\Big{|}=\frac{\beta-\alpha}{|(\alpha-v)(\beta-v)|}$ (19)
-see e.g. [6, pp.10-11]. Taking $f(r)=\frac{t\log(1+\frac{r}{m})}{2\pi},a=0,b=m\varepsilon-1$ and $f^{\prime}(r)=\frac{t}{2\pi(m+r)}$ gives
-$\sum_{-N\leq v\leq 0}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $\ll\sum_{-N\leq v\leq 0}\frac{\beta-\alpha}{(\alpha-v)(\beta-v)}$ (20)
-$=(\beta-\alpha)\sum_{0\leq v\leq N}\frac{1}{(\alpha+v)(\beta+v)}$ (21)
-$\ll(\beta-\alpha)\sum_{0\leq v\leq N}\frac{1}{1+v^{2}}$ (22)
-$\ll\beta-\alpha$ (23)
-$\sim\frac{t\varepsilon}{2\pi m}.$ (24)
-Choose a rational $t\in[T,H]$ so that $\alpha=\frac{t}{2\pi(m+M)}$ and $\beta=\frac{t}{2\pi m}$ are never integers, where $m\in\Big{[}\frac{T}{(1+\varepsilon)\log^{2}T},T(1+\varepsilon)\Big{]}$ is an integer and $M$ is the largest integer $\leq m\varepsilon-1$.
-Thus $v\neq\alpha$ in particular.
-Let $K_{0}$ be the largest integer $<\alpha$.
-We also have
-$\sum_{1\leq v\leq\alpha}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $=\sum_{1\leq v\leq K_{0}}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ (25)
-$\ll\sum_{1\leq v\leq K_{0}}\frac{\beta-\alpha}{(\alpha-v)(\beta-v)}$ (26)
-$\ll(\beta-\alpha)\int_{1}^{K_{0}}\frac{1}{(\alpha-v)(\beta-v)}\mathrm{d}v$ (27)
-$\ll(\beta-\alpha)\log t$ (28)
-$\sim\frac{t\varepsilon}{2\pi m}\log t.$ (29)
-Let $K_{1}$ be the least integer $>\beta$. Note that
-$\sum_{\beta<v\leq N}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $\ll\sum_{K_{1}\leq v\leq N}\frac{\beta-\alpha}{(v-\alpha)(v-\beta)}$ (30)
-$\ll(\beta-\alpha)\int_{K_{1}}^{N}\frac{1}{(v-\alpha)(v-\beta)}\mathrm{d}v$ (31)
-$\ll(\beta-\alpha)\log t$ (32)
-$\sim\frac{t\varepsilon}{2\pi m}\log t.$ (33)
-Since $\beta-\alpha\sim\frac{t\varepsilon}{2\pi m}\ll\frac{\log^{2}T}{T^{1/4}}=o(1)$, note that the range $\alpha<v\leq\beta$ contains no integers. Thus
-$\sum_{\alpha<v\leq\beta}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x=0.$ (34)
-Combining (34), (33), (29), (24) and (18) yields $\sum_{a<r\leq b}e^{2\pi if(r)}=o(1)$.
-Infact, by adapting the argument between (18) and (34), one finds that $\sum_{0<r\leq J}e^{2\pi if(r)}=o(1)$ uniformly for $11\leq J\leq b=m\varepsilon-$ 1.
-Thus by Abel summation it follows that
-$\sum_{a<r\leq b}\frac{1}{r}e^{2\pi if(r)}=\sum_{0<r\leq b}\frac{1}{r}e^{it\log(1+\frac{r}{m})}=o(1).$ (35)
-Inserting (35) into the right-hand side of (17), summing over $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$ and applying the Prime Number Theorem that $\sum_{m\leq x}\Lambda(m)\sim x,$ yields
-$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=o(T).$ (36)
-For $\frac{m}{1+\varepsilon}\leq n\leq m-1$, put $n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ thus $\log(n/m)=-\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=-\frac{m}{r}(1+O(r/m))=-\frac{m}{r}+O(1)$.
+Let $J(u)=\sum_{0<r\leq u}e^{2\pi if(r)}=\sum_{0<r\leq M}e^{2\pi if(r)}$ where $u\leq m\varepsilon,M=M(u)$ is the largest integer $\leq u$, $f(r)=\frac{T\log(1+\frac{r}{m})}{2\pi}$ and $m\in\mathbb{N}$.
+Then $f^{\prime}(r)=\frac{T}{2\pi(r+m)}$.
+Put $\alpha=f^{\prime}(M),\beta=f^{\prime}(0)$ and choose $T\in\mathbb{Q}$ so that both $\alpha$ and $\beta$ are not integers.
+Note that $\beta-\alpha\ll\frac{\log^{2}T}{T^{1/4}}=o(1)$.
+Thus if $\eta$ is any constant $\in(0,\frac{1}{2})$, then the range $[\alpha-\eta,\beta+\eta]$ doesn’t contain any integer.
+Thus $\sum_{\alpha-\eta\leq v\leq\beta+\eta}\int_{0}^{u}e^{2\pi i(f(x)-vx)}\mathrm{d}x=0$ and by Lemma 4.7 of Titchmarsh [7] we have
+$J(u)=\sum_{\alpha-\eta\leq v\leq\beta+\eta}\int_{0}^{u}e^{2\pi i(f(x)-vx)}\mathrm{d}x+O(1)\ll 1$ (18)
+uniformly for $1\leq u\leq m\varepsilon.$ Then
+$\sum_{T^{1/4}<r\leq m\varepsilon-1}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}$ $=\sum_{T^{1/4}<r\leq m\varepsilon-1}\frac{1}{r}e^{2\pi if(r)}$ (19)
+[…]
+[…]
+Since $J(u)\ll 1$ uniformly for $1\leq u\leq m\varepsilon,$ it follows from (21) that
+$\sum_{T^{1/4}<r\leq m\varepsilon-1}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}\ll T^{-1/4}=\varepsilon$. Applying this into (17) yields
+$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}+O(\varepsilon\log m).$ (22)
+Putting $\log(1+\frac{r}{m})=\frac{r}{m}+O((r/m)^{2})$ and $e^{O(\frac{r^{2}\log^{4}T}{m})}=1+O(\frac{r^{2}\log^{4}T}{m})$ into the right-hand side of (22) gives
+$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}+O(\varepsilon\log m).$ (23)
+Inserting (23) into (17) and summing over $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$, yields
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}$ (24)
+$+O(T^{3/4}\log T).$ (25)
+For $\frac{m}{1+\varepsilon}\leq n\leq m-1,$ put $n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ so that $\log(n/m)=-\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=-\frac{m}{r}(1+O(r/m))=-\frac{m}{r}+O(1)$.
 Thus by arguing as above, one arrives at
-$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=o(T).$ (37)
-Summing (9) for $\frac{T}{\log^{2}T}<n<T$, changing the order of summation, invoking (37), (36), (14), (13) and (12) yields
-$\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+o(T)$ (38)
-$=-\frac{T}{\pi}\log\log T+o(T)$ (39)
-since $H=2T+O(1)$ and $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant [5, p.182]. By Lemma 2, we have
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=-\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{-\frac{irT}{m}}$ (26)
+$+O(T^{3/4}\log T).$ (27)
+Combining (27) and (25) gives
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ (28)
+$=\frac{1}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}+O(T^{3/4}\log T).$ (29)
+Let $\{{y\}}$ denote the fractional part of $y$. Recall that if $u$ is not an integer, then $\sum_{n\leq X}\frac{\sin(2\pi nu)}{n}=\pi(\frac{1}{2}-\{{u\}}+O(1/X))$ as $X\rightarrow\infty$. Hence
+$\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}=\sum_{1\leq r\leq T^{1/4}}\frac{2i\sin(rT/m)}{r}=2\pi i\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O(T^{-1/4})\Big{)}.$ (30)
+Plugging (30) into (29) gives
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ (31)
+$=\frac{2\pi i}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O(T^{-1/4})\Big{)}+O(T^{3/4}\log T).$ (32)
+Note that $\sum_{m\leq x}\Lambda(m)\{{x/m\}}=(1-B)x+o(x)$ where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant [5, p.249].
+Also note that $\{{\frac{T}{2\pi m}\}}=\frac{T}{2\pi m}$ for $\frac{T}{2\pi}<m\leq T(1+\varepsilon)$.
+Thus
+$\sum_{m\leq\frac{T}{(1+\varepsilon)\log^{2}T}}\Lambda(m)\{{T/2\pi m\}}-\sum_{m\leq\frac{T}{2\pi}}\Lambda(m)\{{T/2\pi m\}}-\sum_{\frac{T}{2\pi}<m\leq T(1+\varepsilon)}\Lambda(m)\{{T/2\pi m\}}$ (33)
+$+\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\Big{(}\frac{1}{2}+O(T^{-1/4})\Big{)}$ (34)
+$=(B-1)\frac{T}{2\pi}-\frac{T}{2\pi}\sum_{\frac{T}{2\pi}<m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{m}+\frac{T}{2}+o(T)$ (35)
+$=(B-1)\frac{T}{2\pi}-\frac{T}{2\pi}(\log T-\log(T/2\pi))+\frac{T}{2}+o(T)$ (36)
+$=(B-1)\frac{T}{2\pi}-\frac{T}{2\pi}\log(2\pi)+\frac{T}{2}+o(T)$ (37)
+since $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ [5, p.187]. Plugging (37) into (32) gives
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ (38)
+$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{T}{2\pi}-\frac{T}{2\pi}\log(2\pi)+\frac{T}{2}+o(T)\Big{)}.$ (39)
+By a similar argument, one arrives at
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}$ (40)
+$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{H}{2\pi}-\frac{H}{2\pi}\log(2\pi)+\frac{T}{2}+o(T)\Big{)}$ (41)
+$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{T}{\pi}-\frac{T}{\pi}\log(2\pi)+\frac{T}{2}+o(T)\Big{)}$ (42)
+since $H=2T+O(1)$. Combining (42) and (39) gives
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ (43)
+$-\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}$ (44)
+$=\frac{2\pi i}{e}\Big{(}(1-B)\frac{T}{2\pi}+\frac{T}{2\pi}\log(2\pi)+o(T)\Big{)}.$ (45)
+Summing (9) for $\frac{T}{\log^{2}T}<n<T$, changing the order of summation, invoking (45), (14), (13) and (12) yields
+$\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+(1-B)\frac{T}{2\pi}+\frac{T}{2\pi}\log(2\pi)+o(T)$ (46)
+$=-\frac{T}{\pi}\log\log T+(1-B)\frac{T}{2\pi}+\frac{T}{2\pi}\log(2\pi)+o(T).$ (47)
+By Lemma 2, we have
 $\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ (48)
@@
 $\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{2}T}}+\sum_{\frac{T}{\log^{2}T}<n<T}\Big{)}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ (51)
-$\gg T.$ (44)
+$=\frac{T}{2\pi}\log(4\pi)+o(T).$ (52)
 Suppose that $\Theta<1$. Then by Lemma 1 we have

### v43

@@
 By Lemma 12.2 of [5], we know that there exists some $X_{0}\in[X,X+1]$ such that $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX_{0})\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$.
-Thus we can choose $T$ and $H\in[2T,2T+1]$ such that
+Thus we can choose $T$ and $H\in[3T,3T+1]$ such that
 $\max\Big{(}\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iT)\Big{|},\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iH)\Big{|}\Big{)}\ll\log^{2}T$ (1)
@@
 Note that $\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m=2}^{\infty}\frac{\Lambda(m)}{m^{s}}$ for $\sigma>1$. Inserting this into $I_{1}$ gives
-$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}$ (9)
-say.
+$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
 From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise).
@@
 Then $f^{\prime}(r)=\frac{T}{2\pi(r+m)}$.
-Put $\alpha=f^{\prime}(M),\beta=f^{\prime}(0)$ and choose $T\in\mathbb{Q}$ so that both $\alpha$ and $\beta$ are not integers.
-Note that $\beta-\alpha\ll\frac{\log^{2}T}{T^{1/4}}=o(1)$.
-Thus if $\eta$ is any constant $\in(0,\frac{1}{2})$, then the range $[\alpha-\eta,\beta+\eta]$ doesn’t contain any integer.
-Thus $\sum_{\alpha-\eta\leq v\leq\beta+\eta}\int_{0}^{u}e^{2\pi i(f(x)-vx)}\mathrm{d}x=0$ and by Lemma 4.7 of Titchmarsh [7] we have
+Put $\alpha=f^{\prime}(M),\beta=f^{\prime}(0)$ and note that $\beta-\alpha\ll\frac{\log^{2}T}{T^{1/4}}=o(1)$ uniformly for $0\leq u\leq m\varepsilon$.
+Thus if $\eta$ is any constant $\in(0,1)$, then the range $[\alpha-\eta,\beta+\eta]$ contains at most two integers.
+By Lemma 4.2 of Titchmarsh [7], we have $\Big{|}\int_{0}^{u}e^{2\pi i(f(x)-vx)}\mathrm{d}x\Big{|}\leq\frac{4}{f^{\prime}(M)}\ll 1.$ Thus $\sum_{\alpha-\eta\leq v\leq\beta+\eta}\int_{0}^{u}e^{2\pi i(f(x)-vx)}\mathrm{d}x\ll 1$ and by Lemma 4.7 of the same book we have
 $J(u)=\sum_{\alpha-\eta\leq v\leq\beta+\eta}\int_{0}^{u}e^{2\pi i(f(x)-vx)}\mathrm{d}x+O(1)\ll 1$ (18)
@@
 Putting $\log(1+\frac{r}{m})=\frac{r}{m}+O((r/m)^{2})$ and $e^{O(\frac{r^{2}\log^{4}T}{m})}=1+O(\frac{r^{2}\log^{4}T}{m})$ into the right-hand side of (22) gives
-$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}+O(\varepsilon\log m).$ (23)
+$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}+O(\varepsilon\log^{6}T).$ (23)
 Inserting (23) into (17) and summing over $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$, yields
 $\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}$ (24)
-$+O(T^{3/4}\log T).$ (25)
+$+O(T^{3/4}\log^{6}T).$ (25)
 For $\frac{m}{1+\varepsilon}\leq n\leq m-1,$ put $n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ so that $\log(n/m)=-\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=-\frac{m}{r}(1+O(r/m))=-\frac{m}{r}+O(1)$.
@@
 $\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=-\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{-\frac{irT}{m}}$ (26)
-$+O(T^{3/4}\log T).$ (27)
+$+O(T^{3/4}\log^{6}T).$ (27)
 Combining (27) and (25) gives
 $\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ (28)
-$=\frac{1}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}+O(T^{3/4}\log T).$ (29)
+$=\frac{1}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}+O(T^{3/4}\log^{6}T).$ (29)
 Let $\{{y\}}$ denote the fractional part of $y$. Recall that if $u$ is not an integer, then $\sum_{n\leq X}\frac{\sin(2\pi nu)}{n}=\pi(\frac{1}{2}-\{{u\}}+O(1/X))$ as $X\rightarrow\infty$. Hence
@@
 $\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ (31)
-$=\frac{2\pi i}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O(T^{-1/4})\Big{)}+O(T^{3/4}\log T).$ (32)
+$=\frac{2\pi i}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O(T^{-1/4})\Big{)}+O(T^{3/4}\log^{6}T).$ (32)
 Note that $\sum_{m\leq x}\Lambda(m)\{{x/m\}}=(1-B)x+o(x)$ where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant [5, p.249].
@@
 $=(B-1)\frac{T}{2\pi}-\frac{T}{2\pi}\sum_{\frac{T}{2\pi}<m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{m}+\frac{T}{2}+o(T)$ (35)
-$=(B-1)\frac{T}{2\pi}-\frac{T}{2\pi}(\log T-\log(T/2\pi))+\frac{T}{2}+o(T)$ (36)
+$=(B-1)\frac{T}{2\pi}-\frac{T}{2\pi}(\log T-\log(T/2\pi)+O(\varepsilon))+\frac{T}{2}+o(T)$ (36)
 $=(B-1)\frac{T}{2\pi}-\frac{T}{2\pi}\log(2\pi)+\frac{T}{2}+o(T)$ (37)
@@
 $\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}$ (40)
-$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{H}{2\pi}-\frac{H}{2\pi}\log(2\pi)+\frac{T}{2}+o(T)\Big{)}$ (41)
-$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{T}{\pi}-\frac{T}{\pi}\log(2\pi)+\frac{T}{2}+o(T)\Big{)}$ (42)
-since $H=2T+O(1)$. Combining (42) and (39) gives
+$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{H}{2\pi}-\frac{H}{2\pi}\log(3/2\pi)+\frac{T}{2}+o(T)\Big{)}$ (41)
+$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{3T}{2\pi}-\frac{3T}{2\pi}\log(3/2\pi)+\frac{T}{2}+o(T)\Big{)}$ (42)
+since $H=3T+O(1)$. Combining (42) and (39) gives
 $\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ (43)
 $-\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}$ (44)
-$=\frac{2\pi i}{e}\Big{(}(1-B)\frac{T}{2\pi}+\frac{T}{2\pi}\log(2\pi)+o(T)\Big{)}.$ (45)
-Summing (9) for $\frac{T}{\log^{2}T}<n<T$, changing the order of summation, invoking (45), (14), (13) and (12) yields
-$\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+(1-B)\frac{T}{2\pi}+\frac{T}{2\pi}\log(2\pi)+o(T)$ (46)
-$=-\frac{T}{\pi}\log\log T+(1-B)\frac{T}{2\pi}+\frac{T}{2\pi}\log(2\pi)+o(T).$ (47)
+$=\frac{2\pi i}{e}\Big{(}(1-B)\frac{T}{\pi}+\frac{3T}{2\pi}\log 3-\frac{2T}{\pi}\log(2\pi)+o(T)\Big{)}.$ (45)
+Summing (9) for $\frac{T}{\log^{2}T}<n\leq T$, changing the order of summation, invoking (45), (14), (13) and (12) yields
+$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+(1-B)\frac{T}{\pi}+\frac{3T}{2\pi}\log 3-\frac{2T}{\pi}\log(2\pi)+o(T)$ (46)
+$=-\frac{2T}{\pi}\log\log T+(1-B)\frac{T}{\pi}+\frac{3T}{2\pi}\log 3-\frac{2T}{\pi}\log(2\pi)+o(T).$ (47)
 By Lemma 2, we have
-$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ (48)
+$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{\pi n}+O((\log T)\log\log 3n)$ (48)
 uniformly for $2\leq n\ll T$.
@@
 Note that $N(y)=(y/2\pi)\log(y/2\pi)-y/2\pi+O(\log y)$ [5, Corollary 14.3].
-Thus, summing (48) for $1\leq n\leq\frac{T}{\log^{2}T}$ gives
-$\sum_{1\leq n\leq\frac{T}{\log^{2}T}}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=N(H)-N(T)-\frac{T}{2\pi}\sum_{2\leq n\leq\frac{T}{\log^{2}T}}\frac{\Lambda(n)}{n}+o(T)$ (49)
-$=\frac{T}{\pi}\log\log T+\frac{T}{2\pi}(B-1+\log 2)+o(T).$ (50)
+Thus, $N(H)-N(T)=N(3T)-N(T)+O(\log T)=(T/\pi)\log(T/2\pi)+(T/2\pi)(-2+3\log 3)+O(\log T)$ and $(T/\pi)\sum_{2\leq m\leq\frac{T}{\log^{2}T}}\frac{\Lambda(m)}{m}=(T/\pi)\log T-(2T/\pi)\log\log T-(B/\pi)T+o(T).$ Hence, summing (48) for $1\leq n\leq\frac{T}{\log^{2}T}$ gives
+$\sum_{1\leq n\leq\frac{T}{\log^{2}T}}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=N(H)-N(T)-\frac{T}{\pi}\sum_{2\leq n\leq\frac{T}{\log^{2}T}}\frac{\Lambda(n)}{n}+o(T)$ (49)
+$=\frac{2T}{\pi}\log\log T+\frac{T}{2\pi}(3\log 3-2\log(2\pi)+2B-2)+o(T).$ (50)
 Combining (50), (47), (8) and (3) yields
-$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{2}T}}+\sum_{\frac{T}{\log^{2}T}<n<T}\Big{)}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ (51)
-$=\frac{T}{2\pi}\log(4\pi)+o(T).$ (52)
+$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{2}T}}+\sum_{\frac{T}{\log^{2}T}<n\leq T}\Big{)}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ (51)
+$\gg T.$ (52)
 Suppose that $\Theta<1$. Then by Lemma 1 we have

### v44

@@
 By Lemma 12.2 of [5], we know that there exists some $X_{0}\in[X,X+1]$ such that $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX_{0})\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$.
-Thus we can choose $T$ and $H\in[3T,3T+1]$ such that
+Thus we can choose $T$ and $H\in[2T,2T+1]$ such that
 $\max\Big{(}\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iT)\Big{|},\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iH)\Big{|}\Big{)}\ll\log^{2}T$ (1)
 uniformly for $-1\leq\sigma\leq 2$. By the residue theorem, we have
-$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{1+\frac{1}{\log n}+iT}^{1+\frac{1}{\log n}+iH}+\int_{1+\frac{1}{\log n}+iH}^{-1+iH}+\int_{-1+iH}^{-1+iT}+\int_{-1+iT}^{1+\frac{1}{\log n}+iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
+$\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{1+\frac{1}{\log 2n}+iT}^{1+\frac{1}{\log 2n}+iH}+\int_{1+\frac{1}{\log 2n}+iH}^{-1+iH}+\int_{-1+iH}^{-1+iT}+\int_{-1+iT}^{1+\frac{1}{\log 2n}+iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)(2n)^{s-1}\mathrm{d}s$ (2)
 $=I_{1}+I_{2}+I_{3}+I_{4},$ (3)
@@
 Thus since $|\zeta(-it)|\asymp t^{1/2}|\zeta(1+it)|\ll T^{1/2}\log T$, it follows by taking $\sigma=0$ in (4) that
-$S(x):=\sum_{n\leq x}n^{it}\ll T^{\frac{1}{2}}\log^{3}T$ (5)
-uniformly for $\frac{T}{\log^{2}T}\leq x\leq T$. Let $f(n)=n^{\sigma-1}.$ Then by Abel summation, it follows that
-$\sum_{n\leq x}n^{\sigma-1+it}=S(T)f(T)-S(T/\log^{2}T)f(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)f^{\prime}(x)\mathrm{d}x\ll T^{\sigma-\frac{1}{2}}\log^{7}T$ (6)
-for $-1\leq\sigma\leq 2.$ Combining (1) and (6) gives $\sum_{\frac{T}{\log^{2}T}\leq n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{7}T.$ By the log-derivative of the functional equation for $\zeta$, we have
+$S(x):=\sum_{n\leq x}(2n)^{it}\ll T^{\frac{1}{2}}\log^{3}T$ (5)
+uniformly for $\frac{T}{2\log^{2}T}\leq x\leq T$. Let $f(n)=(2n)^{\sigma-1}.$ Then by Abel summation, it follows that
+$\sum_{\frac{T}{\log^{2}T}\leq 2n\leq T}(2n)^{\sigma-1+it}=S(T/2)f(T/2)-S(T/2\log^{2}T)f(T/2\log^{2}T)-\int_{T/2\log^{2}T}^{T/2}S(x)f^{\prime}(x)\mathrm{d}x\ll\frac{\log^{7}T}{T^{\frac{1}{2}-\sigma}}$ (6)
+for $-1\leq\sigma\leq 2.$ Combining (1) and (6) gives $\sum_{\frac{T}{\log^{2}T}\leq 2n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{8}T.$ By the log-derivative of the functional equation for $\zeta$, we have
 $\frac{\zeta^{\prime}}{\zeta}(-\sigma+it)\ll\log T$ (7)
-for $\sigma\geq\sigma_{0}>0$. Combining (6) and (7) yields $\sum_{\frac{T}{\log^{2}T}\leq n\leq T}I_{3}\ll T^{1/2}\log^{7}T.$ Thus we arrive at
-$\sum_{\frac{T}{\log^{2}T}\leq n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{7}T.$ (8)
+for $\sigma\geq\sigma_{0}>0$. Combining (6) and (7) yields $\sum_{\frac{T}{\log^{2}T}\leq 2n\leq T}I_{3}\ll T^{1/2}\log^{8}T.$ Thus we arrive at
+$\sum_{\frac{T}{\log^{2}T}\leq 2n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{7}T.$ (8)
 Note that $\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m=2}^{\infty}\frac{\Lambda(m)}{m^{s}}$ for $\sigma>1$. Inserting this into $I_{1}$ gives
-$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
-From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise).
-For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$.
-For $m\neq n$, let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=\frac{m^{-1-\frac{1}{\log n}}(\log(n)\log(m)-\log^{2}(n)-\log^{2}(m))}{n\log^{2}(n)\log^{2}(n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
-$\sum_{\frac{T}{\log^{2}T}\leq n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}$ $=S(T)g(T)-S(T/\log^{2}T)g(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)g^{\prime}(x)\mathrm{d}x$ (10)
-$\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}-T^{1/2}\log^{3}T\int_{T/\log^{2}T}^{T}g^{\prime}(x)\mathrm{d}x$ (11)
+$I_{1}$ $=\frac{(T-H)\Lambda(2n)}{4\pi n}+\frac{e}{2\pi i}\sum_{m=2,m\neq 2n}^{\infty}\Bigg{(}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}-\frac{\Lambda(m)(2n)^{iH}}{m^{1+\frac{1}{\log 2n}+iH}\log(2n/m)}\Bigg{)}.$ (9)
+From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}\leq 2n\leq T$ (unless specified otherwise).
+For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{|\log(2n/m)|}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$.
+For $m\neq 2n$, let $g(n)=\frac{1}{m^{1+\frac{1}{\log 2n}}\log(2n/m)}$ hence $g^{\prime}(n)=\frac{m^{-1-\frac{1}{\log 2n}}(\log(2n)\log(m)-\log^{2}(2n)-\log^{2}(m))}{n\log^{2}(2n)\log^{2}(2n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
+$\sum_{\frac{T}{\log^{2}T}\leq 2n\leq T}\frac{(2n)^{it}}{m^{1+\frac{1}{\log 2n}}\log(2n/m)}$ $=S(T/2)g(T/2)-S(T/2\log^{2}T)g(T/2\log^{2}T)-\int_{T/2\log^{2}T}^{T/2}S(x)g^{\prime}(x)\mathrm{d}x$ (10)
+$\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}-T^{1/2}\log^{3}T\int_{T/2\log^{2}T}^{T/2}g^{\prime}(x)\mathrm{d}x$ (11)
 $\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (12)
-uniformly for $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon).$ For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll 1/\varepsilon=T^{1/4}$ so it similarly follows from (5) by Abel summation that
-$\sum_{\frac{T}{\log^{2}T}\leq n\leq\frac{m}{1+\varepsilon}\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (13)
+uniformly for $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon).$ For $2n\leq\frac{m}{1+\varepsilon}$ and $2n\geq m(1+\varepsilon)$, $\frac{1}{|\log(2n/m)|}$ is also uniformly $\ll 1/\varepsilon=T^{1/4}$ so it similarly follows from (5) by Abel summation that
+$\sum_{\frac{T}{\log^{2}T}\leq 2n\leq\frac{m}{1+\varepsilon}\leq T}\frac{(2n)^{it}}{m^{1+\frac{1}{\log 2n}}\log(2n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (13)
 and
-$\sum_{\frac{T}{\log^{2}T}\leq m(1+\varepsilon)\leq n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (14)
-We now treat the ranges $m+1\leq n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq n\leq m-1$, where $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$.
-For $m+1\leq n\leq m(1+\varepsilon)$, put $n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(n/m)=\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=\frac{m}{r}(1+O(r/m))=\frac{m}{r}+O(1)$ hence
-$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=\frac{\Lambda(m)}{m^{iT}}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{n^{iT}}{m^{1+\frac{1}{\log n}}\log(n/m)}$ (15)
+$\sum_{\frac{T}{\log^{2}T}\leq m(1+\varepsilon)\leq 2n\leq T}\frac{(2n)^{it}}{m^{1+\frac{1}{\log 2n}}\log(2n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (14)
+We now treat the ranges $m+1\leq 2n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq 2n\leq m-1$, where $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$.
+For $m+1\leq 2n\leq m(1+\varepsilon)$, put $2n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(2n/m)=\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(2n/m)}=\frac{m}{r}(1+O(r/m))=\frac{m}{r}+O(1)$ hence
+$\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\frac{\Lambda(m)}{m^{iT}}\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{(2n)^{iT}}{m^{1+\frac{1}{\log 2n}}\log(2n/m)}$ (15)
 $=\frac{\Lambda(m)}{em^{iT}}\sum_{0<r\leq\varepsilon m-1}\Big{(}\frac{(m+r)^{iT}}{r}+O(1/m)\Big{)}$ (16)
@@
 $\sum_{T^{1/4}<r\leq m\varepsilon-1}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}\ll T^{-1/4}=\varepsilon$. Applying this into (17) yields
-$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}+O(\varepsilon\log m).$ (22)
+$\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}+O(\varepsilon\log m).$ (22)
 Putting $\log(1+\frac{r}{m})=\frac{r}{m}+O((r/m)^{2})$ and $e^{O(\frac{r^{2}\log^{4}T}{m})}=1+O(\frac{r^{2}\log^{4}T}{m})$ into the right-hand side of (22) gives
-$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}+O(\varepsilon\log^{6}T).$ (23)
+$\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}+O(\varepsilon\log^{6}T).$ (23)
 Inserting (23) into (17) and summing over $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$, yields
-$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}$ (24)
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}$ (24)
 $+O(T^{3/4}\log^{6}T).$ (25)
-For $\frac{m}{1+\varepsilon}\leq n\leq m-1,$ put $n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ so that $\log(n/m)=-\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=-\frac{m}{r}(1+O(r/m))=-\frac{m}{r}+O(1)$.
+For $\frac{m}{1+\varepsilon}\leq 2n\leq m-1,$ put $2n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ so that $\log(2n/m)=-\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(2n/m)}=-\frac{m}{r}(1+O(r/m))=-\frac{m}{r}+O(1)$.
 Thus by arguing as above, one arrives at
-$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=-\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{-\frac{irT}{m}}$ (26)
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=-\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{-\frac{irT}{m}}$ (26)
 $+O(T^{3/4}\log^{6}T).$ (27)
 Combining (27) and (25) gives
-$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ (28)
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ (28)
 $=\frac{1}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}+O(T^{3/4}\log^{6}T).$ (29)
-Let $\{{y\}}$ denote the fractional part of $y$. Recall that if $u$ is not an integer, then $\sum_{n\leq X}\frac{\sin(2\pi nu)}{n}=\pi(\frac{1}{2}-\{{u\}}+O(1/X))$ as $X\rightarrow\infty$. Hence
+Let $\{{y\}}$ denote the fractional part of $y$.
+Recall that if $u$ is not an integer, then $\sum_{n\leq X}\frac{\sin(2\pi nu)}{n}=\pi(\frac{1}{2}-\{{u\}}+O(1/X))$ as $X\rightarrow\infty$.
+Hence if $T\in\mathbb{Q}$ so that $T/2\pi m$ is irrational for $m\in\mathbb{N}$, we have
 $\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}=\sum_{1\leq r\leq T^{1/4}}\frac{2i\sin(rT/m)}{r}=2\pi i\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O(T^{-1/4})\Big{)}.$ (30)
 Plugging (30) into (29) gives
-$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ (31)
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ (31)
 $=\frac{2\pi i}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O(T^{-1/4})\Big{)}+O(T^{3/4}\log^{6}T).$ (32)
@@
 since $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ [5, p.187]. Plugging (37) into (32) gives
-$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ (38)
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(2n/m)}$ (38)
 $=\frac{2\pi i}{e}\Big{(}(B-1)\frac{T}{2\pi}-\frac{T}{2\pi}\log(2\pi)+\frac{T}{2}+o(T)\Big{)}.$ (39)
 By a similar argument, one arrives at
-$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}$ (40)
-$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{H}{2\pi}-\frac{H}{2\pi}\log(3/2\pi)+\frac{T}{2}+o(T)\Big{)}$ (41)
-$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{3T}{2\pi}-\frac{3T}{2\pi}\log(3/2\pi)+\frac{T}{2}+o(T)\Big{)}$ (42)
-since $H=3T+O(1)$. Combining (42) and (39) gives
-$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ (43)
-$-\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}$ (44)
-$=\frac{2\pi i}{e}\Big{(}(1-B)\frac{T}{\pi}+\frac{3T}{2\pi}\log 3-\frac{2T}{\pi}\log(2\pi)+o(T)\Big{)}.$ (45)
-Summing (9) for $\frac{T}{\log^{2}T}<n\leq T$, changing the order of summation, invoking (45), (14), (13) and (12) yields
-$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+(1-B)\frac{T}{\pi}+\frac{3T}{2\pi}\log 3-\frac{2T}{\pi}\log(2\pi)+o(T)$ (46)
-$=-\frac{2T}{\pi}\log\log T+(1-B)\frac{T}{\pi}+\frac{3T}{2\pi}\log 3-\frac{2T}{\pi}\log(2\pi)+o(T).$ (47)
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)(2n)^{iH}}{m^{1+\frac{1}{\log 2n}+iH}\log(2n/m)}$ (40)
+$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{H}{2\pi}+\frac{H}{2\pi}\log\pi+\frac{T}{2}+o(T)\Big{)}$ (41)
+$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{T}{\pi}+\frac{T}{\pi}\log\pi+\frac{T}{2}+o(T)\Big{)}$ (42)
+since $H=2T+O(1)$. Combining (42) and (39) gives
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ (43)
+$-\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)(2n)^{iH}}{m^{1+\frac{1}{\log 2n}+iH}\log(2n/m)}$ (44)
+$=\frac{2\pi i}{e}\Big{(}(1-B)\frac{T}{\pi}+\frac{T}{\pi}\log(\pi/2)+o(T)\Big{)}.$ (45)
+Note that $\sum_{m\leq x}\frac{\Lambda(2m)}{2m}=\Big{(}\sum_{m=1}^{\infty}-\sum_{m\geq x+1}\Big{)}\frac{\Lambda(2m)}{2m}=\log 2+O(1/x)$.
+Therefore, summing (9) for $\frac{T}{\log^{2}T}<2n\leq T$, changing the order of summation, invoking (45), (14), (13) and (12) yields
+$\sum_{\frac{T}{\log^{2}T}<2n\leq T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<2n\leq T}\frac{\Lambda(2n)}{2n}+(1-B)\frac{T}{\pi}+\frac{T}{\pi}\log(\pi/2)+o(T)$ (46)
+$=(1-B)\frac{T}{\pi}+\frac{T}{\pi}\log(\pi/2)+o(T)$ (47)
 By Lemma 2, we have
-$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{\pi n}+O((\log T)\log\log 3n)$ (48)
-uniformly for $2\leq n\ll T$.
-Define $N(y):=\sum_{0<\gamma\leq y}1$.
-Note that $N(y)=(y/2\pi)\log(y/2\pi)-y/2\pi+O(\log y)$ [5, Corollary 14.3].
-Thus, $N(H)-N(T)=N(3T)-N(T)+O(\log T)=(T/\pi)\log(T/2\pi)+(T/2\pi)(-2+3\log 3)+O(\log T)$ and $(T/\pi)\sum_{2\leq m\leq\frac{T}{\log^{2}T}}\frac{\Lambda(m)}{m}=(T/\pi)\log T-(2T/\pi)\log\log T-(B/\pi)T+o(T).$ Hence, summing (48) for $1\leq n\leq\frac{T}{\log^{2}T}$ gives
-$\sum_{1\leq n\leq\frac{T}{\log^{2}T}}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=N(H)-N(T)-\frac{T}{\pi}\sum_{2\leq n\leq\frac{T}{\log^{2}T}}\frac{\Lambda(n)}{n}+o(T)$ (49)
-$=\frac{2T}{\pi}\log\log T+\frac{T}{2\pi}(3\log 3-2\log(2\pi)+2B-2)+o(T).$ (50)
+$\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}=-\frac{T\Lambda(2n)}{4\pi n}+O((\log T)\log\log 6n)$ (48)
+uniformly for $1\leq n\ll T$. Hence, summing (48) for $1\leq n\leq\frac{T}{2\log^{2}T}$ gives
+$\sum_{1\leq n\leq\frac{T}{2\log^{2}T}}\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}$ $=-\frac{T}{2\pi}\sum_{1\leq n\leq\frac{T}{2\log^{2}T}}\frac{\Lambda(2n)}{2n}+o(T)$ (49)
+$=-\frac{T}{2\pi}\log 2+o(T).$ (50)
 Combining (50), (47), (8) and (3) yields
-$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{2}T}}+\sum_{\frac{T}{\log^{2}T}<n\leq T}\Big{)}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ (51)
-$\gg T.$ (52)
+$\sum_{1\leq n\leq\frac{T}{2}}\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}$ $=\Big{(}\sum_{1\leq n\leq\frac{T}{2\log^{2}T}}+\sum_{\frac{T}{2\log^{2}T}<n\leq\frac{T}{2}}\Big{)}\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}$ (51)
+$=(1-B+\log\pi)\frac{T}{\pi}+\frac{T}{2\pi}\log 2+o(T)$ (52)
 Suppose that $\Theta<1$. Then by Lemma 1 we have
-$\sum_{1\leq n\leq T}n^{\rho_{T}-1}\ll T^{\Theta-1}$ (53)
+$\sum_{1\leq n\leq\frac{T}{2}}(2n)^{\rho_{T}-1}=2^{\rho_{T}-1}\sum_{1\leq n\leq\frac{T}{2}}n^{\rho_{T}-1}\ll T^{\Theta-1}$ (53)
 uniformly for $T\leq\gamma_{T}\leq H.$ Hence
-$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}\ll T^{\Theta}\log T.$ (54)
+$\sum_{1\leq n\leq\frac{T}{2}}\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq\frac{T}{2}}(2n)^{\rho_{T}-1}\ll T^{\Theta}\log T.$ (54)
 But (54) is a contradiction to (52), thus our supposition must be false. This completes the proof. ∎
@@
 It is therefore instructive to sum (9) carefully, e.g.
-by not rushing to naively bound $n^{it}$ by 1 for $n$ not too close to $m$.
-Doing so will definitely make one lose important information needed for a sufficiently sharp bound for the sums over $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$.
-Note that the analogue of (48) from the Riemann explicit formula is $\sum_{T\leq\gamma_{T}\leq H}\frac{n^{\rho_{T}-1}}{\rho_{T}}\ll\frac{(\log T)(\log\log T)}{T}.$ This has no main term, so it’s not quite useful for our purposes.
+by not rushing to naively bound $(2n)^{it}$ by 1 for $2n$ not too close to $m$.
+Doing so will definitely make one lose important information needed for a sufficiently sharp bound for the sums over $2n\leq\frac{m}{1+\varepsilon}$ and $2n\geq m(1+\varepsilon)$.
+Note that the analogue of (48) from the Riemann explicit formula is $\sum_{T\leq\gamma_{T}\leq H}\frac{(2n)n^{\rho_{T}-1}}{\rho_{T}}\ll\frac{(\log T)(\log\log T)}{T}.$ This has no main term, so it’s not quite useful for our purposes.
 Remark 4.

### v45

@@
 $=\frac{1}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}+O(T^{3/4}\log^{6}T).$ (29)
-Let $\{{y\}}$ denote the fractional part of $y$.
-Recall that if $u$ is not an integer, then $\sum_{n\leq X}\frac{\sin(2\pi nu)}{n}=\pi(\frac{1}{2}-\{{u\}}+O(1/X))$ as $X\rightarrow\infty$.
-Hence if $T\in\mathbb{Q}$ so that $T/2\pi m$ is irrational for $m\in\mathbb{N}$, we have
-$\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}=\sum_{1\leq r\leq T^{1/4}}\frac{2i\sin(rT/m)}{r}=2\pi i\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O(T^{-1/4})\Big{)}.$ (30)
-Plugging (30) into (29) gives
+Let $\{{y\}}$ denote the fractional part of $y$ and $\mid\mid y\mid\mid$ denote the distance to the nearest integer to $y$. Note that if $u>0$ is not an integer, we have
+$\sum_{n\leq X}\frac{\sin(2\pi nu)}{n}=\pi\Big{(}\frac{1}{2}-\{{u\}}+O\Big{(}\frac{1}{1+||u||X}\Big{)}\Big{)}$ (30)
+as $X\rightarrow\infty$. Hence if $T\in\mathbb{Q}$ so that $T/2\pi m$ is irrational for $m\in\mathbb{N}$, we have
+$\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}=\sum_{1\leq r\leq T^{1/4}}\frac{2i\sin(rT/m)}{r}=2\pi i\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O\Big{(}\frac{1}{1+||T/2\pi||T^{1/4}}\Big{)}\Big{)}.$ (31)
+Actually for $20<||u||=u<1/$ 2, the error term in (30) is $\ll\frac{1}{u(X+\frac{1}{2})}$ and for $||u||\geq 1/2$ it remains as it is 222see e.g.
+https://math.stackexchange.com/questions/3331510/error-estimation-for-the-fourier-series-of-the-fractional-part-of-x.
+Thus, applying (31) into (30) with $u=T/2\pi m$ and splitting the error term for $20<||T/2\pi m||<1/$ 2 and $||T/2\pi m||\geq 1/2$ gives
 $\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ (32)
-$=\frac{2\pi i}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O(T^{-1/4})\Big{)}+O(T^{3/4}\log^{6}T).$ (32)
+$=\frac{2\pi i}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\Big{(}\frac{1}{2}-\{{T/2\pi m\}}\Big{)}+O(T^{3/4}\log^{6}T).$ (33)
 Note that $\sum_{m\leq x}\Lambda(m)\{{x/m\}}=(1-B)x+o(x)$ where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant [5, p.249].
@@
 since $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ [5, p.187]. Plugging (38) into (33) gives
-$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(2n/m)}$ (38)
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ (39)
 $=\frac{2\pi i}{e}\Big{(}(B-1)\frac{T}{2\pi}-\frac{T}{2\pi}\log(2\pi)+\frac{T}{2}+o(T)\Big{)}.$ (40)
@@
 $\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)(2n)^{iH}}{m^{1+\frac{1}{\log 2n}+iH}\log(2n/m)}$ (41)
-$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{H}{2\pi}+\frac{H}{2\pi}\log\pi+\frac{T}{2}+o(T)\Big{)}$ (41)
-$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{T}{\pi}+\frac{T}{\pi}\log\pi+\frac{T}{2}+o(T)\Big{)}$ (42)
+$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{H}{2\pi}-\frac{H}{2\pi}\log\pi+\frac{T}{2}+o(T)\Big{)}$ (42)
+$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{T}{\pi}-\frac{T}{\pi}\log\pi+\frac{T}{2}+o(T)\Big{)}$ (43)
 since $H=2T+O(1)$. Combining (43) and (40) gives
@@
 $-\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)(2n)^{iH}}{m^{1+\frac{1}{\log 2n}+iH}\log(2n/m)}$ (45)
-$=\frac{2\pi i}{e}\Big{(}(1-B)\frac{T}{\pi}+\frac{T}{\pi}\log(\pi/2)+o(T)\Big{)}.$ (45)
+$=\frac{2\pi i}{e}\Big{(}(1-B)\frac{T}{2\pi}+\frac{T}{2\pi}\log(\pi/2)+o(T)\Big{)}.$ (46)
 Note that $\sum_{m\leq x}\frac{\Lambda(2m)}{2m}=\Big{(}\sum_{m=1}^{\infty}-\sum_{m\geq x+1}\Big{)}\frac{\Lambda(2m)}{2m}=\log 2+O(1/x)$.
 Therefore, summing (9) for $\frac{T}{\log^{2}T}<2n\leq T$, changing the order of summation, invoking (45), (14), (13) and (12) yields
-$\sum_{\frac{T}{\log^{2}T}<2n\leq T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<2n\leq T}\frac{\Lambda(2n)}{2n}+(1-B)\frac{T}{\pi}+\frac{T}{\pi}\log(\pi/2)+o(T)$ (46)
-$=(1-B)\frac{T}{\pi}+\frac{T}{\pi}\log(\pi/2)+o(T)$ (47)
+$\sum_{\frac{T}{\log^{2}T}<2n\leq T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<2n\leq T}\frac{\Lambda(2n)}{2n}+(1-B)\frac{T}{2\pi}+\frac{T}{2\pi}\log(\pi/2)+o(T)$ (47)
+$=(1-B)\frac{T}{2\pi}+\frac{T}{2\pi}\log(\pi/2)+o(T).$ (48)
 By Lemma 2, we have
@@
 $\sum_{1\leq n\leq\frac{T}{2}}\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}$ $=\Big{(}\sum_{1\leq n\leq\frac{T}{2\log^{2}T}}+\sum_{\frac{T}{2\log^{2}T}<n\leq\frac{T}{2}}\Big{)}\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}$ (52)
-$=(1-B+\log\pi)\frac{T}{\pi}+\frac{T}{2\pi}\log 2+o(T)$ (52)
-Suppose that $\Theta<1$. Then by Lemma 1 we have
+$=(1-B+\log\pi)\frac{T}{2\pi}-\frac{T}{\pi}\log 2+o(T)$ (53)
+for the sequence of $T,H\in[2T,2T+1]$ that satisfy (1). Suppose that $\Theta<1$. Then by Lemma 1 we have
 $\sum_{1\leq n\leq\frac{T}{2}}(2n)^{\rho_{T}-1}=2^{\rho_{T}-1}\sum_{1\leq n\leq\frac{T}{2}}n^{\rho_{T}-1}\ll T^{\Theta-1}$ (54)
 uniformly for $T\leq\gamma_{T}\leq H.$ Hence
-$\sum_{1\leq n\leq\frac{T}{2}}\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq\frac{T}{2}}(2n)^{\rho_{T}-1}\ll T^{\Theta}\log T.$ (54)
-But (54) is a contradiction to (52), thus our supposition must be false. This completes the proof. ∎
+$\sum_{1\leq n\leq\frac{T}{2}}\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq\frac{T}{2}}(2n)^{\rho_{T}-1}\ll T^{\Theta}\log T$ (55)
+uniformly for $T\geq T_{0}$. But (55) is a contradiction to (53), thus our supposition must be false. This completes the proof. ∎
 Corollary 1. There exists an infinite sequence of zeros of $\zeta$, whose real parts converge to 1.
@@
 Doing so will definitely make one lose important information needed for a sufficiently sharp bound for the sums over $2n\leq\frac{m}{1+\varepsilon}$ and $2n\geq m(1+\varepsilon)$.
-Note that the analogue of (48) from the Riemann explicit formula is $\sum_{T\leq\gamma_{T}\leq H}\frac{(2n)n^{\rho_{T}-1}}{\rho_{T}}\ll\frac{(\log T)(\log\log T)}{T}.$ This has no main term, so it’s not quite useful for our purposes.
+Note that the analogue of (49) from the Riemann explicit formula is $\sum_{T\leq\gamma_{T}\leq H}\frac{(2n)^{\rho_{T}-1}}{\rho_{T}}\ll\frac{(\log T)(\log\log T)}{T}.$ This has no main term, so it’s not quite as elegant/useful as (49).
 Remark 4.

### v46

@@
 for $\sigma\geq\sigma_{0}>0$. Combining (6) and (7) yields $\sum_{\frac{T}{\log^{2}T}\leq 2n\leq T}I_{3}\ll T^{1/2}\log^{8}T.$ Thus we arrive at
-$\sum_{\frac{T}{\log^{2}T}\leq 2n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{7}T.$ (8)
+$\sum_{\frac{T}{\log^{2}T}\leq 2n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{8}T.$ (8)
 Note that $\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m=2}^{\infty}\frac{\Lambda(m)}{m^{s}}$ for $\sigma>1$. Inserting this into $I_{1}$ gives
@@
 Let $\{{y\}}$ denote the fractional part of $y$ and $\mid\mid y\mid\mid$ denote the distance to the nearest integer to $y$. Note that if $u>0$ is not an integer, we have
-$\sum_{n\leq X}\frac{\sin(2\pi nu)}{n}=\pi\Big{(}\frac{1}{2}-\{{u\}}+O\Big{(}\frac{1}{1+||u||X}\Big{)}\Big{)}$ (30)
-as $X\rightarrow\infty$. Hence if $T\in\mathbb{Q}$ so that $T/2\pi m$ is irrational for $m\in\mathbb{N}$, we have
-$\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}=\sum_{1\leq r\leq T^{1/4}}\frac{2i\sin(rT/m)}{r}=2\pi i\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O\Big{(}\frac{1}{1+||T/2\pi||T^{1/4}}\Big{)}\Big{)}.$ (31)
-Actually for $20<||u||=u<1/$ 2, the error term in (30) is $\ll\frac{1}{u(X+\frac{1}{2})}$ and for $||u||\geq 1/2$ it remains as it is 222see e.g.
-https://math.stackexchange.com/questions/3331510/error-estimation-for-the-fourier-series-of-the-fractional-part-of-x.
-Thus, applying (31) into (30) with $u=T/2\pi m$ and splitting the error term for $20<||T/2\pi m||<1/$ 2 and $||T/2\pi m||\geq 1/2$ gives
+$\sum_{n\leq X}\frac{\sin(2\pi nu)}{n}=\pi\Big{(}\frac{1}{2}-\{{u\}}+O\Big{(}\frac{1}{1+X||u||}\Big{)}\Big{)}$ (30)
+as $X\rightarrow\infty$.
+222see e.g.
+https://math.stackexchange.com/questions/3331510/error-estimation-for-the-fourier-series-of-the-fractional-part-of-x Hence if $T\in\mathbb{Q}$ so that $T/2\pi m$ is irrational for $m\in\mathbb{N}$, we have
+$\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}=\sum_{1\leq r\leq T^{1/4}}\frac{2i\sin(rT/m)}{r}=2\pi i\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O\Big{(}\frac{1}{1+T^{1/4}||T/2\pi m||}\Big{)}\Big{)}.$ (31)
+Let $U_{m},V_{m}\in\mathbb{N},T=\frac{U_{m}}{V_{m}}$ and $m\mid U_{m}$.
+By Dirichlet’s theorem on rational approximations to irrationals, note that $\frac{1}{2\pi}$ can be approximated by infinitely many irreducible rational numbers $\frac{p_{m}}{q_{m}}$ with error $O({q_{m}}^{-2})$.
+Let: $V_{m}\mid p_{m}$, $\gcd(q_{m},U_{m})=1,\frac{U_{m}}{mV_{m}{q_{m}}^{2}}=o(1/q_{m})$ and $q_{m}\ll(\log T)^{6}$ uniformly for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$.
+Note that: $\frac{(\frac{U_{m}}{m})(\frac{p_{m}}{V_{m}})}{q_{m}}$ is irreducible, $\Big{|}\Big{|}\frac{(\frac{U_{m}}{m})(\frac{p_{m}}{V_{m}})}{q_{m}}\Big{|}\Big{|}\gg\frac{1}{q_{m}}$ and $\frac{T}{2\pi m}=\frac{(\frac{U_{m}}{m})(\frac{p_{m}}{V_{m}})}{q_{m}}+O\Big{(}\frac{U_{m}}{mV_{m}{q_{m}}^{2}}\Big{)}=\frac{(\frac{U_{m}}{m})(\frac{p_{m}}{V_{m}})}{q_{m}}+o(1/q_{m})$.
+Hence $||T/2\pi m||\gg\frac{1}{q_{m}}\gg(\log T)^{-6}$ uniformly for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$.
+Inserting (31) into (29) and applying the fact that $||T/2\pi m||\gg(\log T)^{-6}$, gives
 $\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ (32)
@@
 $\sum_{m\leq\frac{T}{(1+\varepsilon)\log^{2}T}}\Lambda(m)\{{T/2\pi m\}}-\sum_{m\leq\frac{T}{2\pi}}\Lambda(m)\{{T/2\pi m\}}-\sum_{\frac{T}{2\pi}<m\leq T(1+\varepsilon)}\Lambda(m)\{{T/2\pi m\}}$ (34)
-$+\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\Big{(}\frac{1}{2}+O(T^{-1/4})\Big{)}$ (35)
+$+\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{2}$ (35)
 $=(B-1)\frac{T}{2\pi}-\frac{T}{2\pi}\sum_{\frac{T}{2\pi}<m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{m}+\frac{T}{2}+o(T)$ (36)
@@
 $\sum_{1\leq n\leq\frac{T}{2}}\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}$ $=\Big{(}\sum_{1\leq n\leq\frac{T}{2\log^{2}T}}+\sum_{\frac{T}{2\log^{2}T}<n\leq\frac{T}{2}}\Big{)}\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}$ (52)
-$=(1-B+\log\pi)\frac{T}{2\pi}-\frac{T}{\pi}\log 2+o(T)$ (53)
-for the sequence of $T,H\in[2T,2T+1]$ that satisfy (1). Suppose that $\Theta<1$. Then by Lemma 1 we have
+$=(1-B+\log\pi)\frac{T}{2\pi}+o(T)$ (53)
+for the sequence of rational $T,H\in[2T,2T+1]$ that satisfy (1). Suppose that $\Theta<1$. Then by Lemma 1 we have
 $\sum_{1\leq n\leq\frac{T}{2}}(2n)^{\rho_{T}-1}=2^{\rho_{T}-1}\sum_{1\leq n\leq\frac{T}{2}}n^{\rho_{T}-1}\ll T^{\Theta-1}$ (54)
-uniformly for $T\leq\gamma_{T}\leq H.$ Hence
+uniformly for $T\leq\gamma_{T}\leq H.$ For large $x>0$, note that there are $\asymp x\log x$ zeros of $\zeta$ with imaginary parts in $(0,x]$ [5, Corollary 14.3]. Hence
 $\sum_{1\leq n\leq\frac{T}{2}}\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq\frac{T}{2}}(2n)^{\rho_{T}-1}\ll T^{\Theta}\log T$ (55)
@@
 Combine Theorem 1 with the fact that $\zeta(s)\neq 0$ for $\Re(s)\geq 1$ [5, Theorem 6.6]. ∎
-Remark 1. Notice that our argument actually shows that, for all large enough $T\in\mathbb{R}^{+}$, there exists some $\rho_{T}$ with $T\leq\gamma_{T}\leq 2T+1$ and $\Re(\rho_{T})=1-o(1)$.
+Remark 1. Notice that our argument actually shows that, for all large enough $T\in\mathbb{Q}^{+}$, there exists some $\rho_{T}$ with $T\leq\gamma_{T}\leq 2T+1$ and $\Re(\rho_{T})=1-o(1)$.
 Remark 2. Our argument crucially relies on:
@@
 Remark 4.
-The reader might wonder if our approach can give a lower bound for $\Theta_{T}:$ the supremum of the real parts of the zeros $\rho_{T}$ with $T\leq\gamma_{T}\leq H.$ For that, one would need Theorem 4.11 of Titchmarsh [7] to also hold for $\sigma=\sigma_{t}=o(1)$ with an error term $\ll(\log t)^{-1}$.
-However, this is not possible (as can be seen from the approximate functional equation for $\zeta$).
+The reader might wonder if our approach can give a lower bound for $\Theta_{T}:$ the supremum of the real parts of the zeros $\rho_{T}$ with $T\leq\gamma_{T}\leq H.$ For that, one would need our Lemma 1 to also hold for $\sigma=\sigma_{t}=o(1)$ with an error term $\ll(\log t)^{-1}$.
+However, this is not possible (as can be seen e.g by estimating $\zeta(\sigma+it)=\sum_{n\leq\frac{T}{2}}n^{-\sigma-it}=\sum_{n\leq T/2}n^{(-\sigma+\sigma_{t})-\sigma_{t}-it}$ for any fixed $\sigma>0$ via Abel summation, under the assumption that $\sum_{n\leq T/2}n^{-\sigma_{t}-it}=\zeta(\sigma_{t}+it)+O(1/\log t)$).
 It can also be seen that Titchmarsh’s proof of his Theorem 4.11 doesn’t hold for $\sigma=\sigma_{t}=o(1)$.
@@
 Recall that a function can only be holomorphically extended on open sets, and that’s precisely why Titchmarsh assumes that $\sigma$ is bounded below by some constant $\sigma_{0}>0$.
+For the same reason, one cannot apply any of the best known zero-free regions of $\zeta$ (e.g.
+Vinogradov-Koborov) to bound $\Re(\rho_{T})$ in (54).

### v47

@@
 ABSTRACT.
-Let $\Theta$ denote the supremum of the real parts of the zeros of the Riemann zeta function.
-We demonstrate that $\Theta=1$, which entails the existence of infinitely many Riemann zeros off the critical line (thus disproving the Riemann Hypothesis (RH)).
-The paper is concluded by a brief discussion of why our argument doesn’t work for both Weil and Beurling zeta functions whose analogues of the RH are known to be true.
+Let $\Theta$ denote the supremum of the real parts of the zeros of the Riemann zeta function $\zeta(s)$.
+By carefully exploiting the Landau-Gonek explicit formula, the Dirichlet polynomial approximation of $\zeta(s)$ and other classical results in analytic number theory, we rigorously demonstrate that $\Theta=1$.
+This entails the existence of infinitely many Riemann zeros off the critical line (thus disproving the Riemann Hypothesis (RH)).
+The paper is concluded by a brief discussion of why our argument does not work for both Weil and Beurling zeta functions, whose analogues of the RH are known to be true.
 2020 Mathematics Subject Classifications: 11M26, 11M06.
@@
 Euler noticed that for $\Re(s)>1$, $\zeta(s)$ can be expressed as a product $\prod_{p}(1-p^{-s})^{-1}$ over the entire set of primes, which entails that $\zeta(s)\neq 0$ for $\Re(s)>1$.
-It can be shown that $\zeta(s)$ extends to $\mathbb{C}$ as a meromorphic function with only a simple pole at $s=1$, with residue $1$, and satisfies the functional equation $\xi(s)=\xi(1-s)$, where $\xi(z)=\frac{1}{2}z(z-1)\pi^{-z/2}\Gamma(\frac{1}{2}z)\zeta(z)$ and $\Gamma(w)=\int_{0}^{\infty}e^{-x}x^{w-1}\mathrm{d}x$.
+It can be shown that $\zeta(s)$ extends to $\mathbb{C}$ as a meromorphic function with only a simple pole at $s=1$, with residue 11, and satisfies the functional equation $\xi(s)=\xi(1-s)$, where $\xi(z)=\frac{1}{2}z(z-1)\pi^{-z/2}\Gamma(\frac{1}{2}z)\zeta(z)$ and $\Gamma(w)=\int_{0}^{\infty}e^{-x}x^{w-1}\mathrm{d}x$.
 Let $\chi$ be a primitive Dirichlet character and $L(s,\chi)$ be the corresponding L-function.
@@
 For a far more thorough discussion of the RH and GRH, the interested reader is kindly referred to Borwein et al [1].
-Definitions. Let: $\rho$ denote a non-trivial zero of $\zeta,\sigma=\Re(s)$, $\gamma=\Im(\rho)$ and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s.
+Definitions.
+Let: $\rho$ denote a non-trivial zero of $\zeta,\sigma=\Re(s)$, $\gamma=\Im(\rho),\Lambda$ be the von-Mangoldt function and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s.
 Lemma 1.
 (Theorem 4.11 of [7]).
-We have $\zeta(s)=\sum_{n\leq x}n^{-s}-\frac{x^{1-s}}{1-s}+O(x^{-\sigma})$ uniformly for $\sigma\geq\sigma_{0}>0$ and $|\Im(s)|<2\pi x/C$, where $C$ is a given constant $>1$.
+We have $\zeta(s)=\sum_{n\leq x}n^{-s}-\frac{x^{1-s}}{1-s}+O(x^{-\sigma})$ uniformly for $\sigma\geq\sigma_{0}>0$ and $|\Im(s)|<2\pi x/C$, where $C$ is a given constant >1>1.
 Lemma 2, (Theorem 1 of [3]). One has $\sum_{0\leq\gamma\leq y}n^{\rho-1}=-\frac{y\Lambda(n)}{2\pi n}+O((\log y)\log\log 3n)$ uniformly for $2\leq n\ll y.$
@@
 Let: $T\in\mathbb{R}_{>0}$ be arbitrarily large and $X\geq 2$.
-By Lemma 12.2 of [5], we know that there exists some $X_{0}\in[X,X+1]$ such that $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX_{0})\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$.
+By Lemma 12.2 of [5], we know that there exists some $X_{0}\in[X,X+1]$ such that $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX_{0})\ll\log^{2}X$ uniformly for $1\leq\sigma\leq 2$.
 Thus we can choose $T$ and $H\in[2T,2T+1]$ such that
 $\max\Big(\Big|\frac{\zeta^{\prime}}{\zeta}(\sigma+iT)\Big|,\Big|\frac{\zeta^{\prime}}{\zeta}(\sigma+iH)\Big|\Big)\ll\log^{2}T$ (1)
-uniformly for $-1\leq\sigma\leq 2$. By the residue theorem, we have
+uniformly for $1\leq\sigma\leq 2$. By the residue theorem, we have
 $\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg(\int_{1+\frac{1}{\log 2n}+iT}^{1+\frac{1}{\log 2n}+iH}+\int_{1+\frac{1}{\log 2n}+iH}^{-1+iH}+\int_{-1+iH}^{-1+iT}+\int_{-1+iT}^{1+\frac{1}{\log 2n}+iT}\Bigg)\frac{\zeta^{\prime}}{\zeta}(s)(2n)^{s-1}\mathrm{d}s$ (2)
@@
 $\zeta(\sigma-it)=\sum_{n\leq x}n^{-\sigma+it}+\chi(\sigma-it)\sum_{v\leq y}v^{\sigma-1-it}+O(x^{-\sigma}+t^{\frac{1}{2}-\sigma}y^{\sigma-1})$ (4)
-where $00\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h>$ 0 is a constant, $\chi(s)=\frac{\zeta(s)}{\zeta(1-s)}$ and $|\chi(\sigma-it)|=|\chi(\sigma+it)|=\Big{|}\Big{(}\frac{t}{2\pi}\Big{)}^{\sigma-\frac{1}{2}-it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{\frac{1}{2}-\sigma}$ [7, p.78].
+where $0\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h>0$ is a constant, $\chi(s)=\frac{\zeta(s)}{\zeta(1-s)}$ and $|\chi(\sigma-it)|=|\chi(\sigma+it)|=\Big|\Big(\frac{t}{2\pi}\Big)^{\frac{1}{2}-\sigma-it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big|\sim(t/2\pi)^{\frac{1}{2}-\sigma}$ [7, p.78].
 Thus since $|\zeta(-it)|\asymp t^{1/2}|\zeta(1+it)|\ll T^{1/2}\log T$, it follows by taking $\sigma=0$ in (4) that
@@
 $\sum_{\frac{T}{\log^{2}T}\leq 2n\leq T}(2n)^{\sigma-1+it}=S(T/2)f(T/2)-S(T/2\log^{2}T)f(T/2\log^{2}T)-\int_{T/2\log^{2}T}^{T/2}S(x)f^{\prime}(x)\mathrm{d}x\ll\frac{\log^{7}T}{T^{\frac{1}{2}-\sigma}}$ (6)
-for $-1\leq\sigma\leq 2.$ Combining (1) and (6) gives $\sum_{\frac{T}{\log^{2}T}\leq 2n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{8}T.$ By the log-derivative of the functional equation for $\zeta$, we have
+for $1\leq\sigma\leq 2.$ Combining (1) and (6) gives $\sum_{\frac{T}{\log^{2}T}\leq 2n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{8}T.$ By the log-derivative of the functional equation for $\zeta$, we have
 $\frac{\zeta^{\prime}}{\zeta}(-\sigma+it)\ll\log T$ (7)
@@
 We now treat the ranges $m+1\leq 2n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq 2n\leq m-1$, where $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$.
-For $m+1\leq 2n\leq m(1+\varepsilon)$, put $2n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(2n/m)=\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(2n/m)}=\frac{m}{r}(1+O(r/m))=\frac{m}{r}+O(1)$ hence
+For $m+1\leq 2n\leq m(1+\varepsilon)$, put $2n=m+r$ where $1\leq r\leq m\varepsilon-1$ thus $\log(2n/m)=\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(2n/m)}=\frac{m}{r}(1+O(r/m))=\frac{m}{r}+O(1)$ hence
 $\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\frac{\Lambda(m)}{m^{iT}}\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{(2n)^{iT}}{m^{1+\frac{1}{\log 2n}}\log(2n/m)}$ (15)
-$=\frac{\Lambda(m)}{em^{iT}}\sum_{0<r\leq\varepsilon m-1}\Big{(}\frac{(m+r)^{iT}}{r}+O(1/m)\Big{)}$ (16)
+$=\frac{\Lambda(m)}{em^{iT}}\sum_{0<r\leq\varepsilon m-1}\Big(\frac{(m+r)^{iT}}{r}+O\Big(\frac{\varepsilon}{r\log m}\Big)\Big)$ (16)
 $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq\varepsilon m-1}\frac{(1+\frac{r}{m})^{iT}}{r}+O(\varepsilon\log m).$ (17)
-Let $J(u)=\sum_{0<r\leq u}e^{2\pi if(r)}=\sum_{0<r\leq M}e^{2\pi if(r)}$ where $u\leq m\varepsilon,M=M(u)$ is the largest integer $\leq u$, $f(r)=\frac{T\log(1+\frac{r}{m})}{2\pi}$ and $m\in\mathbb{N}$.
-Then $f^{\prime}(r)=\frac{T}{2\pi(r+m)}$.
-Put $\alpha=f^{\prime}(M),\beta=f^{\prime}(0)$ and note that $\beta-\alpha\ll\frac{\log^{2}T}{T^{1/4}}=o(1)$ uniformly for $0\leq u\leq m\varepsilon$.
-Thus if $\eta$ is any constant $\in(0,1)$, then the range $[\alpha-\eta,\beta+\eta]$ contains at most two integers.
-By Lemma 4.2 of Titchmarsh [7], we have $\Big{|}\int_{0}^{u}e^{2\pi i(f(x)-vx)}\mathrm{d}x\Big{|}\leq\frac{4}{f^{\prime}(M)}\ll 1.$ Thus $\sum_{\alpha-\eta\leq v\leq\beta+\eta}\int_{0}^{u}e^{2\pi i(f(x)-vx)}\mathrm{d}x\ll 1$ and by Lemma 4.7 of the same book we have
-$J(u)=\sum_{\alpha-\eta\leq v\leq\beta+\eta}\int_{0}^{u}e^{2\pi i(f(x)-vx)}\mathrm{d}x+O(1)\ll 1$ (18)
-uniformly for $1\leq u\leq m\varepsilon.$ Then
-$\sum_{T^{1/4}<r\leq m\varepsilon-1}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}$ $=\sum_{T^{1/4}<r\leq m\varepsilon-1}\frac{1}{r}e^{2\pi if(r)}$ (19)
-[…]
-[…]
-Since $J(u)\ll 1$ uniformly for $1\leq u\leq m\varepsilon,$ it follows from (21) that
-$\sum_{T^{1/4}<r\leq m\varepsilon-1}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}\ll T^{-1/4}=\varepsilon$. Applying this into (17) yields
-$\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}+O(\varepsilon\log m).$ (22)
-Putting $\log(1+\frac{r}{m})=\frac{r}{m}+O((r/m)^{2})$ and $e^{O(\frac{r^{2}\log^{4}T}{m})}=1+O(\frac{r^{2}\log^{4}T}{m})$ into the right-hand side of (22) gives
-$\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}+O(\varepsilon\log^{6}T).$ (23)
-Inserting (23) into (17) and summing over $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$, yields
+We now estimate
+$J(T):=\sum_{T^{1/4}\leq r\leq m\varepsilon}\frac{1}{r}\exp\!\left(iT\log\left(1+\frac{r}{m}\right)\right).$
+Let $\phi(r)=\frac{T}{2\pi}\log(1+\frac{r}{m})$ thus $\phi^{\prime}(r)=\frac{T}{2\pi(r+m)}.$ Let $\eta\in(0,1/2)$ be a constant, $\beta=\phi^{\prime}(T^{1/4})$ and $\alpha=\phi^{\prime}(m\varepsilon)$.
+By Lemma 4.10 of Titchmarsh [7, p.76], we have
+$J(T)=\sum_{\alpha-\eta\leq v\leq\beta+\eta}\int_{T^{1/4}}^{m\varepsilon}\frac{1}{x}e^{2\pi i(\phi(x)-vx)}\mathrm{d}x+O(T^{-1/4}\log T).$ (18)
+Since $\beta-\alpha\ll\varepsilon\log^{2}T=o(1)$, note that the range $\alpha-\eta,\beta+\eta]$ contains at most one integer $v$. Hence (18) implies that
+$J(T)\ll\Bigg|\int_{T^{1/4}}^{m\varepsilon}\frac{1}{x}e^{2\pi i(\phi(x)-vx)}\mathrm{d}x\Bigg|+T^{-1/4}\log T.$ (19)
+Define
+$J_{0}(T)=\int_{T^{1/4}}^{m\varepsilon}\frac{1}{x}e^{2\pi i(\phi(x)-vx)}\mathrm{d}x=\int_{T^{1/4}}^{m\varepsilon}\frac{1}{x}e^{2\pi iT\Big[\log(1+\frac{x}{m})-\frac{vx}{T}\Big]}\mathrm{d}x.\\$ (20)
+Let $G(x)=\frac{1}{x},F(x)=\log(1+\frac{x}{m})-\frac{vx}{T},\lambda=2\pi T$ and $F^{\prime}(x_{0})=0$ so that $\frac{T}{m+x_{0}}=v$ thus $F^{\prime\prime}(x_{0})=-\frac{v^{2}}{T}$ and $|T-mv|\sim\frac{Tx_{0}}{m}$ where $x_{0}\in[T^{1/4},m\varepsilon]$.
+Then $J_{0}(T)$ is an oscillatory integral of the form $\int_{a}^{b}G(x)\exp(i\lambda F(x))\mathrm{d}x$.
+Thus by the stationary phase approximation 222see e.g.
+https://grokipedia.com/page/Stationary phase approximation (the one-dimensional section)., we have
+$J_{0}(T)\sim G(x_{0})e^{i\lambda F(x_{0})-i\pi/4}\sqrt{\frac{2\pi}{\lambda|F^{\prime\prime}(x_{0})|}}\ll G(x_{0})\sqrt{\frac{2\pi}{\lambda|F^{\prime\prime}(x_{0})|}}\ll\frac{1}{|T-mv|}\ll T^{-1/4}$ (21)
+uniformly for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$ as $\lambda=2\pi T\rightarrow\infty$. Combing (19), (20) and (21) yields
+$\sum_{T^{1/4}\leq r\leq m\varepsilon}\frac{1}{r}\exp\!\left(iT\log\left(1+\frac{r}{m}\right)\right)\ll T^{-1/4}\log T$ (22)
+uniformly for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$. Applying (22) into (17) gives
+$\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}+O(\varepsilon\log^{2}T).$ (23)
+Putting $\log(1+\frac{r}{m})=\frac{r}{m}+O((r/m)^{2})$ and $e^{O({Tr^{2}/m^{2}})}=1+O({Tr^{2}/m^{2}})$ into the right-hand side of (23) gives
+$\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}+O(\varepsilon\log^{2}T).$ (24)
+Summing (24) over $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\ \leq T(1+\varepsilon)$ yields
 $\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}$ (25)
-$+O(T^{3/4}\log^{6}T).$ (25)
-For $\frac{m}{1+\varepsilon}\leq 2n\leq m-1,$ put $2n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ so that $\log(2n/m)=-\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(2n/m)}=-\frac{m}{r}(1+O(r/m))=-\frac{m}{r}+O(1)$.
+$+O(T^{3/4}\log^{2}T).$ (26)
+For $\frac{m}{1+\varepsilon}\leq 2n\leq m-1,$ put 2n=mr2n=m-r where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ so that $\log(2n/m)=-\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(2n/m)}=-\frac{m}{r}(1+O(r/m))=-\frac{m}{r}+O(1)$.
 Thus by arguing as above, one arrives at
 $\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=-\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{-\frac{irT}{m}}$ (27)
-$+O(T^{3/4}\log^{6}T).$ (27)
+$+O(T^{3/4}\log^{2}T).$ (28)
 Combining (28) and (26) gives
 $\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big(\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big)\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ (29)
-$=\frac{1}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}+O(T^{3/4}\log^{6}T).$ (29)
-Let $\{{y\}}$ denote the fractional part of $y$ and $\mid\mid y\mid\mid$ denote the distance to the nearest integer to $y$. Note that if $u>0$ is not an integer, we have
-$\sum_{n\leq X}\frac{\sin(2\pi nu)}{n}=\pi\Big{(}\frac{1}{2}-\{{u\}}+O\Big{(}\frac{1}{1+X||u||}\Big{)}\Big{)}$ (30)
+$=\frac{1}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}+O(T^{3/4}\log^{2}T).$ (30)
+Let $\{{y\}}$ denote the fractional part of $y$ and $\mid\mid y\mid\mid$ denote the distance to the integer nearest to $y$. Note that if $u>0$ is not an integer, we have
+[…]
 as $X\rightarrow\infty$.
-222see e.g.
+333see e.g.
 https://math.stackexchange.com/questions/3331510/error-estimation-for-the-fourier-series-of-the-fractional-part-of-x Hence if $T\in\mathbb{Q}$ so that $T/2\pi m$ is irrational for $m\in\mathbb{N}$, we have
-$\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}=\sum_{1\leq r\leq T^{1/4}}\frac{2i\sin(rT/m)}{r}=2\pi i\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O\Big{(}\frac{1}{1+T^{1/4}||T/2\pi m||}\Big{)}\Big{)}.$ (31)
-Let $U_{m},V_{m}\in\mathbb{N},T=\frac{U_{m}}{V_{m}}$ and $m\mid U_{m}$.
-By Dirichlet’s theorem on rational approximations to irrationals, note that $\frac{1}{2\pi}$ can be approximated by infinitely many irreducible rational numbers $\frac{p_{m}}{q_{m}}$ with error $O({q_{m}}^{-2})$.
-Let: $V_{m}\mid p_{m}$, $\gcd(q_{m},U_{m})=1,\frac{U_{m}}{mV_{m}{q_{m}}^{2}}=o(1/q_{m})$ and $q_{m}\ll(\log T)^{6}$ uniformly for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$.
-Note that: $\frac{(\frac{U_{m}}{m})(\frac{p_{m}}{V_{m}})}{q_{m}}$ is irreducible, $\Big{|}\Big{|}\frac{(\frac{U_{m}}{m})(\frac{p_{m}}{V_{m}})}{q_{m}}\Big{|}\Big{|}\gg\frac{1}{q_{m}}$ and $\frac{T}{2\pi m}=\frac{(\frac{U_{m}}{m})(\frac{p_{m}}{V_{m}})}{q_{m}}+O\Big{(}\frac{U_{m}}{mV_{m}{q_{m}}^{2}}\Big{)}=\frac{(\frac{U_{m}}{m})(\frac{p_{m}}{V_{m}})}{q_{m}}+o(1/q_{m})$.
-Hence $||T/2\pi m||\gg\frac{1}{q_{m}}\gg(\log T)^{-6}$ uniformly for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$.
-Inserting (31) into (29) and applying the fact that $||T/2\pi m||\gg(\log T)^{-6}$, gives
+[…]
+Define
+$U(T)=\sum_{V\leq m\leq T(1+\varepsilon)}\frac{1}{1+T^{1/4}\,\Bigl\|\frac{T}{2\pi m}\Bigr\|},\qquad V:=\frac{T}{(1+\varepsilon)\log^{2}T}.$
+Set $Z=T^{1/4}$ and $x_{m}:=\frac{T}{2\pi m}.$ We split the sum $U(T)$ according to the size of $\|x_{m}\|$.
+For $\|x_{m}\|\leq Z^{-1}$, we have $\frac{1}{1+Z\|x_{m}\|}\asymp 1,$ and the corresponding contribution is bounded by
+$\{m\in[V,T(1+\varepsilon)]:\|x_{m}\|\leq Z^{-1}\}.$
+Fix $k\in\mathbb{N}$.
+The condition $\Bigl|\frac{T}{2\pi m}-k\Bigr|\leq Z^{-1}$ is equivalent to $\frac{T}{2\pi(k+Z^{-1})}\leq m\leq\frac{T}{2\pi(k-Z^{-1})}.$ The length of this interval is $\frac{T}{2\pi}\left(\frac{1}{k-Z^{-1}}-\frac{1}{k+Z^{-1}}\right)=\frac{T}{2\pi}\cdot\frac{2Z^{-1}}{k^{2}-Z^{-2}}\ll\frac{T}{Zk^{2}}$.
+Hence, for fixed $k$, the number of such integers $m$ is $\ll\frac{T}{Zk^{2}}+1.$ Since $x_{m}=\frac{T}{2\pi m}\in\Bigl[\frac{1}{2\pi(1+\varepsilon)},\frac{(1+\varepsilon)\log^{2}T}{2\pi}\Bigr],$ we have $1\leq k\ll(\log T)^{2}$.
+Summing over $k$ gives
+$\{m:\|x_{m}\|\leq Z^{-1}\}\ll\sum_{k\leq(\log T)^{2}}\left(\frac{T}{Zk^{2}}+1\right)\ll\frac{T}{Z}.$
+Thus the contribution from the $x_{m}$ for which $\|x_{m}\|\leq Z^{-1}$, is
+$\sum_{\|x_{m}\|\leq Z^{-1}}\frac{1}{1+Z\|x_{m}\|}\ll\frac{T}{Z}=T^{3/4}.$ (33)
+For $\|x_{m}\|>Z^{-1}$ we use $\frac{1}{1+Z\|x_{m}\|}\leq\frac{1}{Z\|x_{m}\|}.$ Partition $(0,1/2]$ dyadically: $2^{-j-1}<\|x_{m}\|\leq 2^{-j},0\leq j\leq\log_{2}Z.$ For such $m$, $\frac{1}{1+Z\|x_{m}\|}\ll\frac{2^{j}}{Z}$.
+Since $x_{m}$ is monotone in $m$, the number of $m$ in each dyadic interval is $\ll T\cdot 2^{-j}.$ Hence the contribution of each $j$ is $\ll T\cdot 2^{-j}\cdot\frac{2^{j}}{Z}=\frac{T}{Z}.$ Summing over $0\leq j\leq\log_{2}Z$ yields
+$\sum_{\|x_{m}\|>Z^{-1}}\frac{1}{1+Z\|x_{m}\|}\ll\frac{T}{Z}\log Z\ll T^{3/4}\log T.$ (34)
+Combining both ranges gives
+$U(T)=\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{1}{1+T^{1/4}\Bigl\|\tfrac{T}{2\pi m}\Bigr\|}\ll T^{3/4}\log T.$ (35)
+Inserting (32) into (30) and applying (35) gives
 $\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big(\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big)\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ (36)
-$=\frac{2\pi i}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\Big{(}\frac{1}{2}-\{{T/2\pi m\}}\Big{)}+O(T^{3/4}\log^{6}T).$ (33)
+$=\frac{2\pi i}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\Big(\frac{1}{2}-\{{T/2\pi m\}}\Big)+O(T^{3/4}\log^{2}T).$ (37)
 Note that $\sum_{m\leq x}\Lambda(m)\{{x/m\}}=(1-B)x+o(x)$ where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant [5, p.249].
@@
 $=(B-1)\frac{T}{2\pi}-\frac{T}{2\pi}\log(2\pi)+\frac{T}{2}+o(T)$ (42)
-since $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ [5, p.187]. Plugging (38) into (33) gives
+since $\sum_{m\leq x}\Lambda(m)\sim x$ and $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ [5, p.187]. Plugging (42) into (37) gives
 $\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big(\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big)\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ (43)
@@
 $\sum_{1\leq n\leq\frac{T}{2}}\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}$ $=\Big(\sum_{1\leq n\leq\frac{T}{2\log^{2}T}}+\sum_{\frac{T}{2\log^{2}T}<n\leq\frac{T}{2}}\Big)\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}$ (56)
-$=(1-B+\log\pi)\frac{T}{2\pi}+o(T)$ (53)
+$=(1-B+\log(\pi/4))\frac{T}{2\pi}+o(T)$ (57)
 for the sequence of rational $T,H\in[2T,2T+1]$ that satisfy (1). Suppose that $\Theta<1$. Then by Lemma 1 we have
@@
 Remark 1. Notice that our argument actually shows that, for all large enough $T\in\mathbb{Q}^{+}$, there exists some $\rho_{T}$ with $T\leq\gamma_{T}\leq 2T+1$ and $\Re(\rho_{T})=1-o(1)$.
-Remark 2. Our argument crucially relies on:
-(i) The functional equation for $\zeta$, and
-(ii) Theorem 4.11 of [7].
+Remark 2.
+Our argument crucially relies on: (i) the functional equation for $\zeta$, and (ii) theorem 4.11 of [7].
 Both (i) and (ii) do not hold for Beurling zeta functions corresponding to Beurling number systems whose integers are not the ordinary integers.
@@
 The reader might wonder if our approach can give a lower bound for $\Theta_{T}:$ the supremum of the real parts of the zeros $\rho_{T}$ with $T\leq\gamma_{T}\leq H.$ For that, one would need our Lemma 1 to also hold for $\sigma=\sigma_{t}=o(1)$ with an error term $\ll(\log t)^{-1}$.
-However, this is not possible (as can be seen e.g by estimating $\zeta(\sigma+it)=\sum_{n\leq\frac{T}{2}}n^{-\sigma-it}=\sum_{n\leq T/2}n^{(-\sigma+\sigma_{t})-\sigma_{t}-it}$ for any fixed $\sigma>0$ via Abel summation, under the assumption that $\sum_{n\leq T/2}n^{-\sigma_{t}-it}=\zeta(\sigma_{t}+it)+O(1/\log t)$).
+However, this is not possible (as can be seen e.g by estimating $\zeta(a+it)=\sum_{n\leq\frac{T}{2}}n^{-a-it}+O(T^{-a})=\sum_{n\leq T/2}n^{(-a+\sigma_{t})-\sigma_{t}-it}+O(T^{-a})$ for any fixed $a>0$ via Abel summation, under the assumption that $\sum_{n\leq T/2}n^{-\sigma_{t}-it}=\zeta(\sigma_{t}+it)+O(1/\log t)$).
 It can also be seen that Titchmarsh’s proof of his Theorem 4.11 doesn’t hold for $\sigma=\sigma_{t}=o(1)$.

### v48

@@
 We now estimate
-$J(T):=\sum_{T^{1/4}\leq r\leq m\varepsilon}\frac{1}{r}\exp\!\left(iT\log\left(1+\frac{r}{m}\right)\right).$
+$J(T):=\sum_{T^{1/4}\leq r\leq m\varepsilon}\frac{1}{r}\exp\!\left(iT\log\left(1+\frac{r}{m}\right)\right)$ (18)
+where $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$.
 Let $\phi(r)=\frac{T}{2\pi}\log(1+\frac{r}{m})$ thus $\phi^{\prime}(r)=\frac{T}{2\pi(r+m)}.$ Let $\eta\in(0,1/2)$ be a constant, $\beta=\phi^{\prime}(T^{1/4})$ and $\alpha=\phi^{\prime}(m\varepsilon)$.
@@
 $J(T)=\sum_{\alpha-\eta\leq v\leq\beta+\eta}\int_{T^{1/4}}^{m\varepsilon}\frac{1}{x}e^{2\pi i(\phi(x)-vx)}\mathrm{d}x+O(T^{-1/4}\log T).$ (19)
-Since $\beta-\alpha\ll\varepsilon\log^{2}T=o(1)$, note that the range $\alpha-\eta,\beta+\eta]$ contains at most one integer $v$. Hence (18) implies that
-$J(T)\ll\Bigg|\int_{T^{1/4}}^{m\varepsilon}\frac{1}{x}e^{2\pi i(\phi(x)-vx)}\mathrm{d}x\Bigg|+T^{-1/4}\log T.$ (19)
-Define
-$J_{0}(T)=\int_{T^{1/4}}^{m\varepsilon}\frac{1}{x}e^{2\pi i(\phi(x)-vx)}\mathrm{d}x=\int_{T^{1/4}}^{m\varepsilon}\frac{1}{x}e^{2\pi iT\Big[\log(1+\frac{x}{m})-\frac{vx}{T}\Big]}\mathrm{d}x.\\$ (20)
-Let $G(x)=\frac{1}{x},F(x)=\log(1+\frac{x}{m})-\frac{vx}{T},\lambda=2\pi T$ and $F^{\prime}(x_{0})=0$ so that $\frac{T}{m+x_{0}}=v$ thus $F^{\prime\prime}(x_{0})=-\frac{v^{2}}{T}$ and $|T-mv|\sim\frac{Tx_{0}}{m}$ where $x_{0}\in[T^{1/4},m\varepsilon]$.
-Then $J_{0}(T)$ is an oscillatory integral of the form $\int_{a}^{b}G(x)\exp(i\lambda F(x))\mathrm{d}x$.
-Thus by the stationary phase approximation 222see e.g.
-https://grokipedia.com/page/Stationary phase approximation (the one-dimensional section)., we have
-$J_{0}(T)\sim G(x_{0})e^{i\lambda F(x_{0})-i\pi/4}\sqrt{\frac{2\pi}{\lambda|F^{\prime\prime}(x_{0})|}}\ll G(x_{0})\sqrt{\frac{2\pi}{\lambda|F^{\prime\prime}(x_{0})|}}\ll\frac{1}{|T-mv|}\ll T^{-1/4}$ (21)
-uniformly for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$ as $\lambda=2\pi T\rightarrow\infty$. Combing (19), (20) and (21) yields
-$\sum_{T^{1/4}\leq r\leq m\varepsilon}\frac{1}{r}\exp\!\left(iT\log\left(1+\frac{r}{m}\right)\right)\ll T^{-1/4}\log T$ (22)
-uniformly for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$. Applying (22) into (17) gives
-$\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}+O(\varepsilon\log^{2}T).$ (23)
-Putting $\log(1+\frac{r}{m})=\frac{r}{m}+O((r/m)^{2})$ and $e^{O({Tr^{2}/m^{2}})}=1+O({Tr^{2}/m^{2}})$ into the right-hand side of (23) gives
-$\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}+O(\varepsilon\log^{2}T).$ (24)
-Summing (24) over $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\ \leq T(1+\varepsilon)$ yields
+Since $\beta-\alpha\ll\varepsilon\log^{2}T=o(1)$, note that the range $\alpha-\eta,\beta+\eta]$ contains at most one integer $v$.
+Summing (19) over $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$ gives
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}J(T)\ll\int_{T^{1/4}}^{T(1+\varepsilon)\varepsilon}\frac{1}{x}\Bigg|\sum_{m=\max\left(\frac{T}{(1+\varepsilon)\log^{2}T},\frac{x}{\varepsilon}\right)}^{T(1+\varepsilon)}e^{2\pi i\big(\phi(x)-vx\big)}\Bigg|\,dx\;+\;T^{3/4}\log T,$ (20)
+where $v=v(m)$ is the integer nearest to $\frac{T}{2\pi m}$.
+The dependence of $v$ on $m$ must be respected, so we write $k=v(m)$.
+The condition $|k-\frac{T}{2\pi m}|\leq\frac{1}{2}$ implies $m\in I_{k}:=\Bigl[\frac{T}{2\pi(k+\frac{1}{2})},\frac{T}{2\pi(k-\frac{1}{2})}\Bigr].$ For $m\geq T/((1+\varepsilon)\log^{2}T)$ we have $0\leq k\ll\log^{2}T$; hence $k$ runs over $O(\log^{2}T)$ nonnegative integers.
+The length of $I_{k}$ is $|I_{k}|\asymp T(k+1)^{-2}$.
+On $I_{k}$ the phase function is $g_{k}(m)=\frac{T}{2\pi}\log\!\Bigl(1+\frac{x}{m}\Bigr)-kx$ where $k$ is fixed.
+Its derivative satisfies $g_{k}^{\prime}(m)=-\frac{Tx}{2\pi m(m+x)}\asymp-\frac{Tx}{m^{2}}$ and $|g_{k}^{\prime\prime}(m)|\asymp\frac{Tx}{m^{3}}.$ Applying the first‑derivative bound of van der Corput (Lemma 4.19 of Titchmarsh [7]) on each interval $I_{k}$ yields
+$\Bigl|\sum_{m\in I_{k}}e^{2\pi ig_{k}(m)}\Bigr|\ll\frac{1}{\lambda_{1}}+|I_{k}|\lambda_{1},\qquad\lambda_{1}:=\frac{Tx}{m^{2}}\asymp\frac{x(k+1)^{2}}{T}.$
+Since $|I_{k}|\asymp\frac{T}{(1+k)^{2}}$, we have
+$\Bigl|\sum_{m\in I_{k}}e^{2\pi ig_{k}(m)}\Bigr|\ll\frac{T}{x}+x$
+Summing over all admissible $k$ where $0\leq k\ll\log^{2}T$, gives
+$\Bigl|\sum_{m}e^{2\pi i(\phi(x)-\nu x)}\Bigr|\ll\sum_{k}\Bigl(\frac{T}{x}+x\Bigr)\ll\frac{T\log^{2}T}{x}+x\log^{2}T.$ (21)
+Inserting this bound into the integral in (20) gives
+$\int_{T^{1/4}}^{T(1+\varepsilon)\varepsilon}\frac{1}{x}\Bigl|\sum_{m}e^{2\pi i(\phi(x)-\nu x)}\Bigr|\,dx\ll\int_{T^{1/4}}^{T^{3/4}}\Bigl(\frac{T\log^{2}T}{x^{2}}+\log^{2}T\Bigr)dx\ll T^{3/4}\log^{2}T.$
+Plugging this into (20) gives
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}J(T)\ll T^{3/4}\log^{2}T.$ (22)
+Combining (20), (21) and (22) yields
+$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{T^{1/4}\leq r\leq m\varepsilon}\frac{1}{r}\exp\!\left(iT\log\left(1+\frac{r}{m}\right)\right)\ll T^{3/4}\log^{2}T.$ (23)
+For
+$1\leq r\leq T^{1/4}$ (24)
+we have $\log(1+\frac{r}{m})=\frac{r}{m}+O((r/m)^{2})$ and $e^{O({Tr^{2}/m^{2}})}=1+O({Tr^{2}/m^{2}})$. Applying this together with (23) into the right-hand side of (17) yields
 $\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}$ (25)
@@
 as $X\rightarrow\infty$.
-333see e.g.
+222see e.g.
 https://math.stackexchange.com/questions/3331510/error-estimation-for-the-fourier-series-of-the-fractional-part-of-x Hence if $T\in\mathbb{Q}$ so that $T/2\pi m$ is irrational for $m\in\mathbb{N}$, we have
@@
 $\sum_{\|x_{m}\|\leq Z^{-1}}\frac{1}{1+Z\|x_{m}\|}\ll\frac{T}{Z}=T^{3/4}.$ (33)
-For $\|x_{m}\|>Z^{-1}$ we use $\frac{1}{1+Z\|x_{m}\|}\leq\frac{1}{Z\|x_{m}\|}.$ Partition $(0,1/2]$ dyadically: $2^{-j-1}<\|x_{m}\|\leq 2^{-j},0\leq j\leq\log_{2}Z.$ For such $m$, $\frac{1}{1+Z\|x_{m}\|}\ll\frac{2^{j}}{Z}$.
+For $\|x_{m}\|>Z^{-1}$ we use $\frac{1}{1+Z\|x_{m}\|}\leq\frac{1}{Z\|x_{m}\|}.$ Partition $(0,1/2]$ dyadically: $2^{-j-1}<\|x_{m}\|\leq 2^{-j}$ where $0\leq j\leq\log_{2}Z.$ For such $m$, $\frac{1}{1+Z\|x_{m}\|}\ll\frac{2^{j}}{Z}$.
 Since $x_{m}$ is monotone in $m$, the number of $m$ in each dyadic interval is $\ll T\cdot 2^{-j}.$ Hence the contribution of each $j$ is $\ll T\cdot 2^{-j}\cdot\frac{2^{j}}{Z}=\frac{T}{Z}.$ Summing over $0\leq j\leq\log_{2}Z$ yields